### v1

# HoneySat: A Network-based Satellite Honeypot Framework

Abstract
Satellites are the backbone of several mission-critical services, such as GPS that enable our modern society to function. For many years, satellites were assumed to be secure because of their indecipherable architectures and the reliance on security by obscurity. However, technological advancements have made these assumptions obsolete, paving the way for potential attacks, and sparking a renewed interest in satellite security. Unfortunately, to this day, there is no efficient way to collect data on adversarial techniques for satellites, which severely hurts the generation of security intelligence.
In this paper, we present HoneySat, the first high-interaction satellite honeypot framework, which is fully capable of convincingly simulating a real-world CubeSat, a type of Small Satellite (SmallSat) widely used in practice. To provide evidence of the effectiveness of HoneySat, we surveyed experienced SmallSat operators currently in charge of active in-orbit satellite missions.
Results revealed that the majority of satellite operators (71.4%) agreed that HoneySat provides realistic and engaging simulations of CubeSat missions. Further experimental evaluations also showed that HoneySat provides adversaries with extensive interaction opportunities by supporting the majority of adversarial techniques (86.8%) and tactics (100%) that target satellites. Additionally, we also obtained a series of real interactions from actual adversaries by deploying HoneySat on the internet over several months, confirming that HoneySat can operate covertly and efficiently while collecting highly valuable interaction data.
1 Introduction
Artificial satellites are complex devices designed to withstand outer space conditions. They serve multiple purposes or types of missions that include position and navigation, e.g., the Global Positioning System (GPS) constellation [74], Earth observation, e.g., the Sentinel constellation [19], and broadband internet service, e.g., the Starlink constellation. In addition, Spacecraft can range from being as massive as thousands of kilograms, such as the International Space Station (ISS) to one kilogram CubeSats. As a result, the software and hardware components that make up a specific spacecraft vary greatly. Likewise, satellite missions’ vary a lot depending on the owner’s budget. For example, university missions are smaller as they have a limited budget. A cyberattack on a satellite or satellite constellation (group of satellites) could have disastrous consequences on a global scale, which is difficult to comprehend. Such an attack could lead to the cessation of air traffic and widespread communication blackouts. It could also cause food shortages and the freezing of financial transactions [11]. Furthermore, such an attack could exacerbate the Kessler Syndrome [18], a scenario in which collisions between satellites and debris in orbit create a cascade effect, generating even more debris, jeopardizing future satellite launches and operations.
In this increasingly vulnerable environment, the probability of a successful satellite cyberattack continues to rise. This is driven by three key trends [28]: first, satellite deployments have increased at an unprecedented pace. For instance, while an average of 82 launches took place between 2008 and 2017, as many as 197 launches occurred in 2023 alone, each typically carrying multiple satellites [61]. This surge is partly fueled by the rise of cheaper commercial off-the-shelf (COTS) components and the availability of more Space Launch Vehicles (SLVs), which makes access to orbit affordable for smaller institutions, such as universities [67]. Second, ground station technology has become significantly more affordable (and sometimes open source), greatly lowering the communication barrier with satellites [10]. Thus, a broader range of malicious actors can now communicate with satellites. Third, satellite engineers and operators continue to rely on protocol obscurity practices, such as hiding specialized knowledge about the transmission protocol implemented on satellites [81].
Although satellite security research has gained increased attention [81, 60, 53, 79], the relationship between outer space and cyberspace remains poorly understood [52]. At the same time, space-focused threat intelligence remains sparse, and our current methods for identifying tactics, techniques, and procedures (TTPs) used against these critical systems are limited. MITRE ATT&CK currently tracks 152 threat groups but shows only one that targets satellites explicitly [71], highlighting a significant data gap. Although the volume of reported cyber incidents in the space sector has grown, these reports rarely provide sufficient detail [11]. As a result, the security community has limited visibility into adversarial activity aimed at space infrastructure.
Honeypots’ ability to collect real-world cyberattack data makes them an ideal solution to this problem [29]. A honeypot is a decoy computer system intended to lure and entice malicious actors to interact with it [12]; all the while, the honeypot logs all the interactions the attackers make. This log data can later be analyzed to discover new and existing TTPs.
Since the release of the first honeypot, the Deception Toolkit, in 1997 [13], a wide range of honeypots with ever-increasing capabilities have been introduced. These honeypots are used by universities, companies, and nation-states worldwide [20, 8, 27, 62, 37]. Honeypots are also used to deter malicious actors from attacking different types of systems, from industrial control systems [37] to social media platforms [3]. However, as of the time of writing, there is no space-sector specific honeypot in the literature.
In this paper, we design, implement, and test the first satellite honeypot, HoneySat to attract and analyze adversaries who attack space infrastructures over the Internet, a commonly observed threat vector [7, 5, 33, 80]. HoneySat is a modular, high-interaction honeypot framework that realistically simulates a complete satellite system (ground infrastructure and Satellite). Specifically, HoneySat simulates Small Satellites or SmallSats which are spacecraft with a mass of less than 180 kilograms[44]. CubeSats for example, are SmallSats. Additionally, as part of HoneySat, we developed the Satellite Simulator, a Python project to provide generic simulation functionality for users to populate satellite honeypots with believable data. This capability enables the creation of complete testing environments for satellite software that integrate a variety of simulated subsystems and sensors. For instance, Satellite Simulator can simulate orbital mechanics (e.g., position tracking, attitude adjustment) and electrical power systems (e.g., power generation, consumption, and distribution) alongside other subsystems.
We leveraged our framework to create honeypots of real-world CubeSat missions. Our results, backed by our survey of satellite operators, show that HoneySat’s simulation is highly realistic. Our framework is able to simulate an entire real satellite mission, provides realistic telemetry, and supports real telecommands.
In summary, this paper makes the following contributions:
  • A novel framework, HoneySat, a high-interaction, extensible honeypot for small satellites (Sec. 4).
  • The Satellite Simulator, that simulates the physical processes, sensors, and subsystems necessary for a realistic satellite honeypot (Sec. 5).
  • The results of a survey of experienced satellite operators that provide valuable insights into how realistically our honeypot performs, as well as experimental evidence demonstrating that the HoneySat framework can simulate small satellites, collect rich real-world interaction data, and be customized for multiple satellites (Sec. 6).
2 Background
This section lays out key background concepts that are relevant to satellite honeypots. For honeypots: a description of the different existing types (Sec. 2.1), as well as the current state-of-the-art (Sec. 2.2). For satellites: their operation (Sec. 2.3), their architecture (Sec. 2.4), existing Protocol Ecosystems (Sec. 2.5), and tactics, techniques, and procedures (TTPs, Sec. 2.6).
2.1 Types of Honeypots
Honeypots are categorized by interaction levels according to the interaction opportunities they provide. The two main types of honeypots are low-interaction and high-interaction.
Low-Interaction Honeypots. These honeypots offer minimal interaction, simulating real systems through scripts or finite-state machines. Their advantages are ease of setup and maintenance due to low resource consumption, and a reduced risk of adversarial takeover. However, they provide limited interaction opportunities to adversaries which limits the interaction data they provide. Low-interaction honeypot examples include Conpot [78] and Honeyd [54].
High-Interaction Honeypots. These honeypots offer extensive interaction opportunities via emulation or advanced simulations [54]. Their main advantage is providing adversaries with almost limitless interactions, enabling them to provide extensive interaction data. However, they pose a high risk of adversarial takeover as adversaries have more opportunities to hijack the honeypot [39]. High-interaction honeypot examples include Cowrie [50] and HoneyPLC [37].
2.2 Honeypot’s State of the Art
The literature on honeypots includes hundreds of implementations that simulate a diverse set of computer systems [30, 48]. From classic implementations that simulate a host’s TCP/IP stack, such as Honeyd [54], to modern approaches that integrate social media applications, such as HoneyTweet [3]. However, there is no satellite honeypot in the literature. In the absence of a satellite honeypot, we now examine the honeypot approaches most related to satellites: Industrial Control Systems (ICS) [36, 35] and Unmanned Aerial Vehicles (UAV), a.k.a., drones summarized in Table 1.
Satellite systems like ICS must be aware of some physical process, e.g., its position, the sun’s position, etc, via sensors to acquire data about the physical world. Several ICS honeypots have simulated these physical processes. For example, ICSnet [58] and HoneyICS [38] simulate several components such as programmable logic controllers (PLCs), and actuators such as water valves.
HoneyDrone [16] is a UAV honeypot that integrates different simulations, e.g., Ardupilot, to recreate multiple attack scenarios, including the UAV ground control station. Although UAV honeypots share some similarities with satellites [75], satellite honeypots require additional physical simulations and involve a more complex operation environment, which we discuss in Sec. 2.3.
Table 1: Comparison of Existing Honeypots and HoneySat.
Keys: = Supported; $\times$× = Not Supported.
Honeypot/
Feature
Interaction
Level
Included
Protocols
Physics
Sims
Extensi-
bility
Conpot [78]
Low 9 0
HoneyPLC[37]
High 3 0
ICSPot [14]
High 4 1 $\times$×
HoneyICS[38]
High 2 1
HoneyDrone [16]
Medium 4 1 $\times$×
HoneySat
High
4
6
Addressed
in Section
4.4,
4.3
4.3,
5.2.1
4.4,
5.1
6.5
2.3 Anatomy of a Satellite Mission
Figure 1: The components commonly found in a satellite mission in the context of the space and ground segments.
We now describe the components of a satellite mission. Due to satellite missions’ complexity, we explain each component and match it with one of the numbers in Fig. 1. Every satellite mission includes the ground segment from which satellite operators control the satellite and the space segment which includes the satellite itself.
Ground Segment 1. The Ground Segment (GS) covers the terrestrial supporting infrastructure required for a successful satellite operation. It consists of a ground station, responsible for exchanging data with the spacecraft, the computational and network infrastructure required for communication, but also the systems to operate the satellites, e.g., servers, databases, and user interfaces [81]. Several ground stations can be connected in a network and coordinated as part of one GS. The GS includes the Ground Segment Software (GSS) that helps operators schedule and send commands and visualize data that is sent back in response.
Space Segment 2. The space segment comprises a satellite or a constellation of satellites. A satellite is launched into orbit and then establishes communications with the ground segment. During regular operations, satellites may communicate through one or multiple ground stations [81].
Telecommands (TC) 3 and Telemetry (TM) 4. The basic data flow between the space and ground segments are TC and TM [69]. TM is the data the satellite sends to the ground station which may contain the satellite’s status, responses to previous commands, or payload data [69]. TCs are used to operate the satellite and are transmitted and encapsulated in a space protocol (see Sec. 2.5). The design and implementation of TCs varies depending on the satellite mission. From a security perspective, TCs are particularly important as an attacker that can send valid TCs to a satellite can fully take over the mission [81].
Orbital Pass 5. A satellite and its ground station can communicate only during an orbital pass. An orbital pass, or more commonly a pass, is when the satellite rises above a ground station’s horizon and becomes available for communication. A pass’s duration and timing depend on the satellite’s orbit characteristics and any obstructing objects between the satellite and the ground station, e.g., mountains [82]. In some cases, such as geostationary orbit (GEO), the satellite can be above the horizon constantly. The occurrence of passes can be calculated using the spacecraft’s orbital elements. A common representation of the orbital elements is the two-line element set (TLE) [77].
Satellite Mission Operations 6. Satellite mission operations vary widely depending on the owning organization, budget, and technology. Nevertheless, they share some commonalities, which we now describe.
A mission’s operation involves a team of operators that use GSS to operate a satellite or a satellite constellation and ensure the mission’s success [46]. Operations are carried out in a Mission Operations Center (MOC), where satellite operators sit at their workstations to manage TCs sent to the satellite(s).
Satellite operators may also remotely operate satellites 7 by connecting to the ground segment using tools such as desktop sharing like VNC (Virtual Network Computing) [10], or even rely on autonomous or semi-autonomous operations [26].
Satellite operations include two main activities: satellite tracking and TC generation and scheduling. Satellite tracking calculates the satellite’s position in orbit and controls the ground station(s)’ tracking antenna to establish communication between the ground and space segments. TC generation and scheduling crafts command to be programmed and sent to the satellite so it can perform different functions, e.g., read systems status data, schedule orbital maneuvers, schedule payload operations, and download payload data.
Ground (Segment) Software (GSS). GSS allows satellite operators to carry out the satellite mission’s routine operations as described in the previous paragraph. GSS are very diverse. Some satellite missions develop their own GSS while others use open-source [65] or buy proprietary GSS like GSWeb [23]. There are two main types of GSS: Mission Control Software (MCS) and Ground Station Control Software (GSCS).
Mission control software 8 manages TCs and scripts to be sent to the satellite and can display TM. For example, ESA’s SCOS-2000 is an MCS that provides generic functionality that can be customized for a specific mission to cover the functions required for TM reception and processing and TC verification [42]. Some satellite missions develop their own MCS. For example, the SUCHAI mission team developed their MCS application to send and receive TC/TM [65].
Ground station control software (GSCS) 9 helps satellite operators track and visualize the satellite’s orbit and provide detailed information about each satellite’s pass. For example, Gpredict is a popular open-source GSCS that performs real-time satellite tracking and orbit prediction [15]. Gpredict can also interface with and control the radio system and antenna rotor during the satellite pass.
2.4 Satellite Architecture
Satellite architectures are varied and complex, however, here we describe the most common terminology depicted in Fig. 1’s space segment 2.
When referring to satellite architecture there is a distinction between the platform that facilitates the successful operation of the satellite’s day to day-to-day activities and the payload. The platform provides the possibility to run a payload that fulfills the purpose of these missions. This payload differs based on the goal of the mission and can range from instruments to measure physical properties to communications systems providing wireless connectivity.
Platform: The platform, or satellite bus, is a system composed of custom-designed or off-the-shelf subsystems necessary for critical satellite operations. These include, among others, the Attitude Determination and Control System (ADCS) to maintain the satellite’s orientation (i.e. attitude) and position to keep the satellite pointed towards antennas or solar panels illuminated; the Electrical Power Subsystem (EPS) for managing power generation and distribution; the Communication Subsystem (COMM) and the Command and Data Handling (C&DH) subsystems to facilitate communications for receiving TC and sending TM and controlling the satellite operations.
All of these subsystems are then controlled via TCs sent to the satellite. TCs may be, depending on the protocol ecosystem and complexity of the subsystem, interpreted by a central C&DH System or merely forwarded to the recipient subsystem for further processing [84]. This managerial duty is left to Flight Software (FS) running on an embedded system of differing complexity. Some examples include NASA’s Core Flight System (cFS) [41] and F’ (F Prime) [6], KubOS [83], the German Aerospace Center (DLR)’s OUTPOST [17], and the University of Chile’s SUCHAI FS [24]. Attackers aim to gain the ability to send Telecommands to the Flight Software, as this typically grants control over all spacecraft subsystems.
Payload. The payload is the equipment that the satellite employs to fulfill its mission. Due to satellites’ varied missions, payloads are heavily customized [70]. For example, if a satellite’s mission is remote sensing, its payload may include an infrared camera [45]. craft’s essential functions, acts on TCs, and handles internal data transmitted to and from other systems [84]. The CDHS is comprised of Flight Software (FS) running on an OBC [81].
2.5 Small Satellite Protocol Ecosystems
SmallSat missions can often be categorized into the following categories by the adoption of protocols and their corresponding philosophies.
Cubesat Space Protocol. The Cubesat Space Protocol (CSP) family of protocols is a one-stop solution for Small missions [63]. There are not a lot of choices left to the operator/developer and there are two vendors mainly offering components using this protocol and some that offer compatibility solutions [23]. If a mission built on CSP is expanded by a commercial subsystem it will plug in without issue after configuration.
CSP is implemented as an open-source C library called libCSP [63]. CSP follows the TCP/IP model, including transport and routing protocols and multiple layer 2 interfaces such as I2C (Inter-Integrated Circuit), CAN (Controller Area Network), and ZeroMQ (ZMQ) for transmission on TCP/IP networks [72].
In CSP the ground segment interfaces and satellite subsystems are part of a CSP network. Sending TC is as simple as sending a CSP packet with an address corresponding to a node that is part of the satellite. The configured static routing tables of the nodes on the network will then cause the packet to be passed to the ground station and transmitted to the satellite over RF where a system on board will route the packet to its destination. In case the satellite is not currently doing a pass, the packet is dropped.
The straightforwardness of CSP makes it particularly interesting for the honeypot use case as the internal structures of any CSP-based mission will look very similar, so distinguishing two missions is difficult as they will use the protocols in a very similar way.
There are also default services running on standardized ports that are enabled by default when building libCSP [63], for example for network diagnostic (ping) or basic operations (rebooting).
In CSP missions, nodes on the network typically feature a command-line-based interface for configuration and debugging purposes [23]. This may allow someone with access to it to interact with the current node and other nodes on the network.
CCSDS Space Communication Protocols. The CCSDS Space Communication protocols are a vast set of standardized protocols used for different purposes in space communications. A major CCSDS protocol relevant for SmallSat TM and TC is called spacepacket. This protocol is used in combination with the ECSS Packet Utilization Standard (PUS) to define how TCs and TM are encoded and transported. The PUS defines services (and thus sets of TC/TM) for functionality that satellite missions likely require, including large data transfer or event reporting [59]. Moreover, mission designers can tailor the PUS standard by selecting a subset of services and sub-services relevant to their mission needs. Furthermore, it is possible to define custom Services or implement numerous custom functionalities.
Proprietary Protocols. In addition to these widespread ecosystems, vendor-specific sets of protocols may be usable on top of more standard protocols or transport higher-layer packets of known protocols. Some basic small satellites may also have entirely ad hoc protocols that are limited to use with a single mission or a set of satellites by a specific institution. These proprietary protocols are not ideal candidates for constructing honeypots as an ecosystem of one mission that is unique to a small set of missions will look out of place for other missions, preventing generalization.
The choice of ecosystem also influences the choice of GSS as an MCS usually focuses on a single protocol stack. In addition to the space protocol ecosystems above, satellite operators use well-known network protocols to connect with the mission’s ground segment services. These protocols include Telnet, SSH, VNC, FTP, and HTTP. Telnet, SSH, and VNC are used to remotely connect to the MOC workstations. HTTP is used for web interfaces that operators use to visualize mission data, and FTP is used to download TM data from the mission data repository. If an adversary were to compromise any of these services, they could use it as a stepping stone to target the satellite itself.
2.6 Space Systems’ Tactics, Techniques and Procedures (TTPs)
Tactics, Techniques, and Procedures (TTPs) describe the behavior of a malicious actor in a structured scheme to understand how they might execute an attack [56, 47]. Several frameworks have been introduced to standardize space systems’ TTPs. There are two MITRE-style frameworks, the SPARTA matrix [68] and the SPACE-SHIELD matrix [4]. These matrices list and describe space security-specific tactics and techniques such as initial access and ground segment compromise.
3 Threat Model
Following Fig. 1, we assume an adversary willing to compromise a satellite can only interact with the space segment simulation by sending TCs from the ground segment first. To gain initial access to the ground segment, an adversary needs to connect via one of exposed network protocols depicted in Sec. 2.5, namely, VNC, Telnet, TCP/IP, etc., which correspond with the operational protocols used in real satellite missions.
From there, an adversary may try to launch different commands to take full control and/or compromise the services offered by the satellite’s mission. Finally,
in this paper, the modeling of physical radio communication between the space and ground segments, which is commonly used in practice, is considered out of scope and left for future work.
Figure 2: HoneySat framework high-level architecture design.
4 HoneySat Framework Design
In this section, we explain the objectives we aim to achieve (Sec. 4.1), the design principles we follow to meet such objectives (Sec. 4.2), and the overall design of our Space  (Sec. 4.4), and Ground (Sec. 4.3) segments.
4.1 HoneySat’s Design Objectives
Our design aims to achieve the following objectives:
  1. DO-1
    Capability to Capture Rich Interaction Data. As explained in Sec. 1, the number one objective of any honeypot is to capture interaction data from which we can derive knowledge on adversaries’ TTPs. As such, our first objective for HoneySat is to have capability to capture rich interaction data.
  2. DO-2
    Provide Deception. As we discussed in Sec. 1, honeypots’ nature must remain covert to entice adversaries into interacting with it. As such, our second objective is for HoneySat’s nature to remain hidden from adversaries.
  3. DO-3
    Provide Extensibility and Customizability. A framework’s main purpose is to provide generic functionality that can be customized to meet the user’s needs, in HoneySat’s case, we must be able to support multiple SmallSats. For example, a particular SmallSat may use CSP or CCSDS ecosystems. Additionally, each SmallSat mission has a particular orbit that must be customizable. As such, our third objective is for HoneySat to provide extensibility and customizability.
4.2 HoneySat’s Design Principles
To meet our objectives, we selected the following principles:
  1. DP-1
    High-Interaction Simulation. As we described in Sec. 2.1, high-interaction honeypots give adversaries the same or almost identical interaction opportunities as a real satellite. Thus, they can log many of the TTPs discussed in 2.6. For these reasons, we selected high-interaction simulation as our first design principle. This design principle is based on design objective DO-1.
  2. DP-2
    Realistic Simulation. As we discussed in Sec. 2.2, satellites keep track of their orbit location, the position of the sun, among others. These details must be simulated by our honeypot; otherwise, they may alert adversaries that they are interacting with a fake system. This design principle is based on design objective DO-2.
  3. DP-3
    Modularity. As we explained in Sec. 2, satellites are complex and diverse systems. There is no one-size-fits-all solution. To tackle this problem, we designed HoneySat to be modular. Modularity allows us to insert, remove, or change any part of our honeypot framework without implementing a new honeypot from scratch. This design principle is based on our objective DO-3.
We now describe HoneySat’s architecture design and how this architecture integrates the design principles described above. At the highest level, our framework’s architecture, depicted in Fig. 2, consists of two sets of simulations, the space segment simulations and the ground segment simulations. Table 2 illustrates how HoneySat’s simulation components match the components of a real satellite mission.
Segment Real Component HoneySat Component
Ground Remote Operation Protocols VNC, Telnet
Ground Mission Control Software Modified MCS
Ground Ground Station Control Software Gpredict
Ground Web Interface Web Interface
Ground Ground Station Radio Simulator
Space Flight Software F.S. Runtime + Services
Space On Board Computer Docker Container
Space Platform Satellite Simulator
Space Payload Satellite Simulator
Table 2: Correspondence between real satellite mission components and HoneySat’s design components.
4.3 Ground Segment Design
Following Sec. 2.3, the purpose of the HoneySat’s ground segment is to simulate the ground segment assets, e.g., the ground segment software. To accomplish this, our design includes the following components: 1) the Exposed Network Protocols, 2) the Ground Segment Software, 3) the Radio Simulator, 4) the Ground Personality, and 5) the Logging Repository.
1) Exposed Network Protocols. To provide adversaries with feasible access to our honeypot, HoneySat exposes multiple means of interaction over a network such as the Internet. We designed HoneySat to support four different interaction methods using different protocols, namely VNC, Telnet, Web, and Access to the ground station via raw packet transmitting capability (following design principle DP-3).
2) Ground Software. As discussed in Sec. 2.3, the ground software includes mission control software (MCS) and ground station control software (GSCS). We leverage existing ground software, allowing HoneySat to provide a high-interaction simulation (DP-1).
3) Radio Simulator. As discussed in Sec. 2.3, communication between the satellite and the ground station is possible only during a pass. The purpose of the Radio Simulator is to control communication between the ground and space segment simulations to simulate real orbital passes. The Radio Simulator design follows design principles DP-2 and DP-3.
4) Ground Configuration. The ground configuration is a series of settings for the various ground segment-specific configurations, e.g., the satellite mission logo in the Web Interface.
5) Logging Repository. This repository logs data such as the TM/TC traffic to and from the ground station software and the logging attempts received in the web interface. We designed the ground segment simulation logs to be categorized and timestamped. The logging design follows design principle DP-3.
4.4 Space Segment Design
The purpose of the HoneySat’s space segment is to mimic the spacecraft, as discussed in Sec. 2.3. To accomplish this, our design includes the following components: 1) the satellite’s Flight Software Runtime, 2) the Satellite Simulator, 3) the Flight Software Services, 4) the Satellite Personality, and 5) the Logging Repository.
1) Flight Software Runtime. As discussed in Sec. 2.4, the satellite flight software manages all critical functions required for the mission operation, such as interacting with hardware peripherals, processing TCs, and sending TM. For this to work we need an environment where services that handle TC intended to run on flight software can be run. We reuse existing compatibility or testing wrappers to run the relevant parts of flight software for HoneySat to provide nearly identical interactions to an adversary. This produces a high-interaction honeypot simulation that follows design principle DP-1.
2) Satellite Simulator. One of the biggest challenges when designing HoneySat was simulating all the physical processes, e.g., attitude, that satellites need to know. The Satellite Simulator is designed to solve this problem. It simulates all the necessary satellite subsystems, e.g., the Electrical Power System, and the sensors, e.g., a GPS receiver, that a satellite requires to function. The purpose of the Satellite Simulator is to abstract away the various hardware peripherals that communicate with the flight software by simulating these peripheral’s data. For example, suppose a particular flight software was implemented to utilize a specific lithium-ion battery pack. In that case, the Satellite Simulator’s power system simulation simulation can simulate the basic behavior of the battery. We designed the Satellite Simulator to be modular so that new subsystems and sensors can be easily added, modified, or disabled, following our DP-3.
3) Flight Software Services. The flight software services that process TC and provide TM act as the bridge between the Satellite Simulator and the rest of the Honeypot. In case a service requires some data from a subsystem of the satellite or a command is sent to a subsystem, it is passed to the satellite simulator instead. The satellite state module routes the service’s messages to and from the Satellite Simulator. It uses a list of message IDs that can be customized based on the protocol ecosystem and software architecture.
4) Satellite Personality. The satellite personality module is designed to provide a central configuration location for the space segment. It includes FS and Satellite Simulator configuration values such as the satellite’s battery capacity. The satellite personality allows our framework to be easily customized.
5) Logging Repository. We designed each space segment component to provide detailed logs. As shown in Fig. 5, all Satellite Simulator modules can send their own logs.
4.5 Theory of Operation
Figure 3: HoneySat’s Theory of Operation.
In this section, we provide a brief overview of how the designed framework components shown in Figure 2 interact by following the numbers in Fig. 3.
An adversary gets initial access via one of the Exposed Network Protocols ①. On interaction with the Simulated Ground Segment, an attacker is presented with access to the Ground Software ②. The configuration of this ground software is defined by the Ground Configuration, which allows the ground software to look like one of many different missions from its protocol ecosystem to the attacker. The attacker can then interact with the Ground Software, while their actions are reported to the Log Database. They might want to try to gain more privileges on the ground segment or try to send TC to the space segment. Instead of deploying a ground station with an RF transmitter and associated hardware, we employ the Radio Simulator ③ to handle all simulated radio frequency communications.
When an attacker uses the ground segment to send a valid TC or raw packets, the Radio Simulator checks whether the satellite configured in the Satellite Personality is currently passing over the designated ground station location. If so, the Radio Simulator forwards the TC to the Flight Software Runtime ④. When the Flight Software Runtime receives a command, it will run the corresponding Flight Software Service to compute a response ⑤. In case it requires either changes to an on-board system’s state or a sensor value to execute the TC, it will invoke the Satellite Simulator for this ⑥. The Satellite Personality contains the configuration used by the Satellite Simulator. Once a TC is processed, the resulting TM is routed back through the Radio Simulator to either the attacker or the ground software used by the attacker ⑦-⑫, and is also recorded in our Log Database.
5 HoneySat’s Implementation
Having laid out HoneySat’s design we now explain how we implemented the Satellite Simulator (Sec. 5.1), the generic CSP mission honeypot (Sec. 5.2), the ground segment (Sec. 5.2.1), and the space segment (Sec. 5.2.2).
5.1 Satellite Simulator Implementation
We implemented the Satellite Simulator as a Python object-oriented programming application.
We implemented 11 sensors, 4 subsystems, 1 satellite state, and 1 interface, based on our 6 simulations. These modules are generic enough so that we do not have to heavily modified them to support other satellites. Overall, satellite architectures, subsystems, and sensors are varied, so modularity is necessary when constructing a simulation framework. Many of the basic simulations and corresponding sensors can be shared, especially in a honeypot where attackers lack detailed information that would enable them to distinguish between small differences in sensor values. Further details on the simulations can be found in A.2
5.2 Generic CSP Mission Honeypot
Following our design we implemented HoneySat to support the CSP ecosystem. This implementation provides a generic CSP-based mission honeypot that can be customized to simulate a specific CPS-based satellite missions.
5.2.1 Ground Segment Implementation
We now describe how we implemented the five ground segment simulations designed in Sec. 4.3. 1) the Exposed Entry Points, 2) the Ground Software, 3) the Radio Proxy, 4) the Ground Configuration, and 5) Logging Repository.
1) Exposed Network Protocols. We implemented four exposed entry points. A VNC server, a Telnet server, a Web server, and ZeroMQ-based access to the CSP Network. We implemented the VNC server using TigerVNC [51] and used PyZMQ [25] for the Raw Packet Access. The Telnet server was implemented using Python to expose the command-line interface of GSS on the network. The web interface allows for the customization of text presented to the attacker through configuration. Using the above entry points, adversaries can interact with the GSS and the space segment simulations. Among these, the web interface serves as a measure to attract attackers and serves as an initial entry point for attackers. We developed a simple yet convincing web application, presented as an early version of a web interface currently under development. After the login, the attacker would be welcomed by a dashboard that displays information about the satellite and the latest output of the Ground software command line interface. Additionally, the web application provides access to documentation and the VNC server (Fig. 4), facilitating the attacker in finding the other entry points.
Figure 4: The view an attacker would see upon connection to the VNC server. Ground Station Control Software (Red) Ecosystem/Mission Specific MCS (Yellow).
2) Ground Software. To implement the MCS software for our CSP-based honeypot, we require an open-source MCS for the CSP ecosystem. We leveraged the SUCHAI mission control software (SUCHAI MCS). This software is a node on the CSP network that has a command line-based interface that allows basic MCS functionality like scheduling telecommands or saving downlinked telemetry. This is a good choice for the generic CSP MCS as it can be passed off as the CLI of commercial ground segment appliances like the GomSpace NanoCom MS100 [23]. In this generic honeypot, we can set mission-specific parameters through the ground configuration. Additionally, we provide a desktop environment accessible via VNC, complete with supporting applications such as configured Gpredict for the satellite mission. For certain deployments, this environment also includes a mission-specific GUI tool, both of which appear as open windows on the desktop.
3) Radio Simulator. The Radio Simulator was implemented as a ZMQ Hub. The ZMQ Hub functions as a router for the CSP network. Simulating both the ground part of the CSP network including the ground station and space segment. Depending on whether the satellite is currently passing, packets to the space segment are forwarded to the Flight Software Runtime. We decided to use ZeroMQ as it is the protocol to use for transmitting CSP over TCP/IP [64].
4) Ground Configuration. We implemented the ground configuration on each component, and the settings can be modified through a Docker compose file explained in Appendix A.3.
5) Logging Repository. This implementation uses the MongoDB database [43] to aggregate and store logs. The web interface records login attempts, while the Telnet server logs inputs on a per-connection basis. Additionally, network traffic to and from the host running HoneySat is captured.
5.2.2 Space Segment Implementation
1) Flight Software Runtime Although most of the flight software frameworks presented in Section 2.4 are open-source projects, it is difficult to find a real satellite flight software implementation that is available for study and modification [81]. For this reason, we use the University of Chile’s SUCHAI Flight Software (SUCHAI FS) [24], which has been deployed in four satellite missions [21]. The SUCHAI FS provides Flight software based on the CSP ecosystem that already features a mode where it can be run on a POSIX OS.
2) Flight Software Services. The SUCHAI FS uses libCSP [63] and thus features the libCSP default services that present a generic attack surface. Furthermore, there is a set of commands commonly used in satellites as a basic set of tasks to control common satellite operations such as communications. There are 96 of these commands and parameters that are parsed from CSP packets using a text-based protocol.
3) Satellite Personality. The satellite personality was implemented as a Python class that holds multiple variables related to the space segment. The current implementation includes almost 50 configurable variables, such as the satellite’s name, the battery’s nominal cell capacity, etc.
4) Logging Repository. Like the ground segment’s logging repository (Sec 5.2.1), The Satellite Simulator and the FS are connected to the MongoDB instance using a MongoDB client.
Each logged data entry is created with customized tagging and a timestamp. The collected data is structured JSON [32] format and stored in the MongoDB database.
Finally, we containerized and hardened this version of the honeypot implementation for secure and easy deployment, which we describe in Appendix A.3.
6 Evaluation
Having described HoneySat’s implementation in Sec. 5, we now evaluate it. This section lists three experimental questions designed to test HoneySat’s alignment with our design objectives. Next, we present four sets of experiments that provide empirical evidence confirming that our design objectives have been achieved within HoneySat. We describe each experiment’s environment, methodologies, and results.
6.1 Experimental Questions
These questions aim to determine whether or not HoneySat’s meets the design objectives outlined in Sec. 4.1.
  1. Q-1
    Can HoneySat offer extensive interaction opportunities to adversaries?
    Since capturing data on varied techniques is the purpose of any honeypot, we explore the capabilities of HoneySat, as described in Sec. 4. This question is related to design objective DO-1 and is addressed in Sec. 6.2 and 6.4.
  2. Q-2
    Can HoneySat simulate a SmallSat mission well enough to deceive adversaries?
    After enticing an adversary HoneySat must keep its true nature hidden, we need to simulate the satellite’s communication and physics characteristics described in Sec. 2.4. This question relates to design objective DO-2 and is answered in Sec. 6.3.
  3. Q-3
    Can HoneySat simulate different SmallSat missions?
    HoneySat’s customization is important because it would allow users of our framework to implement their own honeypots. This question relates to design objective DO-3, and we answer it in Sec. 6.5.
To answer the above research questions we conducted four experiments. First, in Sec. 6.2 we craft multiple attacks in a controlled environment to quantify the level of interaction that HoneySat provides. Second, in Sec. 6.3 we conduct a survey with experienced satellite operators to evaluate HoneySat’s realism. Third, in Sec. 6.4 we deploy HoneySat and expose it to the Internet to test its deception capabilities. Fourth and final, in Sec. 6.5 we add an additional protocol ecosystem to HoneySat to test its extensibility potential.
6.2 TTP Interaction Experiment
This experiment seeks to answer Q-1 by quantifying the different interactions that HoneySat provides. To achieve this, we leveraged the SPACE-SHIELD matrix (Version 2.0) [4] discussed in Sec. 2.6. SPACE-SHIELD provides a collection of adversary tactics and techniques for space systems. In this experiment, we determined the number of tactics and techniques that HoneySat supports. SPACE-SHIELD consists of 14 tactics and 62 techniques. However, not all of them are applicable to a virtual, network-based honeypot such as HoneySat. For example, the technique Compromise Hardware Supply Chain involves “replacing a hardware component in the supply chain with a custom or counterfeit part” which is out of the scope of a virtual honeypot. Taking this into consideration, 14 tactics, and 33 techniques are applicable to HoneySat.
Experiment Description. Our experimental environment included two hosts. One host running HoneySat and the adversary host. The HoneySat host ran Ubuntu 23.10 and was configured with the SUCHAI-2 satellite and ground personalities. The adversary host ran a Telnet client. Both hosts were connected to the same network.
Experiment Methodology. We designed one interaction or exploit for each of the applicable 33 techniques for our honeypot. The interactions involved ground segment simulations such as the web interface. The exploits were simple, using one TC, or complex with multiple TCs involved.
Experiment Results. We crafted 16 exploits and 17 interactions following the above methodology. For example, the technique T1007 - System Service Discovery’s description reads “Adversaries may try to gather information about registered local system services. Adversaries may obtain information about services using tools and OS utility commands. Adversaries may use the information from System Service Discovery during automated discovery to shape follow-on behaviors, including whether or not the adversary fully infects the target and/or attempts specific actions.”
Based on this description, we designed the following exploit to capture information about the satellite’s running processes:
11: obc_system ps -aux > ps.log
21: tm_send_file 10 ps.log
Similarly, technique T2014 - Backdoor Installation reads “An attacker can interfere with the hardware or the software, integrating or modifying the existing software, hardware configuration, or the transponder configuration to permit himself future access to the resource”. Therefore, we designed the following exploit to upload a malicious script to the satellite software:
1tm_send_file backdoor.sh 1
21: obc_system ./recv_files/backdoor.sh
Due to space limitations, we do not describe all the exploits and interactions here. However, the complete exploit and interaction list is available in Table 6 in Appendix A. The overall results for this experiment are depicted in Table 3. The key findings of our experiment are shown below, providing strong evidence for answering question Q-1 in the affirmative and design objective DO-1 as achieved.
Key findings Q-1 HoneySat supports 86.8% of the SPACE-SHIELD matrix techniques possible in a virtual satellite honeypot. HoneySat supports 100% of the SPACE-SHIELD matrix tactics.
Tactics SPACE-SHIELD Techniques (Applicable to Virtual Honeypots) HoneySat Supported Techniques
Reconnaissance 2 2
Resource Development 2 2
Initial Access 2 2
Execution 2 2
Persistence 2 2
Privilege Escalation 2 1
Defense Evasion 4 4
Credential Access 3 3
Discovery 2 2
Lateral Movement 4 1
Collection 2 2
Command & Control 2 2
Exfiltration 2 1
Impact 7 7
Total 38 33
Table 3: Tactics and techniques supported by HoneySat compared to the ones included in the SPACE-SHIELD matrix that are applicable to virtual honeypots.
6.3 SmallSat Operators Survey
Evaluating the realism of a satellite honeypot is challenging for two main reasons. First, as we discussed in Sec. 1, satellites, including SmallSats are very diverse. Second, there is no established metric or tool, such as Nmap’s OS detection [49], to quantify the level of realism of our honeypot.
To evaluate the realism and deception capabilities of HoneySat, we surveyed experienced, real-world SmallSat operators. Because operators interact with real-world SmallSat missions on a daily basis they are experts and thus are the best population to rigorously evaluate HoneySat.
Satellite Operation Evaluated Component No. Qts.
Telecommand Scheduling Mission Control SW 6
Pass Prediction Ground Station Control SW 6
Antenna Positioning Ground Station Control SW 6
EPS Telemetry Satellite Simulator 6
Temperature Telemetry Satellite Simulator 6
ADCS Telemetry Satellite Simulator 6
RGB Camera Telecommand Satellite Simulator 6
Ping Test Radio Simulator 6
Table 4: Summary of questions in Section 6.3, satellite honeypot operation tasks of the survey (48 total).
6.3.1 Survey Structure
We divided our survey into five sections. The first section collected essential background information about the participants, while the second part focused on their professional experience. In this second section, participants were asked whether they worked in industry, government, or academia and to self-report their cybersecurity skills as well as satellite-related skills (e.g., satellite hardware engineering). The third section of the survey probed participants’ satellite operation experience, including how many missions they had operated and which tools they used.
The largest section of the survey included 48 questions about satellite honeypot operation tasks. Participants watched eight 1–2 minute recordings of HoneySat ’s VNC view (Fig. 4), each showcasing the ground and satellite personality of a real CSP-based CubeSat mission (based on the SUCHAI-2 CubeSat [34]). Every recording highlighted a different feature of HoneySat by replicating a real-world satellite mission operation, as discussed in Section 2. After each recording, participants answered questions designed to assess how realistic they found that particular aspect of the honeypot. Table 4 lists all of the satellite operations demonstrated in these recordings. Finally, once participants had seen multiple elements of our honeypot, the survey concluded with an overall evaluation section. Both the honeypot operation tasks and the overall evaluation used 5-point Likert scale [55] questions to measure participants’ positive or negative reactions.
6.3.2 Participants
We conducted the survey via Qualtrics and distributed the survey directly to active SmallSat mission operators from previously identified missions. In total, we received responses from 14 satellite operators who have between 1 and 5 years of experience operating satellites and have operated between 2 and 7 unique missions. In terms of demographics, 21.4% (3/14) of participants were female, 71.4% (10/14) male and 7.1% (1/14) preferred not to answer. 57.1% (8/14) belonged to the 18-24 age group, 35.7% (5/14) to 25-34 and 7.1% (1/14) to 35-44. 28.5% (4/14) of participants’ highest level of education was high school, 35.7% (5/14) bachelor’s degree, 14.2% (2/14) master’s degree and 21.4% (3/14) Ph.D. In regards to geographic location, 78.5% (11/14) of participants were located in Europe, 14.2% (2/14) in North America and 7.1% (1/14) in South America.
Recruiting participants was challenging due to the rarity and specialized nature of the required expertise. Over a span of four months, we reached out to national and international institutions, as well as private corporations involved in operational satellite missions to identify suitable participants.
6.3.3 Methodology and Key Results
In the survey, we aimed to evaluate three key aspects of our honeypot. First, whether the ground segment simulations are realistic; second, whether the space segment simulations are realistic; and third, whether HoneySat, as a whole, provides a convincing and realistic system.
Before describing the results it is important to emphasize that the participants were informed that the recordings showed a simulation of a satellite mission and not of a real mission.
Ground Segment Realism. To understand if HoneySat’s ground segment is realistic we showed participants recordings of different satellite operations (discussed in Sec. 2.3) performed with HoneySat, to showcase the ground segment components listed in Table 4.
For example, one of the recordings shows the pass prediction operation which involves calculating when and where a satellite will be visible or within range of a specific ground station. This operation is performed using the ground station control software. After the participants watched a recording of this operation, we asked them to rate their perceived level of realism. 85.7% of participants agreed or strongly agreed that the pass prediction operation done in HoneySat resembles that of a real mission.
Another relevant operation is the telecommand scheduling operation which involves the planning and organization of the commands to be sent to the satellite during a pass. 42.8% of participants strongly agreed that the telecommand scheduling operation performed in HoneySat resembles that of a real mission mentioning the use of a terminal-based mission control software. Interestingly, 35.7% of participants neither agreed nor disagreed, citing that they do not use a command line tool but instead a GUI. These results indicate that some of the surveyed operators use a command line-based mission control software while others use a GUI-based confirming the diversity of means of operation in Sec. 2.3.
In summary, according to our survey, most participants perceive the ground segment simulated by HoneySat as comparable to that of a real satellite mission. While a few participants mentioned that they expected to see a GUI-based mission control software instead of a command line-based one, this is something that we expected, as the means of satellite operations vary greatly among missions. Nevertheless, thanks to HoneySat’s modularity HoneySat can be extended to use a GUI-based mission control software such as YAMCS [1].
Space Segment Realism. In a similar manner, to understand if HoneySat’s space segment is realistic, we showed participants recordings of different satellite operations that make use of HoneySat’s space segment simulations (discussed in Sec. 4.4) to showcase multiple components of the space segment listed in Table 4. For example, one of the recordings shows the EPS Telemetry operation which involves the collection of data from the Electrical Power Subsystem discussed in Sec. 2.4. This operation is performed using the mission control software to send telecommands and download the latest EPS telemetry. After the participants watched a recording of this operation performed by HoneySat, we asked them to rate their perceived level of realism of the telemetry output shown during the operation. 57.1% of participants agreed or strongly agreed that the telemetry shown in the EPS Telemetry operation resembles that of a real mission mentioning realistic EPS values such as voltage, and temperature. 14.2% neither agreed nor disagreed mentioning that the EPS temeletry was presented in a different format. Finally, 28.5% disagreed or strongly disagreed mentioning that the telemetry included less values than the mission that they had operated. These results indicate that the EPS telemetry generate by the Satellite Simulator is considered realistic by the majority of participants.
We also asked participants to asses the ping test operation. This operation is a diagnostic procedure used to verify the communication link between a ground station and the satellite. It involves sending CSP packets between two CSP nodes, namely the ground segment and the spacecraft. After the participants watched a recording of this operation performed on HoneySat, we asked them to rate their perceived level of realism. 64.2% of participants agreed or strongly agreed that the ping test operation resembles that of a real mission citing the realistic response times. 21.4% of participants neither agreed nor disagreed stating that in their experience the ping test also downloads additional telemetry. Finally, 14.2% disagreed or strongly disagreed noting that in their mission the ping command has a different name. The results of the ping test operation indicate that HoneySat’s radio simulator (which controls the communication between the ground and space segments) is considered to realistically simulate the communication of a real satellite mission by the majority of participants. A significant key finding of our survey is shown below. Overall, the results obtained provide convincing evidence for answering question Q-2 in the affirmative and design objective DO-2 as achieved.
Key finding Q-2 71.4% of surveyed satellite operators agreed or strongly agreed with the following statement: “If I did not know this was a honeypot simulation of a CubeSat satellite mission, I would have believed it to be an actual CubeSat satellite mission.”
6.4 Internet Interaction Experiment
This experiment explores HoneySat’s capabilities to entice external actors in the wild by deploying our honeypot over the Internet. To accomplish this, we leveraged the Shodan search engine. Shodan is a search engine for Internet-connected devices that scans the whole Internet and indexes exposed servers and their TCP ports. Shodan then reads the banner information for each port. For example, it gathers the web headers and Telnet login banners [40]. Additionally, Shodan tags servers as honeypots if the servers have too many open ports [37].
Experiment Description. We deployed and configured five HoneySat instances over the Internet and exposed TCP port 24 for the Telnet server and 80 for the web interface so that anybody can interact with it and allow Shodan to index our server and its banners. We deployed four HoneySat instances in total.
Honeypot
Source
IP
Cmds
Received
Engaged
Time
Honeypot 1 Egypt 4 2 hr
Honeypot 2 Sweden (Tor) 9 5 min
Honeypot 1 France (Tor) 6 4 min
Honeypot 3 USA 8 3 min
Table 5: Exposed Telnet interactions received.
Experiment Results. Our honeypot servers were successfully indexed by Shodan. Both Telnet and web banners were captured, and none of them were tagged as honeypots.
Our honeypots successfully enticed external actors and captured 4 distinct sessions (show in Table 5) via the exposed Telnet protocol discussed in Sec. 4.3. The exposed Telnet protocol was implemented so that a human had to type “activate” in the terminal before sending commands. This feature was implemented to filter out Internet Telnet bots. As such, the commands that we describe below were sent by a human and not a crawler or Internet bot.
We now describe the commands received in the first interaction session shown in Table 5.
  1. 1.
    help: Shows the ground software available commands.
  2. 2.
    ls: The Linux ls command is not a valid command in the honeypot ground software.
  3. 3.
    fp_show: Prints the list of commands in the current flight plan (The flight plan refers to a pre-programmed sequence of tasks, and commands that guide the satellite’s operation).
  4. 4.
    com_debug: Prints the current CSP network configuration including the current CSP node, interfaces and routing table.
The commands fp_show and com_debug are flight software-specific commands which suggests that the adversary was successfully deceived by our honeypot and purposefully sent flight software commands instead of haphazardly sending unrelated commands. Another factor that indicates that the adversary was deceived is that they engaged with our honeypot over a period of two hours indicating that they did not immediately identified it as a honeypot.
In summary, these results provide strong evidence for answering question Q-2 in the affirmative and design objective DO-2 as achieved.
Key finding Q-2 Human adversaries interacted with three of our HoneySat-powered honeypots during four distinct Telnet sessions resulting in 16 satellite mission-specific commands captured.
6.5 Case Study: Generic CCSDS Mission Honeypot
In Sec. 5.2 we described how we implemented HoneySat using one SmallSat protocol ecosystem, namely, CSP. In this case study we are interested in testing the extensibility capabilities of HoneySat to support additional ecosystems (discussed in Sec. 2.5) by adding a second ecosystem to HoneySat, namely the CCSDS ecosystem.
Experiment Description. We selected the CCSDS ecosystem because it is a standard protocol suite used by other SmallSats [81]. To accomplish this, we leveraged an open-source CCSDS ecosystem-based flight software framework, RACCOON OS [31], and YAMCS, an open-source Mission Control software framework with built-in support for PUS [1].
Methodology. Building upon our HoneySat framework implementation, as detailed in Sec. 5, we enhanced the system by integrating various components of the RACCOON OS and the YAMCS framework.
Results. We successfully extended HoneySat to support the CCSDS ecosystem. Regarding the ground segment we implemented the exposed network protocol using YAMCS’ built-in web interface (Fig. 7), for the mission control software we used YAMCS’s built-in Mission Control Software [1], for the radio simulator we used RACCOON’s communication application and for the ground configuration and logging repository we again used YAMCS built-in features.
Focusing on the space segment, we implemented the satellite flight software runtime using the RACCOON framework, and for the Flight Software Services, we configured the RACCOON framework to connect it to the YAMCS’s MCS on the ground segment. The satellite personality and logging repository were based on the existing HoneySat implementations.
The majority of the effort involved in extending HoneySat to support the CCSDS ecosystem was dedicated to understanding the ecosystem itself, including components such as PUS. Additionally, we had to analyze the RACCOON flight software code and the YAMCS framework documentation. The only extra implementation that was required was the Flight Software Services which we modified using Rust. Other than that we reused several elements of HoneySat like the satellite personality and the Satellite Simulator.
In summary, these results provide strong evidence for answering question Q-3 in the affirmative and design objective DO-3 as achieved.
Key findings Q-3 Out of the box HoneySat supports CSP and CCSDS, the two most widely used standard space ecosystems, and was evaluated by simulating 3 real-world SmallSats.
7 Discussion and Future Work
Limitations. Currently, the functionality of some subsystems within HoneySat’s Satellite Simulator are constrained by the quality of the data provided. For example, the resolution of Earth’s images generated by our camera payload depends on the resolution of its source data (USGS [76]). Consequently, creating a honeypot for a satellite with a high-resolution camera payload would not be feasible without addressing the underlying issues of data source quality.
Broader Applications of HoneySat. While originally designed as a honeypot, our framework offers the flexibility to support a range of applications beyond its initial purpose. One promising use case is the development of digital twins for satellite systems, enabling the simulation of real-world satellite subsystems and communication scenarios. Furthermore, HoneySat can be integrated into cyber range environments to enhance training programs and cybersecurity simulation exercises. Lastly, HoneySat serves as an educational tool, providing researchers, industry professionals, and security enthusiasts with an opportunity to explore and learn about satellite architecture and operations.
8 Conclusion
Although we have yet to witness a Stuxnet-like cyberattack on space systems, security researchers and professionals need to develop effective countermeasures to secure satellites. In this paper, we introduced HoneySat, the first satellite honeypot that provides a much-needed alternative source of empirical data on attackers’ TTPs. We created a satellite honeypot that realistically simulates an entire satellite system, including its sensors and subsystems.
We performed several experiments that provide strong evidence that the framework can obtain rich interaction data, can be extended and customized, and simulate a satellite to keep its honeypot nature hidden from potential adversaries. Finally, we hope that security researchers and professionals take advantage of HoneySat’s open-source implementation and use it as a foundation not only for satellite honeypot deployments but also for simulation, education, and training applications.
9 Ethical Considerations
In this paper we consider the ethical consequences and possible negative outcomes of two main elements of our research. The satellite operator user study and the multiple deployments of our satellite honeypot exposed to the Internet. We now discuss the stakeholders, potential risks and how we mitigated those risks for each of these elements.
9.1 Satellite Operators User Study
Our user study adheres to the ethical guidelines outlined in the USENIX Security ‘25 Ethics Policy. Prior to conducting any research, the study plan, survey protocol, and associated materials were submitted for review by our Institutional Review Board (IRB), which evaluated the study’s potential risks to participants. The protocol received "exempt status" from the IRB, indicating that the research involves no more than minimal risk. This status reflects our careful evaluation of possible harms, which we discuss further below.
Stakeholder Identification. The primary stakeholders in the survey are survey participants. Our study aims to gather voluntary responses regarding the fidelity and realism of our honeypot system. We carefully considered the potential risks to both the participants of the survey.
Risk Mitigation.The main risk for the survey participants is the breach of privacy. To mitigate this risk, we ensured that no identifiable data (e.g., names, contact details) was collected from survey participants and the survey responses remain fully anonymous. Additionally, the survey itself includes an informed consent section that informs the respondents about the voluntary participation, the survey’s purpose, a description of the procedures, the risks involved, and contact information and the option to opt out of the survey. Finally, we provided participants the choice to skip questions using the response options “Prefer not to say” and “I do not know” to applicable questions. Due to the aforementioned risk mitigation strategies the stakeholders’ breach of privacy risk is negligible.
9.2 Honeypot Deployment Experiment
Stakeholder Identification. The primary stakeholders in the honeypot deployment are the external actors who interact with our honeypot and the infrastructure owners on whose infrastructure our honeypot is running. In the honeypot deployment experiment we deployed HoneySat on server infrastructure owned by real-world institutions.
Risk Mitigation. The main risk for the infrastructure owners is that one of the external actors will use our honeypot as an stepping stone to breach their production infrastructure. For the deployment on the premises of the institutions we worked with the local IT infrastructure administrators to ensure that we took all the necessary precautions to avoid any possibility that the external actors could gain access to the infrastructure via our honeypot. First, the infrastructure owners provided us with a server with a clean OS installation that did not have any production applications or sensitive data. Second, the server was isolated on its own network segment. Third, we configured a firewall to only allow the necessary ports. Fourth, the VNC portion of the honeypot is configured so that external actors cannot use it to host malicious services by denying traffic forwarding. Fifth and final, we deployed our honeypot using two sandboxing layers (Docker containers and Virtual Machines (VMs) as we discussed in Appendix A.3. Due to the aforementioned risk mitigation strategies the stakeholders’ breach of privacy risk is negligible.
The main risk for the external actors is the breach of their privacy. However, it is generally accepted that trespassers of a computer system do not have reasonable expectation of privacy [66]. In order to mitigate this risk our honeypot system gives notice and warning to any user connecting to it indicating that "This computer system is for authorized use only." Additionally, the web services of our honeypot are protected by a username and password. This notice system gives external actors a chance to avoid interaction, which helps reduce the risk.
Appendix A Appendix
A.1 Interaction Sequences and Exploits
Table 6 includes all the interaction sequences and exploits we performed during the experiments in Sec. 6.2.
Table 6: Tactics, Techniques and Procedures’ Interaction Experimental Exploits.
Tactic Technique ID Subsystem Exploit
Reconnaissance Active Scanning (RF/Optical) T2001 Threat Model Limitation N.A.
Reconnaissance Gather Victim Mission Information T2002 Web Interface Use the Web Interface to gather mission documentation.
Reconnaissance Gather Victim Org Information T1591 Web interface Use the Web Interface to gather mission documentation.
Reconnaissance In orbit proximity intelligence T2029 Threat Model Limitation N.A.
Reconnaissance Passive Interception (RF/Optical) T2004 Threat Model Limitation N.A.
Reconnaissance Phishing for Information T1598 Tangential N.A.
Resource Development Acquire or Build Infrastructure T1583 Telnet interface Acquire ground segment using the Telnet service.
Resource Development Compromise Account T2038 Tangential N.A.
Resource Development Compromise Infrastructure T1584 Threat Model Limitation N.A.
Resource Development Develop/Obtain Capabilities T2007
Ground software.
Flight software.
Exploit OS, libraries or software vulnarabilities.
Deploy custom CSP application to forge TC/TM.
Initial Access Direct Attack to Space Communication Links T2008
Ground software.
Flight software.
Use the ground software to send/receive TC/TM.
Deploy custom CSP application to forge TC/TM.
Initial Access Ground Segment Compromise T2030
Telnet interface.
Ground software.
Use the telnet interface to access the ground software.
Initial Access Supply Chain Compromise T1195 Tangential N.A.
Initial Access Trusted Relationship T2039 Threat Model Limitation N.A.
Initial Access Valid Credentials T2009 Tangential N.A.
Execution Modification of On Board Control Procedures modification T2010
Ground software.
Flight software.
Upload a script to the satellite software:
tm_send_file code.py 1
1: obc_system python recv_files/code.py
Execution Native API T1106
Ground software.
Flight software.
Execute shell commands or delete system files:
1: obc_system <command>
1: obc_rm -r \$HOME
Execution Payload Exploitation to Execute Commands T2012 Tangential N.A.
Persistence Backdoor Installation T2014
Ground software.
Flight software.
Upload a script to the satellite software:
tm_send_file backdoor.sh 1
1: obc_system ./recv_files/backdoor.sh
Persistence Key Management Infrastructure Manipulation T2013 Tangential N.A.
Persistence Pre-OS Boot T1542
Ground software.
Flight software.
Use obc_system, obc_rm, obc_mkdir commands.
Upload/modify an OS configuration file.
Start/stop/schedule execution of services/daemons.
Persistence Valid Credentials T2009 Tangential N.A.
Privilege Escalation Become Avionics Bus Master T2031 Implementation limitation N.A.
Privilege Escalation Escape to Host T1611
Docker.
Virtual machine.
Escape the container/VM with previously crafted exploits.
Defense Evasion Impair Defenses T1562
Ground software.
Flight software.
Send commands to change operation mode.
1: drp_set_var_name obc_opmode 0
Defense Evasion Indicator Removal on Host T1070
Ground software.
Flight software.
Use commands to remove artifacts, logs, etc.:
1: obc_rm <path>
Defense Evasion Masquerading T2040
Ground software.
Flight software.
Use commands to upload artifacts modify system settings:
tm_send_file articaft 1
1: obc_mv artifact /etc/config/artifact
Defense Evasion Pre-OS Boot T2041
Ground software.
Flight software.
Use obc_system, obc_rm, obc_mkdir commands.
Upload/modify an OS configuration file.
Start/stop/schedule execution of services/daemons.
Credential Access Adversary in the Middle T2042
Ground software.
Flight software.
Deploy a CSP application to capture/inject CSP packets.
Credential Access Brute Force T2043
Ground software.
Flight software.
Brute force valid TC parameters:
1: obc_ebf <KEY>
Credential Access Communication Link Sniffing T2044 Ground software.
Scape the ground software or docker and run tcpdump.
Deploy a CSP application to capture/inject CSP packets.
Credential Access Retrieve TT&C master/session keys T2015 Tangential N.A.
Discovery Key Management Policy Discovery T2032 Tangential N.A.
Discovery Spacecraft’s Components Discovery T2034
Ground software.
Flight software.
Send TC to redirect satellite logs to ground segment:
1: log_set 5 2 10
Discovery System Service Discovery T1007
Ground software.
Flight software.
Capture running processes information
1: obc_system ps -aux >ps.log
1: tm_send_file 10 ps.log
Discovery Trust Relationships Discovery T2033 Tangential N.A.
Lateral Movement Compromise a Payload after compromising the main satellite platform T2045 Implementation Limitation N.A.
Lateral Movement
Compromise of satellite hypervisors
T2017
Docker.
Virtual machine.
Escape the container/VM with previously crafted exploits.
Lateral Movement Compromise the satellite platform starting from a compromised payload. T2046 Implementation Limitation N.A.
Lateral Movement Lateral Movement via common Avionics Bus. T2016 Implementation Limitation N.A.
Collection Adversary in the Middle T1557
Ground software.
Flight software.
Deploy a CSP application to capture/inject CSP packets.
Collection Data from link eavesdropping T2018
Ground software.
Flight software.
Scape the ground software or docker and run tcpdump.
Deploy a CSP application to capture/inject CSP packets.
Command and Control Protocol Tunnelling T2047
Ground software.
Flight software.
Deploy an malicious application that sends data over CSP
Command and Control Telecommand a Spacescraft T2019
Ground software.
Flight software.
Use the ground software to send TC:
1: com_ping 1
Command and Control TT&C over ISL T2048 Threat Model Limitation N.A.
Exfiltration Exfiltration Over Payload Channel T2021 Implementation Limitation N.A.
Exfiltration Exfiltration Over TM Channel T2022
Ground software.
Flight software.
The atacker deploys a custom CSP node or a backdoor
Exfiltration Optical link modification T2037 Threat Model Limitation N.A.
Exfiltration RF modification T2036 Threat Model Limitation N.A.
Exfiltration Side-channel exfiltration T2035 Threat Model Limitation N.A.
Impact Data Manipulation T2054
Ground software.
Flight software.
Send TC to modify/reset TM database:
1: drp_set_var_name drp_ack_ads 10000000
1: drp_reset_payload 1 1010
1: drp_reset_status 1010
Impact Ground Segment Jamming T2050 Threat Model Limitation N.A.
Impact Loss of spacecraft telecommanding T2055
Ground software.
Flight software.
Send TC to change communication paramters.
Modify network configuration in the ground station.
Impact Permanent loss to telecommand satellite T2027
Ground software.
Flight software.
Send TC to destroy filesystem:
1: obc_system rm -rf –no-preserve-root /
Impact Resource damage T2028 Threat Model Limitation N.A.
Impact Resource Hijacking T1496
Ground software.
Flight software.
Upload a script to the satellite software:
tm_send_file code.py 1
1: obc_system python recv_files/code.py
Impact Saturation of Inter Satellite Links T2052 Threat Model Limitation N.A.
Impact Saturation/Exhaustion of Spacecraft Resources T2053
Ground software.
Flight software.
Send TC to create a reset loop:
1: fp_set_cmd_dt 10 2147483647 10 obc_reset
Impact Service Stop T1489
Ground software.
Flight software.
Send TC to lauch a fork bomb or a reset loop
1: obc_system :(){ :|:& };:
1: fp_set_cmd_dt 10 2147483647 10 obc_reset
Impact Spacecraft Jamming T2049 Threat Model Limitation N.A.
Impact Temporary loss to telecommand satellite T2026
Ground software.
Flight software.
Send TC to make the system unresponsive
1: obc_system sleep 3600
Impact Transmitted Data Manipulation T2024 Threat Model Limitation N.A.
A.2 Satellite Simulator Design Details
Figure 5: High-level architecture of the Satellite Simulator. Due to space limitations, we do not show all the available sensors and simulations. Some sensors communicate with multiple simulations. Although the flight software services are not part of the Satellite Simulator we included it to depict the Interface module’s data flow better.
As shown in Fig. 5, the Satellite Simulator includes five module types: simulation modules (light orange), sensor modules (light blue), subsystem state modules (light green), satellite state modules (light gray) and interface modules (light yellow).
Simulation Modules. These modules are abstractions of real-world processes required for a realistic satellite simulation (DP-2). An example of these simulations is the orbital simulation. This simulation uses orbital mechanics to calculate data such as the satellite’s orbital position at any given time.
These simulation modules can communicate with each other to facilitate proper functionality. For example, the power simulation module will query the orbital simulation for orbital data to determine if the satellite is positioned properly in relation to the sun so that it can draw power from its solar panels. The subsystems and the sensors of the Satellite Simulator are designed around these simulations. The Satellite Simulator includes the orbital, rotation, power system, thermal, magnetic, and payload simulations. Due to space restrictions and their complex design, we explain each simulation in Appendix A.2.
Overall, these simulations provide the foundation for any satellite operator to use our framework and implement their honeypot using their satellite flight software. The Satellite Simulator simulations are configured using the satellite personality, which we discuss below.
Sensor Modules. These modules are abstractions of the hardware sensors used by a satellite, i.e., temperature sensor. The sensor modules do not perform any computations, instead they collect data directly from the simulation modules following DP-2. For example, the voltage sensor will query the power simulation to collect data about the current state of the battery.
Subsystem State Modules. These modules are abstractions of the satellite subsystems discussed in Sec. 2.4. A given subsystem state is made up of one or more sensors. For example, in Fig. 5 the Payload State (light green) includes the Camera Sensor (light blue). In this way, subsystem states serve two purposes. They can get data from their sensors or can set a specific configuration on their sensors.
Satellite State Module. This module is an abstraction of an entire satellite. As depicted in Fig. 5 the Satellite State (light gray) is made up of multiple subsystem states. Satellite State module routes messages between the Satellite Simulator and the modified flight software. For example, if the modified flight software sends a message to the Satellite Simulator requesting to provide the present voltage in the EPS, the satellite state will route that message to the EPS State.
Interface Module. This module is an abstraction of the communication protocol between the modified flight software and the Satellite Simulator. As depicted in Fig. 5, the Interface (light yellow) is the module that connects the flight software services with the Satellite Simulator simulations.
This protocol must implement two basic message types, requests and replies. However, to meet DO-3, the underlying implementation of these messages is left open for the user to decide based on their requirements.
To finish, the Satellite Simulator’s built-in modular design allows subsystems, simulations, and sensor modules to be mixed ’n matched to meet the requirements of most satellites and their missions. Additionally, our flexible design allows users to easily add additional simulations, sensors, or subsystems as modules to the honeypot.
  • Orbital Simulation. This simulation calculates and predicts the satellite’s orbit. To achieve this, it uses the two-line element set or TLE data configured in the satellite personality. The TLE data contains the necessary orbital information required to calculate the location of a satellite given a point in time [77]. We rely on the Skyfield python library to perform advanced and precise calculations [57]. This simulation calculates multiple orbital values which include the altitude angle, latitude, longitude of the satellite among other values. The payload, EPS, and magnetic simulations rely on data from this simulation.
  • Rotation Simulation. This simulation simulates the satellite attitude changes to provide the satellite orientation. Determining and controlling the satellite orientation is relevant for payloads such as cameras or communication systems that require pointing to the earth surface or to other celestial bodies. This simulation calculates the rotation kinematics of the satellite using two reference frames: a reference frame fixed to the satellite body and a non-rotational reference frame. The relation between the two reference frames is calculated using the conservation of angular momentum and the rigid-body Euler equation [73]. To simulate the satellite’s rotation, the initial instantaneous angular velocity of the satellite, the current orientation in quaternions, and the satellite’s inertia matrix must be defined. The simulation uses the fourth-order Runge Kutta integration method (RK4) [9] to perform these calculations to avoid divergences in the simulation with time steps greater than 1 second.
  • Power System Simulation. This simulation manages the satellite’s power collection, consumption, and distribution. It tracks the battery capacity, the power draw and simulates battery charging whenever the orbital simulation tells it that the satellite is exposed to sunlight. A battery pack is modeled as a series or parallel configuration of smaller battery packs. The simplest battery pack consists of a single cell. A single cell is simulated to calculate the charge and discharge curves, assuming that the cell’s current will not change. This cell voltage is then used to calculate the pack’s voltage. The electrical power system simulation uses the pack’s voltage together with the input and output power to calculate the battery current. In case the input or output power of the power subsystem changes, a new cell current is calculated, and the battery cell simulation is rerun to reflect the change.
  • Thermal Simulation. This simulation calculates the temperature of the satellite. The simulation is based on two values: the total thermal energy and a user-defined specific heat capacity. Together with emissivity, mass and surface area, the temperature of a satellite can be approximated based on the absorbed sunlight and the energy emitted through thermal radiation. This simulation uses a radiation loss formula [22] to calculate the thermal radiation emission.
  • Magnetic Simulation. This simulation tracks and analyzes the interactions between the Earth’s magnetic field and the satellite’s own magnetic environment. It communicates with the orbital and rotation simulations to determine the satellite’s position and orientation in space so that it can calculate the Earth’s magnetic field components (north, east, vertical) and total intensity at the satellite’s current location.
  • RGB Camera Simulation. This simulation replicates the functionality of an RGB camera pointed at Earth. It uses Earth observation satellite imagery from the EarthExplorer database from the U.S. Geological Survey (USGS) [76]. If a capture image command is issued, it will select an image from one of two pools of images: daytime or nighttime.
    This simulation communicates with the orbital simulation to determine if satellite is in the presence of the sun.
A.3 Deployment and Security Hardening Implementation
Figure 6: Architecture of dockerized Generic CSP Honeypot
Figure 7: Screenshot of the YAMCS web interface, displaying a TM packet, returned from the Generic CCSDS Mission Honeypot.
As we mentioned in Sec. 2.1, high-interaction honeypots such as HoneySat present a high risk of adversary takeover. To mitigate this risk, we implemented HoneySat with two sandboxing layers.
Virtual Machine. VMs provide the highest isolation level among sandboxing techniques. We implemented HoneySat in VM environments to leverage VMs’ robust security.
Containerization. After we completed HoneySat’s development, we used Docker Compose [2] to containerize each of our framework’s applications. Specifically, we created four different containers depicted in Fig. 6. Containerizing HoneySat provides two benefits. First, it creates another sandboxing layer that prevents adversaries from using our honeypot to access the underlying system [39]. Second, it proves a convenient and flexible way to deploy HoneySat.

### v2

@@
 Abstract
-Satellites are the backbone of several mission-critical services, such as GPS that enable our modern society to function.
-For many years, satellites were assumed to be secure because of their indecipherable architectures and the reliance on security by obscurity.
-However, technological advancements have made these assumptions obsolete, paving the way for potential attacks, and sparking a renewed interest in satellite security.
-Unfortunately, to this day, there is no efficient way to collect data on adversarial techniques for satellites, which severely hurts the generation of security intelligence.
-In this paper, we present HoneySat, the first high-interaction satellite honeypot framework, which is fully capable of convincingly simulating a real-world CubeSat, a type of Small Satellite (SmallSat) widely used in practice.
-To provide evidence of the effectiveness of HoneySat, we surveyed experienced SmallSat operators currently in charge of active in-orbit satellite missions.
-Results revealed that the majority of satellite operators (71.4%) agreed that HoneySat provides realistic and engaging simulations of CubeSat missions.
-Further experimental evaluations also showed that HoneySat provides adversaries with extensive interaction opportunities by supporting the majority of adversarial techniques (86.8%) and tactics (100%) that target satellites.
-Additionally, we also obtained a series of real interactions from actual adversaries by deploying HoneySat on the internet over several months, confirming that HoneySat can operate covertly and efficiently while collecting highly valuable interaction data.
-1 Introduction
-Artificial satellites are complex devices designed to withstand outer space conditions.
-They serve multiple purposes or types of missions that include position and navigation, e.g., the Global Positioning System (GPS) constellation [74], Earth observation, e.g., the Sentinel constellation [19], and broadband internet service, e.g., the Starlink constellation.
-In addition, Spacecraft can range from being as massive as thousands of kilograms, such as the International Space Station (ISS) to one kilogram CubeSats.
+Satellites are the backbone of several mission-critical services that enable our modern society to function, for example, GPS.
+For years, satellites were assumed to be secure because of their indecipherable architectures and the reliance on security by obscurity.
+However, technological advancements have made these assumptions obsolete, paving the way for potential attacks, and sparking interest in satellite security.
+Unfortunately, to this day, there is no efficient way to collect data on adversarial techniques for satellites, hurting the generation of security intelligence that can lead to the development of effective countermeasures.
+In this paper, we present HoneySat, the first high-interaction satellite honeypot framework, fully capable of convincingly simulating a real-world CubeSat, a type of Small Satellite (SmallSat).
+To provide evidence of HoneySat’s effectiveness, we surveyed experienced SmallSat operators in charge of in-orbit satellites and deployed HoneySat over the Internet to entice adversaries.
+Our results show that 90% of satellite operators agreed that HoneySat provides a realistic and engaging simulation of a SmallSat mission.
+Additionally, HoneySat successfully deceived human adversaries in the wild and collected 22 real-world satellite-specific adversarial interactions.
+Finally, in a major demonstration of HoneySat’s robustness, we collaborated with an aerospace company to perform a hardware-in-the-loop operation that resulted in HoneySat successfully communicating with an in-orbit, operational SmallSat mission.
+publicationid: pubid: Network and Distributed System Security (NDSS) Symposium 2026 23-27 February 2026, San Diego, CA, USA ISBN 979-8-9894372-8-3 https://dx.doi.org/10.14722/ndss.2026.240537 www.ndss-symposium.org
+I Introduction
+Satellites are complex devices designed to withstand outer space conditions.
+They serve multiple purposes or types of missions that include position and navigation, e.g., the Global Positioning System (GPS) constellation [1], Earth observation, e.g., the Sentinel constellation [2], and broadband internet service, e.g., the Starlink constellation [3].
+In addition, spacecraft can range from being as massive as thousands of kilograms, such as the International Space Station (ISS), to much smaller one-kilogram CubeSats [4].
 As a result, the software and hardware components that make up a specific spacecraft vary greatly.
-Likewise, satellite missions’ vary a lot depending on the owner’s budget.
-For example, university missions are smaller as they have a limited budget.
+Likewise, satellite missions vary a lot depending on the owner’s budget.
+For example, university missions are smaller as they have a limited budget [5].
 A cyberattack on a satellite or satellite constellation (group of satellites) could have disastrous consequences on a global scale, which is difficult to comprehend.
-Such an attack could lead to the cessation of air traffic and widespread communication blackouts.
+Such an attack could lead to the cessation of air traffic and widespread communication blackouts [6].
 It could also cause food shortages and the freezing of financial transactions [7].
@@
 Third, satellite engineers and operators continue to rely on protocol obscurity practices, such as hiding specialized knowledge about the transmission protocol implemented on satellites [12].
-Although satellite security research has gained increased attention [81, 60, 53, 79], the relationship between outer space and cyberspace remains poorly understood [52].
-At the same time, space-focused threat intelligence remains sparse, and our current methods for identifying tactics, techniques, and procedures (TTPs) used against these critical systems are limited.
+Although satellite security research has gained increased attention [12, 13], the relationship between outer space and cyberspace remains poorly understood [14].
+At the same time, space-focused threat intelligence remains sparse, and our current methods for identifying the tactics, techniques, and procedures (TTPs) used against these critical systems are limited.
 MITRE ATT&CK currently tracks 152 threat groups but shows only one that targets satellites explicitly [15], highlighting a significant data gap.
@@
 As a result, the security community has limited visibility into adversarial activity aimed at space infrastructure.
-Honeypots’ ability to collect real-world cyberattack data makes them an ideal solution to this problem [29].
+A honeypot’s ability to collect real-world cyberattack data makes it an ideal solution to this problem [16].
 A honeypot is a decoy computer system intended to lure and entice malicious actors to interact with it [17]; all the while, the honeypot logs all the interactions the attackers make.
@@
 However, as of the time of writing, there is no space-sector specific honeypot in the literature.
-In this paper, we design, implement, and test the first satellite honeypot, HoneySat to attract and analyze adversaries who attack space infrastructures over the Internet, a commonly observed threat vector [7, 5, 33, 80].
+In this paper, we design, implement, and evaluate the first satellite honeypot, HoneySat, to attract and analyze adversaries who attack space infrastructures over the Internet, a commonly observed threat vector [25, 26, 27, 28].
 HoneySat is a modular, high-interaction honeypot framework that realistically simulates a complete satellite system (ground infrastructure and satellite).
-Specifically, HoneySat simulates Small Satellites or SmallSats which are spacecraft with a mass of less than 180 kilograms[44].
-CubeSats for example, are SmallSats.
+Specifically, HoneySat simulates Small Satellites or SmallSats, which are spacecraft with a mass of less than 180 kilograms [29].
+CubeSats, for example, are SmallSats.
 Additionally, as part of HoneySat, we developed the Satellite Simulator, a Python project to provide generic simulation functionality for users to populate satellite honeypots with believable data.
@@
 Our results, backed by our survey of satellite operators, show that HoneySat’s simulation is highly realistic.
-Our framework is able to simulate an entire real satellite mission, provides realistic telemetry, and supports real telecommands.
+Our framework simulates an entire satellite mission, provides realistic telemetry, and supports real telecommands.
+Moreover, in a significant showcase of HoneySat’s practical relevance, a leading multinational space agency has integrated HoneySat into operational test environments to evaluate its effectiveness in securing mission-critical space systems.
+Finally, we collaborated with an aerospace company to perform a hardware-in-the-loop experiment where HoneySat successfully communicated with an in-orbit, operational SmallSat.
 In summary, this paper makes the following contributions:
-  • A novel framework, HoneySat, a high-interaction, extensible honeypot for small satellites (Sec. 4).
-  • The Satellite Simulator, that simulates the physical processes, sensors, and subsystems necessary for a realistic satellite honeypot (Sec. 5).
-  • The results of a survey of experienced satellite operators that provide valuable insights into how realistically our honeypot performs, as well as experimental evidence demonstrating that the HoneySat framework can simulate small satellites, collect rich real-world interaction data, and be customized for multiple satellites (Sec.
-6).
-2 Background
+  • We introduce HoneySat, a novel, high-interaction, extensible honeypot framework for small satellites (Sec. IV).
+  • We present the Satellite Simulator, a library that simulates the physical processes, sensors, and subsystems necessary for a realistic satellite honeypot (Sec. V).
+  • We demonstrate HoneySat’s high level of realism through a hands-on survey where 10 experienced satellite operators interacted with HoneySat in real time (Sec. VI-C).
+  • We show HoneySat’s extensibility features by integrating two real-world satellite flight software and integrating HoneySat into a hardware-in-the-loop operation allowing our honeypot to communicate with an in-orbit operational SmallSat (Sec.
+VI-E and Sec.
+VI-F).
+II Background
 This section lays out key background concepts that are relevant to satellite honeypots.
 For honeypots: a description of the different existing types (Sec.
-2.1), as well as the current state-of-the-art (Sec.
-2.2).
+II-A), as well as the current state-of-the-art (Sec.
+II-B).
 For satellites: their operation (Sec.
-2.3), their architecture (Sec.
-2.4), existing Protocol Ecosystems (Sec.
-2.5), and tactics, techniques, and procedures (TTPs, Sec.
-2.6).
-2.1 Types of Honeypots
+II-C), architecture (Sec.
+II-D), existing protocol ecosystems (Sec.
+II-E), and tactics, techniques, and procedures (TTPs) (Sec.
+II-F).
+II-A Types of Honeypots
 Honeypots are categorized by interaction levels according to the interaction opportunities they provide. The two main types of honeypots are low-interaction and high-interaction.
 Low-Interaction Honeypots.
-These honeypots offer minimal interaction, simulating real systems through scripts or finite-state machines.
-Their advantages are ease of setup and maintenance due to low resource consumption, and a reduced risk of adversarial takeover.
-However, they provide limited interaction opportunities to adversaries which limits the interaction data they provide.
+These honeypots offer minimal interaction, simulating real systems through scripts.
+They are easy to setup and maintain due to low resource consumption, and have a reduced risk of adversarial takeover.
+However, they provide limited interaction opportunities to adversaries, which limits the data they provide.
 Low-interaction honeypot examples include Conpot [30] and Honeyd [31].
@@
 These honeypots offer extensive interaction opportunities via emulation or advanced simulations [31].
-Their main advantage is providing adversaries with almost limitless interactions, enabling them to provide extensive interaction data.
+Their main advantage is providing adversaries with almost limitless interactions, providing extensive interaction data.
 However, they pose a high risk of adversarial takeover as adversaries have more opportunities to hijack the honeypot [32].
 High-interaction honeypot examples include Cowrie [33] and HoneyPLC [23].
-2.2 Honeypot’s State of the Art
-The literature on honeypots includes hundreds of implementations that simulate a diverse set of computer systems [30, 48].
-From classic implementations that simulate a host’s TCP/IP stack, such as Honeyd [54], to modern approaches that integrate social media applications, such as HoneyTweet [3].
-However, there is no satellite honeypot in the literature.
-In the absence of a satellite honeypot, we now examine the honeypot approaches most related to satellites: Industrial Control Systems (ICS) [36, 35] and Unmanned Aerial Vehicles (UAV), a.k.a., drones summarized in Table 1.
-Satellite systems like ICS must be aware of some physical process, e.g., its position, the sun’s position, etc, via sensors to acquire data about the physical world.
+II-B Honeypot’s State of the Art
+The literature on honeypots includes hundreds of implementations that simulate a diverse set of systems [34, 35].
+From implementations that simulate a host’s TCP/IP stack, such as Honeyd [31], to modern approaches that integrate social media applications, such as HoneyTweet [24].
+In the absence of a satellite honeypot, we examine the approaches most related to satellites: Industrial Control Systems (ICS) and Unmanned Aerial Vehicles (UAV), summarized in Table I.
+Satellite systems like ICS must be aware of some physical process, e.g., the sun’s position, via sensors to acquire data about the physical world.
 Several ICS honeypots have simulated these physical processes.
-For example, ICSnet [58] and HoneyICS [38] simulate several components such as programmable logic controllers (PLCs), and actuators such as water valves.
-HoneyDrone [16] is a UAV honeypot that integrates different simulations, e.g., Ardupilot, to recreate multiple attack scenarios, including the UAV ground control station.
+For example, ICSnet [36] and HoneyICS [37] simulate components such as programmable logic controllers (PLCs) and actuators such as valves.
+HoneyDrone [38] is a UAV honeypot that integrates different simulations, e.g., Ardupilot, to recreate multiple attack scenarios.
 Although UAV honeypots share some similarities with satellites [39], satellite honeypots require additional physical simulations and involve a more complex operation environment, which we discuss in Sec.
-2.3.
-Table 1: Comparison of Existing Honeypots and HoneySat.
-Keys: = Supported; $\times$× = Not Supported.
+II-C.
+TABLE I: Comparison of Existing Honeypots and HoneySat.
+Keys: = Supported; $\times$ = Not Supported.
 Honeypot/
@@
 ICSPot [40]
-High 4 1 $\times$×
+High 4 1 $\times$
 HoneyICS[37]
@@
 HoneyDrone [38]
-Medium 4 1 $\times$×
+Medium 4 1 $\times$
 HoneySat
@@
 in Section
-4.4,
-4.3
-4.3,
-5.2.1
-4.4,
-5.1
-6.5
-2.3 Anatomy of a Satellite Mission
-Figure 1: The components commonly found in a satellite mission in the context of the space and ground segments.
+IV-D,
+IV-C
+IV-C,
+V-B1
+IV-D,
+V-A
+VI-E
+VI-F
+II-C Anatomy of a Satellite Mission
 We now describe the components of a satellite mission.
-Due to satellite missions’ complexity, we explain each component and match it with one of the numbers in Fig.
+Due to satellite missions’ complexity and diversity, we explain each component and match it with one of the numbers in Fig.
 1.
@@
 Ground Segment 1.
-The Ground Segment (GS) covers the terrestrial supporting infrastructure required for a successful satellite operation.
-It consists of a ground station, responsible for exchanging data with the spacecraft, the computational and network infrastructure required for communication, but also the systems to operate the satellites, e.g., servers, databases, and user interfaces [81].
-Several ground stations can be connected in a network and coordinated as part of one GS.
-The GS includes the Ground Segment Software (GSS) that helps operators schedule and send commands and visualize data that is sent back in response.
+The Ground Segment (GS) covers the terrestrial infrastructure required for a successful satellite operation.
+It consists of a ground station, responsible for exchanging data with the spacecraft, the computational and network infrastructure required for communication, and the systems to operate the satellites, e.g., servers [12].
+The GS includes the Ground Segment Software (GSS) that helps operators schedule and send commands and visualize data.
 Space Segment 2.
@@
 The basic data flow between the space and ground segments are TC and TM [41].
-TM is the data the satellite sends to the ground station which may contain the satellite’s status, responses to previous commands, or payload data [69].
+TM is the data the satellite sends to the ground station which may contain the satellite’s status or payload data [41].
 TCs are used to operate the satellite and are transmitted and encapsulated in a space protocol (see Sec.
-2.5).
+II-E).
 The design and implementation of TCs varies depending on the satellite mission.
@@
 In some cases, such as geostationary orbit (GEO), the satellite can be above the horizon constantly.
-The occurrence of passes can be calculated using the spacecraft’s orbital elements.
-A common representation of the orbital elements is the two-line element set (TLE) [77].
+The occurrence of passes can be calculated using the spacecraft’s orbital elements in the two-line element (TLE) data [43].
 Satellite Mission Operations 6.
-Satellite mission operations vary widely depending on the owning organization, budget, and technology.
+Satellite mission operations vary widely depending on the owning organization’s budget, and technology [5].
 Nevertheless, they share some commonalities, which we now describe.
-A mission’s operation involves a team of operators that use GSS to operate a satellite or a satellite constellation and ensure the mission’s success [46].
-Operations are carried out in a Mission Operations Center (MOC), where satellite operators sit at their workstations to manage TCs sent to the satellite(s).
-Satellite operators may also remotely operate satellites 7 by connecting to the ground segment using tools such as desktop sharing like VNC (Virtual Network Computing) [10], or even rely on autonomous or semi-autonomous operations [26].
+A mission’s operation involves a team of operators that use ground segment software (GSS) to operate a satellite(s) and ensure the mission’s success [44].
+Operations are carried out in a Mission Operations Center (MOC), where operators sit at their workstations to manage TCs sent to the satellite(s).
+Operators may also remotely operate satellites 7 by connecting to the ground segment using tools such as VNC (Virtual Network Computing) [11].
 Satellite operations include two main activities: satellite tracking and TC generation and scheduling.
-Satellite tracking calculates the satellite’s position in orbit and controls the ground station(s)’ tracking antenna to establish communication between the ground and space segments.
-TC generation and scheduling crafts command to be programmed and sent to the satellite so it can perform different functions, e.g., read systems status data, schedule orbital maneuvers, schedule payload operations, and download payload data.
+Satellite tracking calculates the satellite’s position in orbit and controls the ground station’s tracking antenna to establish communication between the ground and space segments.
+TC generation and scheduling crafts commands to be sent to the satellite to perform different functions, e.g., download payload data.
+Figure 1: The components commonly found in a satellite mission in the context of the space and ground segments.
 Ground (Segment) Software (GSS).
-GSS allows satellite operators to carry out the satellite mission’s routine operations as described in the previous paragraph.
+GSS allows operators to carry out the mission’s routine operations.
 GSS are very diverse.
-Some satellite missions develop their own GSS while others use open-source [65] or buy proprietary GSS like GSWeb [23].
+Some satellite missions develop their own GSS while others use open-source [45] or proprietary GSS [46].
 There are two main types of GSS: Mission Control Software (MCS) and Ground Station Control Software (GSCS).
-Mission control software 8 manages TCs and scripts to be sent to the satellite and can display TM.
-For example, ESA’s SCOS-2000 is an MCS that provides generic functionality that can be customized for a specific mission to cover the functions required for TM reception and processing and TC verification [42].
-Some satellite missions develop their own MCS.
-For example, the SUCHAI mission team developed their MCS application to send and receive TC/TM [65].
-Ground station control software (GSCS) 9 helps satellite operators track and visualize the satellite’s orbit and provide detailed information about each satellite’s pass.
+Mission control software 8 manages TCs and scripts to be sent to the satellite and display TM.
+For example, ESA’s SCOS-2000 is an MCS that provides generic functionality that can be customized for a specific mission operation [47].
+Some missions develop their own MCS.
+For example, the SUCHAI mission developed their own MCS application [45].
+Ground station control software (GSCS) 9 helps satellite operators track and visualize the satellite’s orbit and provide information about each satellite’s pass.
 For example, Gpredict is a popular open-source GSCS that performs real-time satellite tracking and orbit prediction [48].
-Gpredict can also interface with and control the radio system and antenna rotor during the satellite pass.
-2.4 Satellite Architecture
-Satellite architectures are varied and complex, however, here we describe the most common terminology depicted in Fig. 1’s space segment 2.
-When referring to satellite architecture there is a distinction between the platform that facilitates the successful operation of the satellite’s day to day-to-day activities and the payload.
+II-D Satellite Architecture
+Satellite architectures are varied and complex [49]; however, here we describe the most common terminology depicted in Fig. 1’s space segment 2.
+When referring to satellite architecture there is a distinction between the platform that facilitates the successful operation of the satellite’s activities and the payload.
 The platform provides the possibility to run a payload that fulfills the purpose of these missions.
-This payload differs based on the goal of the mission and can range from instruments to measure physical properties to communications systems providing wireless connectivity.
-Platform: The platform, or satellite bus, is a system composed of custom-designed or off-the-shelf subsystems necessary for critical satellite operations.
-These include, among others, the Attitude Determination and Control System (ADCS) to maintain the satellite’s orientation (i.e.
+Payloads are different depending on the satellite’s mission and can range from instruments to measure physical properties to communications systems providing wireless connectivity.
+Platform: The platform is composed of custom-designed or off-the-shelf subsystems necessary for critical satellite operations.
+These include the Attitude Determination and Control System (ADCS) to maintain the satellite’s orientation (i.e.
 attitude) and position to keep the satellite pointed towards antennas or solar panels illuminated; the Electrical Power Subsystem (EPS) for managing power generation and distribution; the Communication Subsystem (COMM) and the Command and Data Handling (C&DH) subsystems to facilitate communications for receiving TC and sending TM and controlling the satellite operations.
-All of these subsystems are then controlled via TCs sent to the satellite.
-TCs may be, depending on the protocol ecosystem and complexity of the subsystem, interpreted by a central C&DH System or merely forwarded to the recipient subsystem for further processing [84].
-This managerial duty is left to Flight Software (FS) running on an embedded system of differing complexity.
-Some examples include NASA’s Core Flight System (cFS) [41] and F’ (F Prime) [6], KubOS [83], the German Aerospace Center (DLR)’s OUTPOST [17], and the University of Chile’s SUCHAI FS [24].
-Attackers aim to gain the ability to send Telecommands to the Flight Software, as this typically grants control over all spacecraft subsystems.
+These subsystems are controlled via TCs sent to the satellite.
+TCs may be interpreted by a central C&DH System or merely forwarded to the recipient subsystem for further processing [50].
+This managerial duty is handled by the Flight Software (FS) running on an embedded system.
+Some examples include NASA’s Core Flight System (cFS) [51] and the University of Chile’s SUCHAI FS [52].
+Attackers aim to gain the ability to send TCs to the flight software as this may grant control over all spacecraft subsystems.
 Payload.
-The payload is the equipment that the satellite employs to fulfill its mission.
+The payload is the equipment that a satellite employs to fulfill its mission.
 Due to satellites’ varied missions, payloads are heavily customized [53].
 For example, if a satellite’s mission is remote sensing, its payload may include an infrared camera [54].
-craft’s essential functions, acts on TCs, and handles internal data transmitted to and from other systems [84].
-The CDHS is comprised of Flight Software (FS) running on an OBC [81].
-2.5 Small Satellite Protocol Ecosystems
-SmallSat missions can often be categorized into the following categories by the adoption of protocols and their corresponding philosophies.
+II-E Small Satellite Protocol Ecosystems
+SmallSat missions can often be categorized by the adoption of protocols and their corresponding philosophies.
+Currently, there are two major protocol ecosystems which SmallSat missions can adopt: CSP and CCSDS.
 Cubesat Space Protocol.
-The Cubesat Space Protocol (CSP) family of protocols is a one-stop solution for Small missions [63].
-There are not a lot of choices left to the operator/developer and there are two vendors mainly offering components using this protocol and some that offer compatibility solutions [23].
-If a mission built on CSP is expanded by a commercial subsystem it will plug in without issue after configuration.
-CSP is implemented as an open-source C library called libCSP [63].
-CSP follows the TCP/IP model, including transport and routing protocols and multiple layer 2 interfaces such as I2C (Inter-Integrated Circuit), CAN (Controller Area Network), and ZeroMQ (ZMQ) for transmission on TCP/IP networks [72].
-In CSP the ground segment interfaces and satellite subsystems are part of a CSP network.
-Sending TC is as simple as sending a CSP packet with an address corresponding to a node that is part of the satellite.
-The configured static routing tables of the nodes on the network will then cause the packet to be passed to the ground station and transmitted to the satellite over RF where a system on board will route the packet to its destination.
-In case the satellite is not currently doing a pass, the packet is dropped.
-The straightforwardness of CSP makes it particularly interesting for the honeypot use case as the internal structures of any CSP-based mission will look very similar, so distinguishing two missions is difficult as they will use the protocols in a very similar way.
+The Cubesat Space Protocol (CSP) family of protocols is a one-stop solution for SmallSat missions [55].
+CSP is implemented as an open-source C library called libCSP [55] and follows the TCP/IP model, including transport and routing protocols and multiple layer 2 interfaces such as I2C (Inter-Integrated Circuit), and ZeroMQ (ZMQ) for transmission on TCP/IP networks [56].
+In the context of a satellite mission, CSP connects the ground segment and satellite subsystems as part of a CSP network where each subsystem is identified as a node.
+Sending a TC is as simple as sending a CSP packet with the address of the corresponding subsystem node.
+In case the satellite is not currently performing a pass, the packet is dropped.
 There are also default services running on standardized ports that are enabled by default when building libCSP [55], for example for network diagnostic (ping) or basic operations (rebooting).
-In CSP missions, nodes on the network typically feature a command-line-based interface for configuration and debugging purposes [23].
-This may allow someone with access to it to interact with the current node and other nodes on the network.
+The straightforwardness of CSP makes it particularly relevant for the honeypot use case as the internal structures of any CSP-based mission will look similar, so distinguishing two missions may be difficult.
 CCSDS Space Communication Protocols.
-The CCSDS Space Communication protocols are a vast set of standardized protocols used for different purposes in space communications.
-A major CCSDS protocol relevant for SmallSat TM and TC is called spacepacket.
-This protocol is used in combination with the ECSS Packet Utilization Standard (PUS) to define how TCs and TM are encoded and transported.
-The PUS defines services (and thus sets of TC/TM) for functionality that satellite missions likely require, including large data transfer or event reporting [59].
-Moreover, mission designers can tailor the PUS standard by selecting a subset of services and sub-services relevant to their mission needs.
-Furthermore, it is possible to define custom Services or implement numerous custom functionalities.
+The Consultative Committee for Space Data Systems (CCSDS) Space Communication protocols are a set of standardized protocols used for different purposes in space communications [57].
+A relevant CCSDS protocol for SmallSat TM and TC is the Space Packet protocol.
+This protocol is used in combination with the ECSS Packet Utilization Standard (PUS) [58] to define how TCs and TM are encoded and transported.
+PUS defines services (and thus sets of TC/TM) for functionality that satellite missions require, including large data transfer or event reporting [57].
 Proprietary Protocols.
-In addition to these widespread ecosystems, vendor-specific sets of protocols may be usable on top of more standard protocols or transport higher-layer packets of known protocols.
-Some basic small satellites may also have entirely ad hoc protocols that are limited to use with a single mission or a set of satellites by a specific institution.
-These proprietary protocols are not ideal candidates for constructing honeypots as an ecosystem of one mission that is unique to a small set of missions will look out of place for other missions, preventing generalization.
-The choice of ecosystem also influences the choice of GSS as an MCS usually focuses on a single protocol stack.
+In addition to CSP and CCSDS, vendor-specific protocols may be used on top of more standard protocols.
+Some basic small satellites may have entirely ad-hoc protocols used on a single mission, for example the ESTCube-1 uses a custom protocol [59].
+These proprietary protocols are not ideal candidates for constructing honeypots as the ecosystem of one mission will look out of place for other missions, preventing generalization.
+The choice of ecosystem also influences the choice of GSS as an MCS usually supports a single protocol stack [45].
 In addition to the space protocol ecosystems above, satellite operators use well-known network protocols to connect with the mission’s ground segment services.
 These protocols include Telnet, SSH, VNC, FTP, and HTTP.
-Telnet, SSH, and VNC are used to remotely connect to the MOC workstations.
-HTTP is used for web interfaces that operators use to visualize mission data, and FTP is used to download TM data from the mission data repository.
+Telnet, SSH, and VNC are used to remotely connect to the MOC workstations [11].
+HTTP is used for web interfaces that operators use to visualize mission data.
+FTP is used to download TM data from the mission data repository.
 If an adversary were to compromise any of these services, they could use it as a stepping stone to target the satellite itself.
-2.6 Space Systems’ Tactics, Techniques and Procedures (TTPs)
-Tactics, Techniques, and Procedures (TTPs) describe the behavior of a malicious actor in a structured scheme to understand how they might execute an attack [56, 47].
+II-F Space Systems’ Tactics, Techniques and Procedures (TTPs)
+Tactics, Techniques, and Procedures (TTPs) describe a malicious actor’s behavior in a structured scheme to understand how they might execute an attack [60, 61].
 Several frameworks have been introduced to standardize space systems’ TTPs.
@@
 These matrices list and describe space security-specific tactics and techniques such as initial access and ground segment compromise.
-3 Threat Model
+For the remainder of this paper we use the SPACE-SHIELD matrix to analyze HoneySat.
+III Threat Model
 Following Fig.
-1, we assume an adversary willing to compromise a satellite can only interact with the space segment simulation by sending TCs from the ground segment first.
-To gain initial access to the ground segment, an adversary needs to connect via one of exposed network protocols depicted in Sec.
-2.5, namely, VNC, Telnet, TCP/IP, etc., which correspond with the operational protocols used in real satellite missions.
-From there, an adversary may try to launch different commands to take full control and/or compromise the services offered by the satellite’s mission. Finally,
-in this paper, the modeling of physical radio communication between the space and ground segments, which is commonly used in practice, is considered out of scope and left for future work.
-Figure 2: HoneySat framework high-level architecture design.
-4 HoneySat Framework Design
+2, we assume that an adversary willing to compromise a satellite can only interact with the space segment simulation by sending TCs from the ground segment first.
+To gain initial access to the ground segment, an adversary needs to connect via one of the exposed network protocols discussed in Sec.
+II-E, namely, VNC, Telnet, TCP/IP, etc., which correspond to the operational protocols used in real satellite missions.
+From there, an adversary may try to launch different commands to take full control and/or compromise the services offered by the satellite’s mission as depicted in Fig.
+3.
+Finally, in this paper, the modeling of physical radio communication between the space and ground segments, which is commonly used in practice, is considered out of scope and left for future work.
+The threat model is further referenced in Sec.
+IV-E.
+IV HoneySat Framework Design
 In this section, we explain the objectives we aim to achieve (Sec.
-4.1), the design principles we follow to meet such objectives (Sec.
-4.2), and the overall design of our Space  (Sec.
-4.4), and Ground (Sec.
-4.3) segments.
-4.1 HoneySat’s Design Objectives
+IV-A), the design principles we follow to meet such objectives (Sec.
+IV-B), and the overall design of our Space  (Sec.
+IV-D), and Ground (Sec.
+IV-C) segments.
+IV-A HoneySat’s Design Objectives
 Our design aims to achieve the following objectives:
@@
 As explained in Sec.
-1, the number one objective of any honeypot is to capture interaction data from which we can derive knowledge on adversaries’ TTPs.
-As such, our first objective for HoneySat is to have capability to capture rich interaction data.
+I, the number one objective of any honeypot is to capture interaction data from which we can derive knowledge on adversaries’ TTPs.
+As such, our first objective for HoneySat is to have the capability to capture rich interaction data.
   2. DO-2
@@
 As we discussed in Sec.
-1, honeypots’ nature must remain covert to entice adversaries into interacting with it.
+I, honeypots’ nature must remain covert to entice adversaries to interact with it.
 As such, our second objective is for HoneySat’s nature to remain hidden from adversaries.
@@
     Provide Extensibility and Customizability.
-A framework’s main purpose is to provide generic functionality that can be customized to meet the user’s needs, in HoneySat’s case, we must be able to support multiple SmallSats.
+A framework’s main purpose is to provide generic functionality that can be customized to meet the user’s needs.
+In HoneySat’s case, we must be able to support multiple SmallSats.
 For example, a particular SmallSat may use CSP or CCSDS ecosystems.
@@
 As such, our third objective is for HoneySat to provide extensibility and customizability.
-4.2 HoneySat’s Design Principles
+Figure 2: HoneySat framework high-level architecture design.
+IV-B HoneySat’s Design Principles
 To meet our objectives, we selected the following principles:
@@
 As we described in Sec.
-2.1, high-interaction honeypots give adversaries the same or almost identical interaction opportunities as a real satellite.
-Thus, they can log many of the TTPs discussed in 2.6.
+II-A, high-interaction honeypots give adversaries the same or almost identical interaction opportunities as a real satellite.
 For these reasons, we selected high-interaction simulation as our first design principle.
@@
 As we discussed in Sec.
-2.2, satellites keep track of their orbit location, the position of the sun, among others.
+II-B, satellites keep track of their orbit location, the position of the sun, among other details.
 These details must be simulated by our honeypot; otherwise, they may alert adversaries that they are interacting with a fake system.
+For this reason, we selected realistic simulation as our second design principle.
 This design principle is based on design objective DO-2.
@@
 As we explained in Sec.
-2, satellites are complex and diverse systems.
+II, satellites are complex and diverse systems.
 There is no one-size-fits-all solution.
@@
 2, consists of two sets of simulations, the space segment simulations and the ground segment simulations.
-Table 2 illustrates how HoneySat’s simulation components match the components of a real satellite mission.
+Table II illustrates how HoneySat’s simulation components match the components of a real satellite mission.
+TABLE II: Correspondence between real satellite mission components and HoneySat’s design components.
 Segment Real Component HoneySat Component
@@
 Space Payload Satellite Simulator
-Table 2: Correspondence between real satellite mission components and HoneySat’s design components.
-4.3 Ground Segment Design
+IV-C Ground Segment Design
 Following Sec.
-2.3, the purpose of the HoneySat’s ground segment is to simulate the ground segment assets, e.g., the ground segment software.
+II-C, the purpose of the HoneySat’s ground segment (depicted in Fig.
+IV) is to simulate the ground segment assets, e.g., the ground segment software.
 To accomplish this, our design includes the following components: 1) the Exposed Network Protocols, 2) the Ground Segment Software, 3) the Radio Simulator, 4) the Ground Personality, and 5) the Logging Repository.
@@
 To provide adversaries with feasible access to our honeypot, HoneySat exposes multiple means of interaction over a network such as the Internet.
-We designed HoneySat to support four different interaction methods using different protocols, namely VNC, Telnet, Web, and Access to the ground station via raw packet transmitting capability (following design principle DP-3).
+We designed HoneySat to support four different interaction methods using different protocols, namely VNC, Telnet, Web, and Access to the ground station via raw packet transmitting capability; following the design principle DP-3.
+We selected these protocols based on data obtained from our satellite operator survey, discussed in Sec.
+VI-C, which revealed that satellite operators do use remote access tools, such as web interfaces and screen sharing, to operate satellites.
 2) Ground Software.
 As discussed in Sec.
-2.3, the ground software includes mission control software (MCS) and ground station control software (GSCS).
-We leverage existing ground software, allowing HoneySat to provide a high-interaction simulation (DP-1).
+II-C, the ground software includes mission control software (MCS) and ground station control software (GSCS).
+We leverage existing ground software, allowing HoneySat to provide a high-interaction simulation; thereby following design principle DP-1.
 3) Radio Simulator.
 As discussed in Sec.
-2.3, communication between the satellite and the ground station is possible only during a pass.
-The purpose of the Radio Simulator is to control communication between the ground and space segment simulations to simulate real orbital passes.
-The Radio Simulator design follows design principles DP-2 and DP-3.
-4) Ground Configuration. The ground configuration is a series of settings for the various ground segment-specific configurations, e.g., the satellite mission logo in the Web Interface.
+II-C, communication between the satellite and the ground station is possible only during a pass.
+The purpose of the Radio Simulator is to mimic real orbital passes by enabling and disabling communication between HoneySat’s ground and space segment simulations at the appropriate times.
+The Radio Simulator design follows the design principles DP-2 and DP-3.
+4) Ground Configuration. The ground configuration is a series of settings for the ground segment-specific configurations, e.g., the satellite mission logo in the Web Interface.
 5) Logging Repository.
-This repository logs data such as the TM/TC traffic to and from the ground station software and the logging attempts received in the web interface.
+This repository records data such as the TM/TC traffic to and from the ground station software and the logging attempts received in the web interface.
 We designed the ground segment simulation logs to be categorized and timestamped.
-The logging design follows design principle DP-3.
-4.4 Space Segment Design
+The logging repository design follows design principle DP-3.
+IV-D Space Segment Design
 The purpose of the HoneySat’s space segment is to mimic the spacecraft, as discussed in Sec.
-2.3.
+II-C.
 To accomplish this, our design includes the following components: 1) the satellite’s Flight Software Runtime, 2) the Satellite Simulator, 3) the Flight Software Services, 4) the Satellite Personality, and 5) the Logging Repository.
@@
 As discussed in Sec.
-2.4, the satellite flight software manages all critical functions required for the mission operation, such as interacting with hardware peripherals, processing TCs, and sending TM.
+II-D, the satellite flight software manages all critical functions required for the mission operation, such as interacting with hardware peripherals, processing TCs, and sending TM.
 For this to work we need an environment where services that handle TC intended to run on flight software can be run.
@@
 One of the biggest challenges when designing HoneySat was simulating all the physical processes, e.g., attitude, that satellites need to know.
-The Satellite Simulator is designed to solve this problem.
-It simulates all the necessary satellite subsystems, e.g., the Electrical Power System, and the sensors, e.g., a GPS receiver, that a satellite requires to function.
-The purpose of the Satellite Simulator is to abstract away the various hardware peripherals that communicate with the flight software by simulating these peripheral’s data.
-For example, suppose a particular flight software was implemented to utilize a specific lithium-ion battery pack.
-In that case, the Satellite Simulator’s power system simulation simulation can simulate the basic behavior of the battery.
-We designed the Satellite Simulator to be modular so that new subsystems and sensors can be easily added, modified, or disabled, following our DP-3.
+The Satellite Simulator is designed to solve this problem by simulating all the necessary satellite subsystems, e.g., the Electrical Power System, and the sensors, e.g., a GPS receiver, that a satellite requires to function.
+The purpose of the Satellite Simulator is to abstract away the various hardware peripherals that communicate with the flight software by simulating these peripherals’ data.
+For example, suppose that a particular flight software was implemented to utilize a specific lithium-ion battery pack.
+In that case, the Satellite Simulator’s power system simulation can simulate the basic behavior of the battery.
+We designed the Satellite Simulator to be modular so that new subsystems and sensors can be easily added, modified, or disabled according to our design principle DP-3.
 3) Flight Software Services.
@@
 The satellite state module routes the service’s messages to and from the Satellite Simulator.
-It uses a list of message IDs that can be customized based on the protocol ecosystem and software architecture.
+It uses a list of message IDs that can be customized based on the protocol ecosystem and software architecture complying with our design principles DP-1 and DP-3.
 4) Satellite Personality.
@@
 It includes FS and Satellite Simulator configuration values such as the satellite’s battery capacity.
-The satellite personality allows our framework to be easily customized.
+The satellite personality makes our framework easily customizable; thus following our design principle DP-3.
 5) Logging Repository. We designed each space segment component to provide detailed logs. As shown in Fig. 6, all Satellite Simulator modules can send their own logs.
-4.5 Theory of Operation
+IV-E Theory of Operation
 Figure 3: HoneySat’s Theory of Operation.
@@
 An adversary gets initial access via one of the Exposed Network Protocols ①.
-On interaction with the Simulated Ground Segment, an attacker is presented with access to the Ground Software ②.
+On interacting with the Simulated Ground Segment, an attacker is presented with access to the Ground Software ②.
 The configuration of this ground software is defined by the Ground Configuration, which allows the ground software to look like one of many different missions from its protocol ecosystem to the attacker.
-The attacker can then interact with the Ground Software, while their actions are reported to the Log Database.
+The attacker can then interact with the Ground Software, while their actions are reported to the Logging Repository.
 They might want to try to gain more privileges on the ground segment or try to send TC to the space segment.
@@
 When the Flight Software Runtime receives a command, it will run the corresponding Flight Software Service to compute a response ⑤.
-In case it requires either changes to an on-board system’s state or a sensor value to execute the TC, it will invoke the Satellite Simulator for this ⑥.
+In case it requires either changes to an on-board system’s state or a sensor value to execute the TC, it will invoke the Satellite Simulator ⑥.
 The Satellite Personality contains the configuration used by the Satellite Simulator.
-Once a TC is processed, the resulting TM is routed back through the Radio Simulator to either the attacker or the ground software used by the attacker ⑦-⑫, and is also recorded in our Log Database.
-5 HoneySat’s Implementation
+Once a TC is processed, the resulting TM is routed back through the Radio Simulator to either the attacker or the ground software used by the attacker ⑦-⑫, and is also recorded in our Logging Repository.
+V HoneySat’s Implementation
 Having laid out HoneySat’s design we now explain how we implemented the Satellite Simulator (Sec.
-5.1), the generic CSP mission honeypot (Sec.
-5.2), the ground segment (Sec.
-5.2.1), and the space segment (Sec.
-5.2.2).
-5.1 Satellite Simulator Implementation
+V-A), the generic CSP mission honeypot (Sec.
+V-B), the ground segment (Sec.
+V-B1), and the space segment (Sec.
+V-B2).
+V-A Satellite Simulator Implementation
 We implemented the Satellite Simulator as a Python object-oriented programming application.
@@
 Many of the basic simulations and corresponding sensors can be shared, especially in a honeypot where attackers lack detailed information that would enable them to distinguish between small differences in sensor values.
-Further details on the simulations can be found in A.2
-5.2 Generic CSP Mission Honeypot
-Following our design we implemented HoneySat to support the CSP ecosystem.
-This implementation provides a generic CSP-based mission honeypot that can be customized to simulate a specific CPS-based satellite missions.
-5.2.1 Ground Segment Implementation
+Further details
+can be found in Appendix A-B.
+V-B Generic CSP Mission Honeypot
+Following our design, we implemented HoneySat to support the CSP ecosystem.
+This implementation provides a generic CSP-based mission honeypot that can be customized to simulate a specific CPS-based satellite mission.
+V-B1 Ground Segment Implementation
 We now describe how we implemented the five ground segment simulations designed in Sec.
-4.3.
+IV-C.
 1) the Exposed Entry Points, 2) the Ground Software, 3) the Radio Proxy, 4) the Ground Configuration, and 5) Logging Repository.
@@
 We implemented the VNC server using TigerVNC [64] and used PyZMQ [65] for the Raw Packet Access.
-The Telnet server was implemented using Python to expose the command-line interface of GSS on the network.
+The Telnet server was implemented using Python to expose the command-line interface of the Ground Segment Software (GSS) on the network.
 The web interface allows for the customization of text presented to the attacker through configuration.
 Using the above entry points, adversaries can interact with the GSS and the space segment simulations.
-Among these, the web interface serves as a measure to attract attackers and serves as an initial entry point for attackers.
-We developed a simple yet convincing web application, presented as an early version of a web interface currently under development.
-After the login, the attacker would be welcomed by a dashboard that displays information about the satellite and the latest output of the Ground software command line interface.
-Additionally, the web application provides access to documentation and the VNC server (Fig.
-4), facilitating the attacker in finding the other entry points.
 Figure 4: The view an attacker would see upon connection to the VNC server. Ground Station Control Software (Red) Ecosystem/Mission Specific MCS (Yellow).
 2) Ground Software.
-To implement the MCS software for our CSP-based honeypot, we require an open-source MCS for the CSP ecosystem.
+To implement the MCS software for our CSP-based honeypot, we require an MCS that supports the CSP ecosystem.
 We leveraged the SUCHAI mission control software (SUCHAI MCS).
@@
 This is a good choice for the generic CSP MCS as it can be passed off as the CLI of commercial ground segment appliances like the GomSpace NanoCom MS100 [46].
-In this generic honeypot, we can set mission-specific parameters through the ground configuration.
-Additionally, we provide a desktop environment accessible via VNC, complete with supporting applications such as configured Gpredict for the satellite mission.
+Given HoneySat being a generic honeypot, we can set mission-specific parameters through the ground configuration.
+Additionally, we provide a desktop environment accessible via VNC, complete with supporting applications such as a configured Gpredict for the satellite mission.
 For certain deployments, this environment also includes a mission-specific GUI tool, both of which appear as open windows on the desktop.
 3) Radio Simulator.
-The Radio Simulator was implemented as a ZMQ Hub.
-The ZMQ Hub functions as a router for the CSP network.
-Simulating both the ground part of the CSP network including the ground station and space segment.
+The Radio Simulator was implemented as a ZMQ Hub [66].
+The ZMQ Hub functions as a router for the CSP network simulating both the ground part of the CSP network including the ground station and the space segment.
 Depending on whether the satellite is currently passing, packets to the space segment are forwarded to the Flight Software Runtime.
-We decided to use ZeroMQ as it is the protocol to use for transmitting CSP over TCP/IP [64].
-4) Ground Configuration. We implemented the ground configuration on each component, and the settings can be modified through a Docker compose file explained in Appendix A.3.
+We decided to use ZeroMQ as it is the protocol to use for transmitting CSP packets over TCP/IP [66].
+4) Ground Configuration. We implemented the ground configuration as settings that can be modified through a Docker compose file explained in Appendix A-C.
 5) Logging Repository.
@@
 Additionally, network traffic to and from the host running HoneySat is captured.
-5.2.2 Space Segment Implementation
-1) Flight Software Runtime Although most of the flight software frameworks presented in Section 2.4 are open-source projects, it is difficult to find a real satellite flight software implementation that is available for study and modification [81].
-For this reason, we use the University of Chile’s SUCHAI Flight Software (SUCHAI FS) [24], which has been deployed in four satellite missions [21].
+V-B2 Space Segment Implementation
+We now describe how we implemented the four space segment simulations designed in Sec.
+IV-D.
+1) the Flight Software Runtime, 2) the Flight Software Services, 3) the Satellite Personality,and 4) the Logging Repository.
+1) Flight Software Runtime.
+Although most of the flight software frameworks presented in Section II-D are open-source projects, it is difficult to find a real satellite flight software implementation that is available for study and modification [12].
+Thus, we use the University of Chile’s SUCHAI Flight Software (SUCHAI FS) [52], which has been deployed in four satellite missions [68].
 The SUCHAI FS provides flight software based on the CSP ecosystem that already features a mode where it can be run on a POSIX OS.
@@
 There are 96 of these commands and parameters that are parsed from CSP packets using a text-based protocol.
+We modified the command handlers to use the Satellite Simulator instead of commanding subsystems and querying real sensors.
 3) Satellite Personality.
@@
 The current implementation includes almost 50 configurable variables, such as the satellite’s name, the battery’s nominal cell capacity, etc.
-4) Logging Repository. Like the ground segment’s logging repository (Sec 5.2.1), The Satellite Simulator and the FS are connected to the MongoDB instance using a MongoDB client.
-Each logged data entry is created with customized tagging and a timestamp. The collected data is structured JSON [32] format and stored in the MongoDB database.
-Finally, we containerized and hardened this version of the honeypot implementation for secure and easy deployment, which we describe in Appendix A.3.
-6 Evaluation
-Having described HoneySat’s implementation in Sec.
-5, we now evaluate it.
-This section lists three experimental questions designed to test HoneySat’s alignment with our design objectives.
-Next, we present four sets of experiments that provide empirical evidence confirming that our design objectives have been achieved within HoneySat.
-We describe each experiment’s environment, methodologies, and results.
-6.1 Experimental Questions
-These questions aim to determine whether or not HoneySat’s meets the design objectives outlined in Sec. 4.1.
+4) Logging Repository. Like the ground segment’s logging repository (Sec V-B1), the Satellite Simulator and the FS are connected to the MongoDB instance using a MongoDB client.
+Each logged data entry is created with a customized tag and a timestamp. The collected data is structured JSON [69] format and stored in the MongoDB database.
+Finally, we containerized and hardened this version of the honeypot implementation for secure and easy deployment, which we describe in Appendix A-C.
+VI Evaluation
+In order to evaluate HoneySat we start by listing three experimental questions designed to test its alignment with our design objectives.
+Next, we present four sets of experiments that provide empirical evidence confirming that our design objectives have been addressed as intended.
+We then describe each experiment’s environment, methodologies, and results.
+VI-A Experimental Questions
+The following questions aim to determine if HoneySat meets the design objectives outlined in Sec. IV-A.
   1. Q-1
@@
     Since capturing data on varied techniques is the purpose of any honeypot, we explore the capabilities of HoneySat, as described in Sec.
-4.
+IV.
 This question is related to design objective DO-1 and is addressed in Sec.
-6.2 and 6.4.
+VI-B.
   2. Q-2
     Can HoneySat simulate a SmallSat mission well enough to deceive adversaries?
-    After enticing an adversary HoneySat must keep its true nature hidden, we need to simulate the satellite’s communication and physics characteristics described in Sec.
-2.4.
+    After enticing an adversary, HoneySat must keep its true nature hidden.
+Thus, HoneySat needs to simulate the satellite’s communication and physics characteristics described in Sec.
+II-D.
 This question relates to design objective DO-2 and is answered in Sec.
-6.3.
+VI-C and VI-D.
   3. Q-3
-    Can HoneySat simulate different SmallSat missions?
-    HoneySat’s customization is important because it would allow users of our framework to implement their own honeypots. This question relates to design objective DO-3, and we answer it in Sec. 6.5.
-To answer the above research questions we conducted four experiments.
+    Can HoneySat be extended to simulate different SmallSat missions and support hardware-in-the-loop simulations?
+    HoneySat’s customization is important because it would allow users of our framework to implement their own honeypots and even support satellite hardware integration.
+This question relates to design objective DO-3, and we answer it in Sec.
+VI-E and Sec.
+VI-F.
+To answer the above research questions, we conducted five experiments.
 First, in Sec.
-6.2 we craft multiple attacks in a controlled environment to quantify the level of interaction that HoneySat provides.
+VI-B we craft multiple attacks in a controlled environment to quantify the level of interaction that HoneySat provides.
 Second, in Sec.
-6.3 we conduct a survey with experienced satellite operators to evaluate HoneySat’s realism.
+VI-C we conduct a survey with experienced satellite operators to evaluate HoneySat’s realism.
 Third, in Sec.
-6.4 we deploy HoneySat and expose it to the Internet to test its deception capabilities.
-Fourth and final, in Sec.
-6.5 we add an additional protocol ecosystem to HoneySat to test its extensibility potential.
-6.2 TTP Interaction Experiment
-This experiment seeks to answer Q-1 by quantifying the different interactions that HoneySat provides.
+VI-D we deploy HoneySat and expose it to the Internet to test its deception capabilities.
+Fourth, in Sec.
+VI-E we test HoneySat’s extensibility by integrating a completely different flight software.
+Fifth and final, in Sec.
+VI-F we test HoneySat’s support for hardware-in-the-loop operations.
+VI-B TTP Interaction Experiment
+This experiment seeks to answer Q-1 by quantifying the interactions provided by HoneySat.
 To achieve this, we leveraged the SPACE-SHIELD matrix (Version 2.0) [63] discussed in Sec.
-2.6.
+II-F.
 SPACE-SHIELD provides a collection of adversary tactics and techniques for space systems.
@@
 For example, the technique Compromise Hardware Supply Chain involves “replacing a hardware component in the supply chain with a custom or counterfeit part” which is out of the scope of a virtual honeypot.
-Taking this into consideration, 14 tactics, and 33 techniques are applicable to HoneySat.
+Taking this into consideration, 14 tactics are applicable to HoneySat.
+From these tactics, HoneySat can feasibly offer up to 38 techniques as possible interactions for adversaries.
 Experiment Description.
@@
 Experiment Methodology.
-We designed one interaction or exploit for each of the applicable 33 techniques for our honeypot.
+We designed one interaction or exploit for each of the 38 applicable techniques for our honeypot.
 The interactions involved ground segment simulations such as the web interface.
 The exploits were simple, using one TC, or complex with multiple TCs involved.
-Experiment Results.
-We crafted 16 exploits and 17 interactions following the above methodology.
-For example, the technique T1007 - System Service Discovery’s description reads “Adversaries may try to gather information about registered local system services.
+For example, the technique System Service Discovery (ID T1007)’s description reads “Adversaries may try to gather information about registered local system services.
 Adversaries may obtain information about services using tools and OS utility commands.
@@
 21: tm_send_file 10 ps.log
-Similarly, technique T2014 - Backdoor Installation reads “An attacker can interfere with the hardware or the software, integrating or modifying the existing software, hardware configuration, or the transponder configuration to permit himself future access to the resource”.
-Therefore, we designed the following exploit to upload a malicious script to the satellite software:
-1tm_send_file backdoor.sh 1
-21: obc_system ./recv_files/backdoor.sh
-Due to space limitations, we do not describe all the exploits and interactions here.
-However, the complete exploit and interaction list is available in Table 6 in Appendix A.
-The overall results for this experiment are depicted in Table 3.
+Experiment Results.
+We successfully crafted and ran an interaction or exploit on HoneySat for 33 techniques out of the feasible 38 as depicted in Table III.
+The remaining 5 techniques were not implemented due to limitations in HoneySat’s implementation.
+For example, the Retrieve TT&C master/session keys (T2015.002) technique depends on a cryptographic protocol implementation which HoneySat’s flight software currently does not support.
+This is not surprising, as satellite flight software rarely if ever does [12].
+Due to space limitations, we do not describe all the exploits and interactions here. However, the complete exploit and interaction list is available in Table VIII in Appendix A.
 The key findings of our experiment are shown below, providing strong evidence for answering question Q-1 in the affirmative and design objective DO-1 as achieved.
-Key findings Q-1 HoneySat supports 86.8% of the SPACE-SHIELD matrix techniques possible in a virtual satellite honeypot. HoneySat supports 100% of the SPACE-SHIELD matrix tactics.
+Key findings Q-1 HoneySat supports 87% of the SPACE-SHIELD matrix techniques possible in a virtual satellite honeypot. HoneySat supports 100% of the SPACE-SHIELD matrix tactics.
+TABLE III: Tactics and techniques supported by HoneySat compared to the ones included in the SPACE-SHIELD matrix that are applicable to virtual honeypots.
 Tactics SPACE-SHIELD Techniques (Applicable to Virtual Honeypots) HoneySat Supported Techniques
@@
 Total 38 33
-Table 3: Tactics and techniques supported by HoneySat compared to the ones included in the SPACE-SHIELD matrix that are applicable to virtual honeypots.
-6.3 SmallSat Operators Survey
+VI-C SmallSat Operators Survey
 Evaluating the realism of a satellite honeypot is challenging for two main reasons.
-First, as we discussed in Sec.
-1, satellites, including SmallSats are very diverse.
+First, as discussed in Sec.
+I, satellites, including SmallSats, are very diverse.
 Second, there is no established metric or tool, such as Nmap’s OS detection [70], to quantify the level of realism of our honeypot.
-To evaluate the realism and deception capabilities of HoneySat, we surveyed experienced, real-world SmallSat operators.
+To evaluate the realism and deception capabilities of HoneySat, we surveyed experienced SmallSat operators.
 Because operators interact with real-world SmallSat missions on a daily basis they are experts and thus are the best population to rigorously evaluate HoneySat.
+TABLE IV: Summary of questions in Section VI-C, satellite honeypot operation tasks of the survey.
 Satellite Operation Evaluated Component No. Qts.
-Telecommand Scheduling Mission Control SW 6
-Pass Prediction Ground Station Control SW 6
-Antenna Positioning Ground Station Control SW 6
-EPS Telemetry Satellite Simulator 6
-Temperature Telemetry Satellite Simulator 6
-ADCS Telemetry Satellite Simulator 6
-RGB Camera Telecommand Satellite Simulator 6
-Ping Test Radio Simulator 6
-Table 4: Summary of questions in Section 6.3, satellite honeypot operation tasks of the survey (48 total).
-6.3.1 Survey Structure
+Telecommand Scheduling Mission Control SW 5
+Pass Prediction Ground Station Control SW 5
+Telemetry Download Satellite Simulator 7
+Ping Test Radio Simulator 5
+VI-C1 Survey Structure
 We divided our survey into five sections.
-The first section collected essential background information about the participants, while the second part focused on their professional experience.
-In this second section, participants were asked whether they worked in industry, government, or academia and to self-report their cybersecurity skills as well as satellite-related skills (e.g., satellite hardware engineering).
-The third section of the survey probed participants’ satellite operation experience, including how many missions they had operated and which tools they used.
-The largest section of the survey included 48 questions about satellite honeypot operation tasks.
-Participants watched eight 1–2 minute recordings of HoneySat ’s VNC view (Fig.
-4), each showcasing the ground and satellite personality of a real CSP-based CubeSat mission (based on the SUCHAI-2 CubeSat [34]).
-Every recording highlighted a different feature of HoneySat by replicating a real-world satellite mission operation, as discussed in Section 2.
-After each recording, participants answered questions designed to assess how realistic they found that particular aspect of the honeypot.
-Table 4 lists all of the satellite operations demonstrated in these recordings.
-Finally, once participants had seen multiple elements of our honeypot, the survey concluded with an overall evaluation section.
-Both the honeypot operation tasks and the overall evaluation used 5-point Likert scale [55] questions to measure participants’ positive or negative reactions.
-6.3.2 Participants
-We conducted the survey via Qualtrics and distributed the survey directly to active SmallSat mission operators from previously identified missions.
-In total, we received responses from 14 satellite operators who have between 1 and 5 years of experience operating satellites and have operated between 2 and 7 unique missions.
-In terms of demographics, 21.4% (3/14) of participants were female, 71.4% (10/14) male and 7.1% (1/14) preferred not to answer.
-57.1% (8/14) belonged to the 18-24 age group, 35.7% (5/14) to 25-34 and 7.1% (1/14) to 35-44.
-28.5% (4/14) of participants’ highest level of education was high school, 35.7% (5/14) bachelor’s degree, 14.2% (2/14) master’s degree and 21.4% (3/14) Ph.D.
-In regards to geographic location, 78.5% (11/14) of participants were located in Europe, 14.2% (2/14) in North America and 7.1% (1/14) in South America.
+The first section collected the participants’ background information including demographic data.
+The second section focused on their professional experience, including whether they worked in industry, government, or academia and to self-report their satellite-related and cybersecurity skills (e.g., satellite hardware engineering).
+The third section probed participants’ satellite operation experience, including how many missions they had operated and which tools they used.
+The fourth survey section involved giving participants hands-on access to a live instance of HoneySat.
+In this section, participants were asked to perform 4 hands-on tasks that correspond to specific satellite mission operations using a live instance of HoneySat as depicted in Fig.
+4.
+Each hands-on task was designed to feature different components of HoneySat by replicating a real-world satellite mission operation, as discussed in Section II.
+After participants completed each hands-on task, they answered questions designed to assess how realistic they found different aspects of the honeypot.
+Table IV lists the satellite operations included in the survey.
+Finally, in section five, once participants had interacted with multiple elements of our honeypot via these tasks, the survey concluded with participants answering questions about HoneySat’s overall realism and deception.
+Questions related to both the honeypot operation tasks and the overall evaluation used a 5-point Likert scale [71] to measure the participants’ reactions.
+VI-C2 Participants
+We conducted the survey via Qualtrics and Zoom videoconferencing and distributed the survey directly to experienced SmallSat mission operators from previously identified missions.
+In total, we received responses from 10 satellite operators who have between 1 to 10 years of experience operating satellites and have operated between 1 to 5 unique missions.
+In terms of demographics, 20% (2/10) of the participants were female and 80% (8/10) male.
+30% (3/10) belonged to the 18-24 age group, 40% (4/10) to the 25-34 age group, and 30% (3/10) to the 35-44 age group.
+In regards to geographic location, 40% (4/10) of participants were located in Europe, 20% (2/10) in North America and 40% (4/10) in South America.
 Recruiting participants was challenging due to the rarity and specialized nature of the required expertise.
 Over a span of four months, we reached out to national and international institutions, as well as private corporations involved in operational satellite missions to identify suitable participants.
-6.3.3 Methodology and Key Results
-In the survey, we aimed to evaluate three key aspects of our honeypot.
-First, whether the ground segment simulations are realistic; second, whether the space segment simulations are realistic; and third, whether HoneySat, as a whole, provides a convincing and realistic system.
-Before describing the results it is important to emphasize that the participants were informed that the recordings showed a simulation of a satellite mission and not of a real mission.
+VI-C3 Methodology and Key Results
+In the survey, we evaluated three key aspects of our honeypot.
+First, whether the ground segment simulations are realistic; second, whether the space segment simulations are realistic; and third, whether HoneySat, as a whole, provides a convincing and realistic SmallSat mission simulation.
+Before describing the results, it is important to emphasize that the participants were informed that the system they interacted with was a simulation of a satellite mission and not a real one.
+This is because it was not feasible to synchronize participants’ availability to take the survey with the timing of the real satellite’s pass, which may happen only a few times per day and lasts only a few minutes.
 Ground Segment Realism.
-To understand if HoneySat’s ground segment is realistic we showed participants recordings of different satellite operations (discussed in Sec.
-2.3) performed with HoneySat, to showcase the ground segment components listed in Table 4.
-For example, one of the recordings shows the pass prediction operation which involves calculating when and where a satellite will be visible or within range of a specific ground station.
+To understand if HoneySat’s ground segment is realistic, participants interacted with HoneySat by performing different satellite operations (discussed in Sec.
+II-C), to showcase the ground segment components listed in Table IV.
+One of these tasks is the pass prediction operation which involves calculating when and where a satellite will be within the communication range of a specific ground station.
 This operation is performed using the ground station control software.
-After the participants watched a recording of this operation, we asked them to rate their perceived level of realism.
-85.7% of participants agreed or strongly agreed that the pass prediction operation done in HoneySat resembles that of a real mission.
-Another relevant operation is the telecommand scheduling operation which involves the planning and organization of the commands to be sent to the satellite during a pass.
-42.8% of participants strongly agreed that the telecommand scheduling operation performed in HoneySat resembles that of a real mission mentioning the use of a terminal-based mission control software.
-Interestingly, 35.7% of participants neither agreed nor disagreed, citing that they do not use a command line tool but instead a GUI.
-These results indicate that some of the surveyed operators use a command line-based mission control software while others use a GUI-based confirming the diversity of means of operation in Sec.
-2.3.
-In summary, according to our survey, most participants perceive the ground segment simulated by HoneySat as comparable to that of a real satellite mission.
-While a few participants mentioned that they expected to see a GUI-based mission control software instead of a command line-based one, this is something that we expected, as the means of satellite operations vary greatly among missions.
-Nevertheless, thanks to HoneySat’s modularity HoneySat can be extended to use a GUI-based mission control software such as YAMCS [1].
+After the participants performed this operation on HoneySat, we asked them to rate their perceived level of realism.
+90% of the participants strongly agreed and 10% agreed that the pass prediction operation they performed on HoneySat resembles that of a real mission.
+Another relevant operation is the telecommand scheduling operation which involves the planning and queuing of the commands to be sent to the satellite during a pass.
+Again, after performing this operation on HoneySat, 90% of the participants strongly agreed and 10% agreed that the telecommand scheduling operation performed resembles that of a real mission mentioning the use of a command line interface-based mission control software as a contributing factor.
+In summary, according to these results, the vast majority of participants, who are experienced satellite operators, perceive the ground segment simulated by HoneySat as highly resembling a real satellite mission.
 Space Segment Realism.
-In a similar manner, to understand if HoneySat’s space segment is realistic, we showed participants recordings of different satellite operations that make use of HoneySat’s space segment simulations (discussed in Sec.
-4.4) to showcase multiple components of the space segment listed in Table 4.
-For example, one of the recordings shows the EPS Telemetry operation which involves the collection of data from the Electrical Power Subsystem discussed in Sec.
-2.4.
-This operation is performed using the mission control software to send telecommands and download the latest EPS telemetry.
-After the participants watched a recording of this operation performed by HoneySat, we asked them to rate their perceived level of realism of the telemetry output shown during the operation.
-57.1% of participants agreed or strongly agreed that the telemetry shown in the EPS Telemetry operation resembles that of a real mission mentioning realistic EPS values such as voltage, and temperature.
-14.2% neither agreed nor disagreed mentioning that the EPS temeletry was presented in a different format.
-Finally, 28.5% disagreed or strongly disagreed mentioning that the telemetry included less values than the mission that they had operated.
-These results indicate that the EPS telemetry generate by the Satellite Simulator is considered realistic by the majority of participants.
-We also asked participants to asses the ping test operation.
+Similarly, to understand HoneySat’s space segment realism level, we gave participants hands-on access to a live instance of HoneySat and asked them to perform different operations that make use of HoneySat’s space segment simulations (discussed in Sec.
+IV-D).
+For example in the telemetry download operation, participants issued multiple telecommands to HoneySat’s space segment simulation to download a plethora of telemetry data, including EPS, temperature and ADCS data, which was generated in real time by our Satellite Simulator.
+After the participants performed the telemetry download operation, we asked them to rate their perceived level of realism of the telemetry output shown during the operation.
+90% of the participants agreed or strongly agreed that the telemetry shown during the live telemetry download operation resembles that of a real mission, with the temperature telemetry achieving a 100% strongly agree response rate with some participants mentioning that the temperature values resemble a normal operation and align with the expected temperature values of a satellite.
+For the EPS telemetry, 80% strongly agreed, 10% agreed, and the last 10% neither agreed nor disagreed when asked if they perceived the EPS telemetry as realistic.
+In this case, one participant pointed out that in their specific mission they have more batteries, which is expected as satellite missions are highly diverse as we discussed in Sec.
+II-C.
+These results indicate that the telemetry generated by the Satellite Simulator is considered highly realistic by the vast majority of participants.
+We also asked the participants to assess the ping test operation.
 This operation is a diagnostic procedure used to verify the communication link between a ground station and the satellite.
 It involves sending CSP packets between two CSP nodes, namely the ground segment and the spacecraft.
-After the participants watched a recording of this operation performed on HoneySat, we asked them to rate their perceived level of realism.
-64.2% of participants agreed or strongly agreed that the ping test operation resembles that of a real mission citing the realistic response times.
-21.4% of participants neither agreed nor disagreed stating that in their experience the ping test also downloads additional telemetry.
-Finally, 14.2% disagreed or strongly disagreed noting that in their mission the ping command has a different name.
-The results of the ping test operation indicate that HoneySat’s radio simulator (which controls the communication between the ground and space segments) is considered to realistically simulate the communication of a real satellite mission by the majority of participants.
-A significant key finding of our survey is shown below.
-Overall, the results obtained provide convincing evidence for answering question Q-2 in the affirmative and design objective DO-2 as achieved.
-Key finding Q-2 71.4% of surveyed satellite operators agreed or strongly agreed with the following statement: “If I did not know this was a honeypot simulation of a CubeSat satellite mission, I would have believed it to be an actual CubeSat satellite mission.”
-6.4 Internet Interaction Experiment
-This experiment explores HoneySat’s capabilities to entice external actors in the wild by deploying our honeypot over the Internet.
-To accomplish this, we leveraged the Shodan search engine.
-Shodan is a search engine for Internet-connected devices that scans the whole Internet and indexes exposed servers and their TCP ports.
-Shodan then reads the banner information for each port.
-For example, it gathers the web headers and Telnet login banners [40].
-Additionally, Shodan tags servers as honeypots if the servers have too many open ports [37].
-Experiment Description.
-We deployed and configured five HoneySat instances over the Internet and exposed TCP port 24 for the Telnet server and 80 for the web interface so that anybody can interact with it and allow Shodan to index our server and its banners.
-We deployed four HoneySat instances in total.
-Honeypot
-Source
+After the participants completed this operation’s hands-on task on HoneySat, we asked them to rate their perceived level of realism.
+70% of the participants strongly agreed and 30% agreed that the ping test operation resembles that of a real mission with some participants citing that communication was only possible when the satellite was performing a pass.
+The results of the ping test operation indicate that HoneySat’s radio simulator (which controls the communication between the ground and space segments) is considered to realistically simulate the communication of a real satellite mission by all participants.
+In summary, according to these results, the vast majority of participants, who are experienced satellite operators, perceive the space segment simulated by HoneySat as highly resembling a real satellite mission.
+HoneySat’s Overall Realism.
+Finally, to understand HoneySat’s overall realism, we asked the participants a series of overall evaluation questions after they had interacted with HoneySat.
+For example, we asked participants to evaluate the realism of the telecommands used in all hands-on tasks.
+70% strongly agreed, 20% agreed and 10% neither agreed nor disagreed that HoneySat’s telecommands are realistic.
+Additionally, when asked if they would be able to distinguish between HoneySat’s satellite mission simulation and a real mission, 70% strongly agreed, 20% agreed and 10% neither agreed nor disagreed that they would not be able to and again to emphasize that these are experienced satellite operators.
+Overall, our survey results provide convincing evidence for answering question Q-2 in the affirmative and design objective DO-2 as achieved.
+Key finding Q-2 90% of surveyed experienced satellite operators agreed or strongly agreed with the following statement: “I would not be able to distinguish between HoneySat’s satellite honeypot system from the real CubeSat satellite mission it is based on.”
+VI-D Internet Interaction Experiment
+This experiment explores HoneySat’s capabilities to entice external actors to interact with it by deploying our honeypot over the Internet.
+Experiment Methodology.
+We leveraged HoneySat’s customization and deployability features to deploy five instances of HoneySat over the Internet and exposed the necessary TCP ports used by HoneySat’s exposed network protocols discussed in Sec.
+IV-C, e.g., Telnet’s port 23.
+Initially, we deployed our first HoneySat instance to simulate the SUCHAI-2 SmallSat on the premises of the University of Chile, where the mission’s real ground segment is.
+This on-premises deployment provides the most convincing environment to adversaries.
+After coordinating with the University of Chile’s local stakeholders for three months, we deployed HoneySat on an onsite server using a University of Chile IP address starting on July 2024.
+Although we went to great lengths to make our deployment as realistic as possible, after a few months, our honeypot did not entice any external actors.
+Therefore, in January 2025, we deployed four additional HoneySat instances on cloud servers.
+We configured two instances to simulate NASA’s Advanced Composite Solar Sail System (ACS3) CubeSat [72], and two simulating DLR’s PIXL-1 CubeSat [73].
+We selected these two particular CubeSats because they use the CSP ecosystem (Sec.
+II-E) and are currently in-orbit and operational.
+We ensured these additional deployments were believable by configuring their satellite personalities and giving them public IP addresses in the regions belonging to the satellite mission origin, namely Germany and the United States.
+Table V summarizes all five HoneySat deployment details.
+TABLE V: HoneySat Internet deployments’ details.
+Deployment
+Type
+Satellite
+Personality
+HoneySat IP
+Location
+Duration
+Cloud PIXL-1 Germany 6
+Cloud PIXL-1 Germany 6
+Cloud ACS3 United States 6
+Cloud ACS3 United States 6
+On-prem SUCHAI-2 University of Chile 12
+TABLE VI: Exposed Telnet interactions received.
+Satellite
+Personality
+HoneySat
+IP
+Attacker
 IP
@@
 Time
-Honeypot 1 Egypt 4 2 hr
-Honeypot 2 Sweden (Tor) 9 5 min
-Honeypot 1 France (Tor) 6 4 min
-Honeypot 3 USA 8 3 min
-Table 5: Exposed Telnet interactions received.
-Experiment Results. Our honeypot servers were successfully indexed by Shodan. Both Telnet and web banners were captured, and none of them were tagged as honeypots.
-Our honeypots successfully enticed external actors and captured 4 distinct sessions (show in Table 5) via the exposed Telnet protocol discussed in Sec.
-4.3.
-The exposed Telnet protocol was implemented so that a human had to type “activate” in the terminal before sending commands.
-This feature was implemented to filter out Internet Telnet bots.
-As such, the commands that we describe below were sent by a human and not a crawler or Internet bot.
-We now describe the commands received in the first interaction session shown in Table 5.
-  1. 1.
-    help: Shows the ground software available commands.
-  2. 2.
-    ls: The Linux ls command is not a valid command in the honeypot ground software.
-  3. 3.
-    fp_show: Prints the list of commands in the current flight plan (The flight plan refers to a pre-programmed sequence of tasks, and commands that guide the satellite’s operation).
-  4. 4.
-    com_debug: Prints the current CSP network configuration including the current CSP node, interfaces and routing table.
-The commands fp_show and com_debug are flight software-specific commands which suggests that the adversary was successfully deceived by our honeypot and purposefully sent flight software commands instead of haphazardly sending unrelated commands.
-Another factor that indicates that the adversary was deceived is that they engaged with our honeypot over a period of two hours indicating that they did not immediately identified it as a honeypot.
+ACS3 United States Egypt 3 2 hr
+PIXL-1 Germany Sweden (Tor Node) 6 5 min
+ACS3 United States France (Tor Node) 5 4 min
+ACS3 United States USA 8 3 min
+Experiment Results.
+Three of our honeypots successfully enticed external actors and captured four distinct interaction sessions (show in Table VI) via the exposed Telnet protocol discussed in Sec.
+IV-C.
+The four interaction sessions yielded 22 satellite flight software-specific commands.
+These flight software-specific commands show that the adversaries that interacted with HoneySat were purposefully trying to exploit our honeypot using commands that the flight software recognized, not just haphazardly sending random commands.
+We now describe one of the most complex interaction sessions HoneySat captured and analyze it in terms of the SPACE-SHIELD matrix (discussed in Sec.
+II-F).
+In this session, the adversary targeted one of the ACS3 honeypots located in the United States.
+First, they connected to the exposed Telnet server thus getting access to the ground segment, this corresponds to the Ground Segment Compromise (ID T1584.001) technique.
+Then they issued the “help” and “test” commands which list the available commands supported by the MCS and flight software; these commands correspond to the Gather Victim Mission Information (ID T2002) technique.
+Later, the adversary issued the “tm dump 0x0000”, which writes telemetry data to a file, showing an attempt to extract mission data, which matches the Exfiltration Over TM Channel (ID T2022) technique.
+Next, they sent the “com_ping” command which attempts to initiate an interaction with the target spacecraft, matching the Active Scanning (RF/Optical) (ID T2001) technique.
+The adversary again implemented the Exfiltration Over TM Channel (ID T2022) technique by issuing the “tm_parse_beacon” command which prints beacon telemetry (lightweight periodic telemetry) and the “obc_get_sensor” command which instructs the OBC to print a particular sensor’s data.
+Finally, the adversary attempted to tamper with the spacecraft’s OBC by issuing the “obc_update_status” command.
+This command updates the OBC status variables to any arbitrary value, thus matching the Modification of On Board Control Procedures (ID T2010) technique.
+Overall, this interaction session alone involved a chain of five SPACE-SHIELD matrix techniques.
+However, after analyzing all four interaction sessions, we identified the Spacecraft’s Components Discovery (ID T2034) technique, totaling six SPACE-SHIELD techniques based on real-world data captured by HoneySat after successfully enticing and deceiving human adversaries.
+It is worth noting that satellites and space systems in general are far less common compared to CPS such as ICS and UAVs making these interactions collected by HoneySat all the more valuable.
+Additionally, these interactions were produced by human adversaries and not crawlers.
+We know this because HoneySat’s exposed Telnet was implemented so that a human had to type “activate” before sending commands effectively filtering out Internet Telnet bots.
 In summary, these results provide strong evidence for answering question Q-2 in the affirmative and design objective DO-2 as achieved.
-Key finding Q-2 Human adversaries interacted with three of our HoneySat-powered honeypots during four distinct Telnet sessions resulting in 16 satellite mission-specific commands captured.
-6.5 Case Study: Generic CCSDS Mission Honeypot
+Key findings Q-2 HoneySat deceived human adversaries who sent 22 satellite flight software-specific commands.
+The interactions collected by our HoneySat Internet deployments account for six SPACE-SHIELD matrix techniques.
+VI-E Case Study: Extending HoneySat to CCSDS Ecosystem
+TABLE VII: SmallSat ecosystems (2) and example SmallSats (3) integrated to HoneySat.
+Honeypot
+Framework
+Generic
+SmallSat
+Ground
+Configuration
+Institution
+HoneySat CSP
+ACS3
+NASA
+PIXL-1
+DLR
+SUCHAI-2
+U of Chile
+CCSDS
+PUS
+N/A
 In Sec.
-5.2 we described how we implemented HoneySat using one SmallSat protocol ecosystem, namely, CSP.
-In this case study we are interested in testing the extensibility capabilities of HoneySat to support additional ecosystems (discussed in Sec.
-2.5) by adding a second ecosystem to HoneySat, namely the CCSDS ecosystem.
+V-B we described how we implemented HoneySat using one SmallSat protocol ecosystem, namely, CSP.
+In this case study, we are interested in testing the extensibility capabilities of HoneySat to support additional ecosystems (discussed in Sec.
+II-E) by adding a second ecosystem to HoneySat, namely the CCSDS ecosystem.
 Experiment Description.
@@
 To accomplish this, we leveraged an open-source CCSDS ecosystem-based flight software framework, RACCOON OS [74], and YAMCS, an open-source Mission Control software framework with built-in support for PUS [75].
-Methodology. Building upon our HoneySat framework implementation, as detailed in Sec. 5, we enhanced the system by integrating various components of the RACCOON OS and the YAMCS framework.
-Results.
+Experiment Methodology. Building upon our HoneySat framework implementation, as detailed in Sec. V, we enhanced the system by integrating various components of the RACCOON OS and the YAMCS framework.
+Regarding the ground segment, we implemented the exposed network protocol using YAMCS’ built-in web interface (Fig.
+8).
+For the mission control software, we used YAMCS’s built-in Mission Control Software [75].
+For the radio simulator, we used RACCOON’s communication application.
+For both the ground configuration and the logging repository, we again used YAMCS built-in features.
+For the space segment, we implemented the flight software runtime using the RACCOON framework.
+For the flight software services, we configured the RACCOON framework to connect it to the YAMCS’s MCS on the ground segment.
+The satellite personality and logging repository were based on the existing HoneySat implementations.
+Experiment Results.
 We successfully extended HoneySat to support the CCSDS ecosystem.
-Regarding the ground segment we implemented the exposed network protocol using YAMCS’ built-in web interface (Fig.
-7), for the mission control software we used YAMCS’s built-in Mission Control Software [1], for the radio simulator we used RACCOON’s communication application and for the ground configuration and logging repository we again used YAMCS built-in features.
-Focusing on the space segment, we implemented the satellite flight software runtime using the RACCOON framework, and for the Flight Software Services, we configured the RACCOON framework to connect it to the YAMCS’s MCS on the ground segment.
-The satellite personality and logging repository were based on the existing HoneySat implementations.
-The majority of the effort involved in extending HoneySat to support the CCSDS ecosystem was dedicated to understanding the ecosystem itself, including components such as PUS.
-Additionally, we had to analyze the RACCOON flight software code and the YAMCS framework documentation.
+The implementation was completed by a graduate student with no prior satellite-related in the span of two weeks.
+However, the majority of the effort involved in extending HoneySat to support the CCSDS ecosystem was dedicated to understanding the ecosystem itself, and analyzing the RACCOON flight software code and the YAMCS framework documentation.
 The only extra implementation that was required was the flight software services which we modified using Rust.
-Other than that we reused several elements of HoneySat like the satellite personality and the Satellite Simulator.
+Other than that, we reused several HoneySat modules such as the satellite personality and the Satellite Simulator.
 In summary, these results provide strong evidence for answering question Q-3 in the affirmative and design objective DO-3 as achieved.
 Key findings Q-3 Out of the box HoneySat supports CSP and CCSDS, the two most widely used standard space ecosystems, and was evaluated by simulating 3 real-world SmallSats.
-7 Discussion and Future Work
+VI-F Case Study: Hardware-in-the-loop Experiment
+This experiment is designed to produce evidence of the robustness of HoneySat to support real-world satellite hardware-in-the-loop (HIL) operations.
+Experiment Description.
+In this experiment we integrated HoneySat with an in-orbit, operational SmallSat mission.
+To achieve this, we collaborated with Quantum Galactics GmbH [76], an aerospace company that develops hardware subsystems for SmallSats.
+The experiment involved sending a telecommand from HoneySat’s ground segment simulation which was then sent to the in-orbit SmallSat, which would then process the telecommand and respond with sanitized telemetry.
+Due to confidentiality and security restrictions, we are unable to disclose the name of the in-orbit SmallSat which we refer to as SmallSat X.
+Experiment Methodology.
+In order to achieve the HIL integration, we customized both HoneySat and SmallSat X’s mission.
+For HoneySat, we modified the Satellite Simulator to support SmallSat X’s mission control software (YAMCS).
+Specifically, we deployed a proxy server that connected the Satellite Simulator to the SmallSat X’s mission production YAMCS instance.
+Conversely, the SmallSat X’s environment was customized by Quantum Galactics’ team by deploying a script in their YAMCS instance to complete the integration.
+Once the integration was completed, we coordinated with Quantum Galactics to send a telecommand from HoneySat during one of SmallSat X’s passes.
+Experiment Results.
+The telecommand successfully reached SmallSat X during one of the passes and HoneySat received the appropriate telemetry.
+Due to security restrictions, we are not able to disclose further details on the telecommand and the telemetry received.
+However, the entire communication was initiated by HoneySat which in turn received sanitized telemetry from SmallSat X, effectively closing the loop in the HIL experiment.
+Additionally, Quantum Galactics provided us with a snippet of decoded radio signal data shown in Fig.
+5, which confirms that the experiment was successful.
+In summary, these results highlight HoneySat’s robust extensibility features and provide strong evidence for answering question Q-3 in the affirmative and design objective DO-3 as achieved.
+Figure 5: Waterfall plot that shows the radio signal spectrogram when HoneySat communicated with SmallSat X.
+The signal on top (blue box) belongs to the telecommand sent by HoneySat, and the signal on the bottom (red box) belongs to the telemetry received from the SmallSat X.
+Key findings Q-3 HoneySat was successfully integrated into a satellite hardware-in-the-loop operation and communicated with an in-orbit, operational SmallSat.
+VII Discussion and Future Work
+HoneySat’s Generalizability.
+HoneySat was designed to be extensible and modular as we discussed in Sec.
+IV-A which makes our framework generalizable to additional SmallSat missions.
+We showed this by extending HoneySat to support the two most widely used SmallSat protocol ecosystems, namely CSP and CCSDS (Sec.
+VI-E), two flight software, SUCHAI and RACOON (Sec.
+VI-E), and we customized HoneySat to simulate three different SmallSats (Sec.
+VI-D).
+Additionally, we provide a plethora of customization tools including the Satellite Simulator (Sec IV-D), the satellite personality configuration, the ground configuration, and multiple exposed network protocols (Sec V-B2).
+Finally, in Table IX, we provide examples of different extensibility use cases, including high-level steps and the HoneySat modules involved so that researchers can use and extend HoneySat for their own experiments.
 Limitations.
@@
 Broader Applications of HoneySat.
-While originally designed as a honeypot, our framework offers the flexibility to support a range of applications beyond its initial purpose.
+Originally designed as a honeypot, our framework can potentially support a range of applications beyond its initial purpose.
 One promising use case is the development of digital twins for satellite systems, enabling the simulation of real-world satellite subsystems and communication scenarios.
-Furthermore, HoneySat can be integrated into cyber range environments to enhance training programs and cybersecurity simulation exercises.
-Lastly, HoneySat serves as an educational tool, providing researchers, industry professionals, and security enthusiasts with an opportunity to explore and learn about satellite architecture and operations.
-8 Conclusion
-Although we have yet to witness a Stuxnet-like cyberattack on space systems, security researchers and professionals need to develop effective countermeasures to secure satellites.
-In this paper, we introduced HoneySat, the first satellite honeypot that provides a much-needed alternative source of empirical data on attackers’ TTPs.
-We created a satellite honeypot that realistically simulates an entire satellite system, including its sensors and subsystems.
-We performed several experiments that provide strong evidence that the framework can obtain rich interaction data, can be extended and customized, and simulate a satellite to keep its honeypot nature hidden from potential adversaries.
-Finally, we hope that security researchers and professionals take advantage of HoneySat’s open-source implementation and use it as a foundation not only for satellite honeypot deployments but also for simulation, education, and training applications.
-9 Ethical Considerations
-In this paper we consider the ethical consequences and possible negative outcomes of two main elements of our research.
+Furthermore, HoneySat can be integrated into cyber range environments to enhance cybersecurity training exercises.
+VIII Conclusion
+Although we have yet to witness a Stuxnet-like cyberattack on space systems, security researchers need to develop effective countermeasures to secure satellites.
+In this paper, we introduced HoneySat, the first satellite honeypot that realistically simulates an entire satellite mission and provided strong evidence that shows that our honeypot can obtain real-world interaction data and can be extended and customized with real-world satellite software and hardware.
+Finally, we hope that security researchers use HoneySat’s open-source implementation as a foundation not only for satellite honeypot deployments but also for simulation, education, and training applications.
+IX Ethical Considerations
+In this paper, we consider the ethical consequences and possible negative outcomes of two main elements of our research.
 The satellite operator user study and the multiple deployments of our satellite honeypot exposed to the Internet.
-We now discuss the stakeholders, potential risks and how we mitigated those risks for each of these elements.
-9.1 Satellite Operators User Study
-Our user study adheres to the ethical guidelines outlined in the USENIX Security ‘25 Ethics Policy.
+We now discuss the stakeholders, potential risks, and how we mitigated those risks for each of these elements.
+IX-A Satellite Operators User Study
 Prior to conducting any research, the study plan, survey protocol, and associated materials were submitted for review by our Institutional Review Board (IRB), which evaluated the study’s potential risks to participants.
-The protocol received "exempt status" from the IRB, indicating that the research involves no more than minimal risk.
+The protocol received ”exempt status” from the IRB, indicating that the research involves no more than minimal risk.
 This status reflects our careful evaluation of possible harms, which we discuss further below.
@@
 Risk Mitigation.The main risk for the survey participants is the breach of privacy.
-To mitigate this risk, we ensured that no identifiable data (e.g., names, contact details) was collected from survey participants and the survey responses remain fully anonymous.
-Additionally, the survey itself includes an informed consent section that informs the respondents about the voluntary participation, the survey’s purpose, a description of the procedures, the risks involved, and contact information and the option to opt out of the survey.
+To mitigate this risk, we ensured that no identifiable data (e.g., names, contact details) was collected from survey participants, and the survey responses remain fully anonymous.
+Additionally, the survey itself includes an informed consent section that informs the respondents about the voluntary participation, the survey’s purpose, a description of the procedures, the risks involved, contact information, and the option to opt out of the survey.
 Finally, we provided participants the choice to skip questions using the response options “Prefer not to say” and “I do not know” to applicable questions.
-Due to the aforementioned risk mitigation strategies the stakeholders’ breach of privacy risk is negligible.
-9.2 Honeypot Deployment Experiment
+Due to the aforementioned risk mitigation strategies, the stakeholders’ breach of privacy risk is negligible.
+IX-B Honeypot Deployment Experiment
 Stakeholder Identification.
@@
 The main risk for the infrastructure owners is that one of the external actors will use our honeypot as an stepping stone to breach their production infrastructure.
-For the deployment on the premises of the institutions we worked with the local IT infrastructure administrators to ensure that we took all the necessary precautions to avoid any possibility that the external actors could gain access to the infrastructure via our honeypot.
+For the deployment on the premises of the institutions, we worked with the local IT infrastructure administrators to ensure that we took all the necessary precautions to avoid any possibility that the external actors could gain access to the infrastructure via our honeypot.
 First, the infrastructure owners provided us with a server with a clean OS installation that did not have any production applications or sensitive data.
@@
 Fourth, the VNC portion of the honeypot is configured so that external actors cannot use it to host malicious services by denying traffic forwarding.
-Fifth and final, we deployed our honeypot using two sandboxing layers (Docker containers and Virtual Machines (VMs) as we discussed in Appendix A.3.
-Due to the aforementioned risk mitigation strategies the stakeholders’ breach of privacy risk is negligible.
+Fifth and final, we deployed our honeypot using two sandboxing layers (Docker containers and Virtual Machines (VMs) as we discussed in Appendix A-C.
+Due to the aforementioned risk mitigation strategies, the stakeholders’ breach of privacy risk is negligible.
 The main risk for the external actors is the breach of their privacy.
 However, it is generally accepted that trespassers of a computer system do not have reasonable expectation of privacy [78].
-In order to mitigate this risk our honeypot system gives notice and warning to any user connecting to it indicating that "This computer system is for authorized use only." Additionally, the web services of our honeypot are protected by a username and password.
+In order to mitigate this risk, our honeypot system gives notice and warning to any user connecting to it indicating that ”This computer system is for authorized use only.” Additionally, the web services of our honeypot are protected by a username and password.
 This notice system gives external actors a chance to avoid interaction, which helps reduce the risk.
 Appendix A Appendix
-A.1 Interaction Sequences and Exploits
-Table 6 includes all the interaction sequences and exploits we performed during the experiments in Sec. 6.2.
-Table 6: Tactics, Techniques and Procedures’ Interaction Experimental Exploits.
+A-A Interaction Sequences and Exploits
+Table VIII includes all the interaction sequences and exploits we performed during the experiments in Sec. VI-B.
+TABLE VIII: Tactics, Techniques and Procedures’ Interaction Experimental Exploits.
 Tactic Technique ID Subsystem Exploit
@@
 Send TC to lauch a fork bomb or a reset loop
-1: obc_system :(){ :|:& };:
+1: obc_system :(){ :—:& };:
 1: fp_set_cmd_dt 10 2147483647 10 obc_reset
@@
 Impact Transmitted Data Manipulation T2024 Threat Model Limitation N.A.
-A.2 Satellite Simulator Design Details
+A-B Satellite Simulator Design Details
 Figure 6: High-level architecture of the Satellite Simulator.
@@
 As shown in Fig.
-5, the Satellite Simulator includes five module types: simulation modules (light orange), sensor modules (light blue), subsystem state modules (light green), satellite state modules (light gray) and interface modules (light yellow).
+6, the Satellite Simulator includes five module types: simulation modules (light orange), sensor modules (light blue), subsystem state modules (light green), satellite state modules (light gray) and interface modules (light yellow).
 Simulation Modules.
@@
 The Satellite Simulator includes the orbital, rotation, power system, thermal, magnetic, and payload simulations.
-Due to space restrictions and their complex design, we explain each simulation in Appendix A.2.
+Due to space restrictions and their complex design, we explain each simulation in Appendix A-B.
 Overall, these simulations provide the foundation for any satellite operator to use our framework and implement their honeypot using their satellite flight software.
@@
 These modules are abstractions of the satellite subsystems discussed in Sec.
-2.4.
+II-D.
 A given subsystem state is made up of one or more sensors.
@@
 As depicted in Fig.
-5, the Interface (light yellow) is the module that connects the flight software services with the Satellite Simulator simulations.
+6, the Interface (light yellow) is the module that connects the flight software services with the Satellite Simulator simulations.
 This protocol must implement two basic message types, requests and replies.
@@
 This simulation communicates with the orbital simulation to determine if satellite is in the presence of the sun.
-A.3 Deployment and Security Hardening Implementation
+A-C Deployment and Security Hardening Implementation
 Figure 7: Architecture of dockerized Generic CSP Honeypot
-Figure 7: Screenshot of the YAMCS web interface, displaying a TM packet, returned from the Generic CCSDS Mission Honeypot.
-As we mentioned in Sec. 2.1, high-interaction honeypots such as HoneySat present a high risk of adversary takeover. To mitigate this risk, we implemented HoneySat with two sandboxing layers.
+As we mentioned in Sec. II-A, high-interaction honeypots such as HoneySat present a high risk of adversary takeover. To mitigate this risk, we implemented HoneySat with two sandboxing layers.
 Virtual Machine. VMs provide the highest isolation level among sandboxing techniques. We implemented HoneySat in VM environments to leverage VMs’ robust security.
@@
 Specifically, we created four different containers depicted in Fig.
-6.
+7.
 Containerizing HoneySat provides two benefits.
@@
 Second, it proves a convenient and flexible way to deploy HoneySat.
+Figure 8: Screenshot of the HoneySat YAMCS web interface, displaying a TM packet, returned from the Generic CCSDS Mission Honeypot.
+A-D Extensibility Example Use Cases
+TABLE IX: HoneySat example use cases for further extensibility.
+Extensibility Use Case Relevant HoneySat Module(s) High-level Steps
+Add new sensor Satellite Simulator To add a new sensor you can leverage the existing Sensor Abstract Base Class (ABC) included in the Satellite Simulator source code and write the logic for your new sensor.
+Add new payload Satellite Simulator To add a new payload you can leverage the existing PayloadSimulation Abstract Base Class (ABC) included in the Satellite Simulator source code and write the logic for your new payload including where to get the payload data.
+Integrate new flight software Satellite Simulator, ground configuration, satellite personality, radio simulator, flight software services To integrate a new flight software into HoneySat the first step is to analyze and understand the target flight software architecture and source code even before attempting integration.
+Then the target flight software must be integrated using HoneySat’s flight software services module.
+After integrating the flight software services, the Satellite Simulator can be connected using the Interface Python ABC locanted in the Satellite Simulator source code.
+Add a new exposed network protocol Ground Software, Docker compose A new exposed network protocol can be added by modifying the Docker compose file included with HoneySat.
+The compose file allows for new services and TCP ports to be easily added.

### v3

@@
 Abstract
-Satellites are the backbone of several mission-critical services that enable our modern society to function, for example, GPS.
+Satellites are the backbone of mission-critical services that enable our modern society to function, for example, GPS.
 For years, satellites were assumed to be secure because of their indecipherable architectures and the reliance on security by obscurity.
-However, technological advancements have made these assumptions obsolete, paving the way for potential attacks, and sparking interest in satellite security.
-Unfortunately, to this day, there is no efficient way to collect data on adversarial techniques for satellites, hurting the generation of security intelligence that can lead to the development of effective countermeasures.
-In this paper, we present HoneySat, the first high-interaction satellite honeypot framework, fully capable of convincingly simulating a real-world CubeSat, a type of Small Satellite (SmallSat).
-To provide evidence of HoneySat’s effectiveness, we surveyed experienced SmallSat operators in charge of in-orbit satellites and deployed HoneySat over the Internet to entice adversaries.
-Our results show that 90% of satellite operators agreed that HoneySat provides a realistic and engaging simulation of a SmallSat mission.
-Additionally, HoneySat successfully deceived human adversaries in the wild and collected 22 real-world satellite-specific adversarial interactions.
-Finally, in a major demonstration of HoneySat’s robustness, we collaborated with an aerospace company to perform a hardware-in-the-loop operation that resulted in HoneySat successfully communicating with an in-orbit, operational SmallSat mission.
-publicationid: pubid: Network and Distributed System Security (NDSS) Symposium 2026 23-27 February 2026, San Diego, CA, USA ISBN 979-8-9894372-8-3 https://dx.doi.org/10.14722/ndss.2026.240537 www.ndss-symposium.org
+However, technological advancements have made these assumptions obsolete, paving the way for potential attacks.
+Unfortunately, there is no way to collect data on satellite adversarial techniques, hindering the generation of intelligence that leads to the development of countermeasures.
+In this paper, we present HoneySat, the first high-interaction satellite honeypot framework, capable of convincingly simulating a real-world CubeSat, a type of Small Satellite (SmallSat).
+To provide evidence of HoneySat’s effectiveness, we surveyed SmallSat operators and deployed HoneySat over the Internet.
+Our results show that 90% of satellite operators agreed that HoneySat provides a realistic simulation.
+Additionally, HoneySat successfully deceived adversaries in the wild and collected 22 real-world adversarial interactions.
+Finally, we performed a hardware-in-the-loop operation where HoneySat successfully communicated with an in-orbit, operational SmallSat mission.
+publicationid: pubid: Network and Distributed System Security (NDSS) Symposium 2026 23-27 February 2026, San Diego, CA, USA ISBN 979-8-9919276-8-0 https://dx.doi.org/10.14722/ndss.2026.240537 www.ndss-symposium.org
 I Introduction
 Satellites are complex devices designed to withstand outer space conditions.
-They serve multiple purposes or types of missions that include position and navigation, e.g., the Global Positioning System (GPS) constellation [1], Earth observation, e.g., the Sentinel constellation [2], and broadband internet service, e.g., the Starlink constellation [3].
-In addition, spacecraft can range from being as massive as thousands of kilograms, such as the International Space Station (ISS), to much smaller one-kilogram CubeSats [4].
+They serve multiple purposes or types of missions that include position and navigation, e.g., the Global Positioning System (GPS) constellation [space2023global].
+In addition, spacecraft can range from being as massive as thousands of kilograms, such as the International Space Station (ISS), to much smaller one-kilogram CubeSats [uwe1_wuerzburg_2005].
 As a result, the software and hardware components that make up a specific spacecraft vary greatly.
-Likewise, satellite missions vary a lot depending on the owner’s budget.
-For example, university missions are smaller as they have a limited budget [5].
 A cyberattack on a satellite or satellite constellation (group of satellites) could have disastrous consequences on a global scale, which is difficult to comprehend.
-Such an attack could lead to the cessation of air traffic and widespread communication blackouts [6].
-It could also cause food shortages and the freezing of financial transactions [7].
-Furthermore, such an attack could exacerbate the Kessler Syndrome [8], a scenario in which collisions between satellites and debris in orbit create a cascade effect, generating even more debris, jeopardizing future satellite launches and operations.
+Such an attack could lead to the cessation of air traffic and widespread communication blackouts [holmes2024growing].
+It could also cause food shortages and the freezing of financial transactions [vancamp2022world].
+Furthermore, such an attack could exacerbate the Kessler Syndrome [drmola2018kessler], a scenario in which collisions between satellites and debris in orbit create a cascade effect, generating even more debris, jeopardizing future satellite launches and operations.
 In this increasingly vulnerable environment, the probability of a successful satellite cyberattack continues to rise.
-This is driven by three key trends [6]: first, satellite deployments have increased at an unprecedented pace.
-For instance, while an average of 82 launches took place between 2008 and 2017, as many as 197 launches occurred in 2023 alone, each typically carrying multiple satellites [9].
-This surge is partly fueled by the rise of cheaper commercial off-the-shelf (COTS) components and the availability of more Space Launch Vehicles (SLVs), which makes access to orbit affordable for smaller institutions, such as universities [10].
-Second, ground station technology has become significantly more affordable (and sometimes open source), greatly lowering the communication barrier with satellites [11].
+This is driven by three key trends [holmes2024growing]: first, satellite deployments have increased at an unprecedented pace.
+For instance, while an average of 82 launches took place between 2008 and 2017, as many as 197 launches occurred in 2023 alone, each typically carrying multiple satellites [semanik2023private].
+Second, ground station technology has become significantly more affordable (and sometimes open source), greatly lowering the communication barrier with satellites [california2024earth].
 Thus, a broader range of malicious actors can now communicate with satellites.
-Third, satellite engineers and operators continue to rely on protocol obscurity practices, such as hiding specialized knowledge about the transmission protocol implemented on satellites [12].
-Although satellite security research has gained increased attention [12, 13], the relationship between outer space and cyberspace remains poorly understood [14].
-At the same time, space-focused threat intelligence remains sparse, and our current methods for identifying the tactics, techniques, and procedures (TTPs) used against these critical systems are limited.
-MITRE ATT&CK currently tracks 152 threat groups but shows only one that targets satellites explicitly [15], highlighting a significant data gap.
-Although the volume of reported cyber incidents in the space sector has grown, these reports rarely provide sufficient detail [7].
+Third, satellite engineers and operators continue to rely on protocol obscurity practices, such as hiding specialized knowledge about the transmission protocol implemented on satellites [willbold2023space].
+Although satellite security research has gained increased attention [willbold2023space, scharnowski2023case], the relationship between outer space and cyberspace remains poorly understood [pavur2022building].
+Although the volume of reported cyber incidents in the space sector has grown, these reports rarely provide sufficient detail [vancamp2022world].
 As a result, the security community has limited visibility into adversarial activity aimed at space infrastructure.
-A honeypot’s ability to collect real-world cyberattack data makes it an ideal solution to this problem [16].
-A honeypot is a decoy computer system intended to lure and entice malicious actors to interact with it [17]; all the while, the honeypot logs all the interactions the attackers make.
-This log data can later be analyzed to discover new and existing TTPs.
-Since the release of the first honeypot, the Deception Toolkit, in 1997 [18], a wide range of honeypots with ever-increasing capabilities have been introduced.
-These honeypots are used by universities, companies, and nation-states worldwide [19, 20, 21, 22, 23].
-Honeypots are also used to deter malicious actors from attacking different types of systems, from industrial control systems [23] to social media platforms [24].
-However, as of the time of writing, there is no space-sector specific honeypot in the literature.
-In this paper, we design, implement, and evaluate the first satellite honeypot, HoneySat, to attract and analyze adversaries who attack space infrastructures over the Internet, a commonly observed threat vector [25, 26, 27, 28].
-HoneySat is a modular, high-interaction honeypot framework that realistically simulates a complete satellite system (ground infrastructure and satellite).
-Specifically, HoneySat simulates Small Satellites or SmallSats, which are spacecraft with a mass of less than 180 kilograms [29].
-CubeSats, for example, are SmallSats.
-Additionally, as part of HoneySat, we developed the Satellite Simulator, a Python project to provide generic simulation functionality for users to populate satellite honeypots with believable data.
-This capability enables the creation of complete testing environments for satellite software that integrate a variety of simulated subsystems and sensors.
-For instance, Satellite Simulator can simulate orbital mechanics (e.g., position tracking, attitude adjustment) and electrical power systems (e.g., power generation, consumption, and distribution) alongside other subsystems.
 We leveraged our framework to create honeypots of real-world CubeSat missions.
@@
 Our framework simulates an entire satellite mission, provides realistic telemetry, and supports real telecommands.
-Moreover, in a significant showcase of HoneySat’s practical relevance, a leading multinational space agency has integrated HoneySat into operational test environments to evaluate its effectiveness in securing mission-critical space systems.
 Finally, we collaborated with an aerospace company to perform a hardware-in-the-loop experiment where HoneySat successfully communicated with an in-orbit, operational SmallSat.
@@
 VI-F).
+In the spirit of open and reproducible science, HoneySat’s source code and experimental results are available online [^1] https://doi.org/10.5281/zenodo.17871431, [^2] https://github.com/HoneySat.
 II Background
@@
 These honeypots offer minimal interaction, simulating real systems through scripts.
-They are easy to setup and maintain due to low resource consumption, and have a reduced risk of adversarial takeover.
-However, they provide limited interaction opportunities to adversaries, which limits the data they provide.
-Low-interaction honeypot examples include Conpot [30] and Honeyd [31].
+They are easy to setup and maintain, and have a reduced risk of adversarial takeover.
+However, they provide limited interaction opportunities, which limits the data they provide.
+Some examples include Conpot [vestegaard2014conpot] and Honeyd [provos2007virtual].
 High-Interaction Honeypots.
-These honeypots offer extensive interaction opportunities via emulation or advanced simulations [31].
-Their main advantage is providing adversaries with almost limitless interactions, providing extensive interaction data.
-However, they pose a high risk of adversarial takeover as adversaries have more opportunities to hijack the honeypot [32].
-High-interaction honeypot examples include Cowrie [33] and HoneyPLC [23].
+These honeypots offer extensive interaction opportunities via emulation or simulation [provos2007virtual].
+Their main advantage is providing adversaries with almost limitless interactions.
+However, they pose a high takeover risk [marra2024feasibility].
+Examples include Cowrie [oosterhof2024cowrie] and HoneyPLC [lopez2022honeyplc].
 II-B Honeypot’s State of the Art
-The literature on honeypots includes hundreds of implementations that simulate a diverse set of systems [34, 35].
-From implementations that simulate a host’s TCP/IP stack, such as Honeyd [31], to modern approaches that integrate social media applications, such as HoneyTweet [24].
+The literature on honeypots includes hundreds of implementations that simulate a diverse set of systems [ilg2023survey, nazario2024awesome].
+From implementations that simulate a host’s TCP/IP stack, such as Honeyd [provos2007virtual], to modern approaches that integrate social media applications, such as HoneyTweet [acharya2024conning].
 In the absence of a satellite honeypot, we examine the approaches most related to satellites: Industrial Control Systems (ICS) and Unmanned Aerial Vehicles (UAV), summarized in Table I.
 Satellite systems like ICS must be aware of some physical process, e.g., the sun’s position, via sensors to acquire data about the physical world.
-Several ICS honeypots have simulated these physical processes.
-For example, ICSnet [36] and HoneyICS [37] simulate components such as programmable logic controllers (PLCs) and actuators such as valves.
-HoneyDrone [38] is a UAV honeypot that integrates different simulations, e.g., Ardupilot, to recreate multiple attack scenarios.
-Although UAV honeypots share some similarities with satellites [39], satellite honeypots require additional physical simulations and involve a more complex operation environment, which we discuss in Sec.
-II-C.
+ICS honeypots, such as HoneyICS [lucchese2023honeyics], simulate these physical processes.
+In addition to ICS, UAV honeypots such as HoneyDrone [daubert2018honeydrone] achieve this via simulations, and recreate attack scenarios.
 TABLE I: Comparison of Existing Honeypots and HoneySat.
@@
 Protocols
-Physics
-Sims
-Extensi-
-bility
-Conpot [30]
+Physics Sims
+Extensibility
+Conpot [vestegaard2014conpot]
 Low 9 0
-HoneyPLC[23]
+HoneyPLC[lopez2022honeyplc]
 High 3 0
-ICSPot [40]
+ICSPot [conti2022icspot]
 High 4 1 $\times$
-HoneyICS[37]
+HoneyICS[lucchese2023honeyics]
 High 2 1
-HoneyDrone [38]
+HoneyDrone [daubert2018honeydrone]
 Medium 4 1 $\times$
@@
 The Ground Segment (GS) covers the terrestrial infrastructure required for a successful satellite operation.
-It consists of a ground station, responsible for exchanging data with the spacecraft, the computational and network infrastructure required for communication, and the systems to operate the satellites, e.g., servers [12].
+It consists of a ground station, responsible for exchanging data with the spacecraft, the computational and network infrastructure required for communication, and the systems to operate the satellites, e.g., servers [willbold2023space].
 The GS includes the Ground Segment Software (GSS) that helps operators schedule and send commands and visualize data.
@@
 A satellite is launched into orbit and then establishes communications with the ground segment.
-During regular operations, satellites may communicate through one or multiple ground stations [12].
+During regular operations, satellites may communicate through one or multiple ground stations [willbold2023space].
 Telecommands (TC) 3 and Telemetry (TM) 4.
-The basic data flow between the space and ground segments are TC and TM [41].
-TM is the data the satellite sends to the ground station which may contain the satellite’s status or payload data [41].
+The basic data flow between the space and ground segments are TC and TM [esa2013telemetry].
+TM is the data the satellite sends to the ground station which may contain the satellite’s status or payload data [esa2013telemetry].
 TCs are used to operate the satellite and are transmitted and encapsulated in a space protocol (see Sec.
@@
 The design and implementation of TCs varies depending on the satellite mission.
-From a security perspective, TCs are particularly important as an attacker that can send valid TCs to a satellite can fully take over the mission [12].
+From a security perspective, TCs are particularly important as an attacker that can send valid TCs to a satellite can fully take over the mission [willbold2023space].
 Orbital Pass 5.
 A satellite and its ground station can communicate only during an orbital pass.
-An orbital pass, or more commonly a pass, is when the satellite rises above a ground station’s horizon and becomes available for communication.
-A pass’s duration and timing depend on the satellite’s orbit characteristics and any obstructing objects between the satellite and the ground station, e.g., mountains [42].
-In some cases, such as geostationary orbit (GEO), the satellite can be above the horizon constantly.
-The occurrence of passes can be calculated using the spacecraft’s orbital elements in the two-line element (TLE) data [43].
+An orbital pass, or simply a pass, is when the satellite rises above a ground station’s horizon and becomes available for communication.
+A pass’s duration and timing depend on the satellite’s orbit characteristics and any obstructing objects, e.g., mountains [wood2006introduction].
+Passes can be predicted using two-line element (TLE) data [vallado2012two].
 Satellite Mission Operations 6.
-Satellite mission operations vary widely depending on the owning organization’s budget, and technology [5].
+Satellite mission operations vary widely depending on the owning organization’s budget, and technology [maya1_inquirer_2018].
 Nevertheless, they share some commonalities, which we now describe.
-A mission’s operation involves a team of operators that use ground segment software (GSS) to operate a satellite(s) and ensure the mission’s success [44].
+A mission’s operation involves a team of operators that use ground segment software (GSS) to operate a satellite(s) and ensure the mission’s success [nasa2024flight].
 Operations are carried out in a Mission Operations Center (MOC), where operators sit at their workstations to manage TCs sent to the satellite(s).
-Operators may also remotely operate satellites 7 by connecting to the ground segment using tools such as VNC (Virtual Network Computing) [11].
+Operators may also remotely operate satellites 7 by connecting to the ground segment using tools such as VNC (Virtual Network Computing) [california2024earth].
 Satellite operations include two main activities: satellite tracking and TC generation and scheduling.
@@
 TC generation and scheduling crafts commands to be sent to the satellite to perform different functions, e.g., download payload data.
-Figure 1: The components commonly found in a satellite mission in the context of the space and ground segments.
+Figure 1: The components commonly found in a satellite mission.
 Ground (Segment) Software (GSS).
@@
 GSS are very diverse.
-Some satellite missions develop their own GSS while others use open-source [45] or proprietary GSS [46].
+Some satellite missions develop their own GSS while others use open-source [suchai-flight-software] or proprietary GSS [gomspace2021nanocom].
 There are two main types of GSS: Mission Control Software (MCS) and Ground Station Control Software (GSCS).
 Mission control software 8 manages TCs and scripts to be sent to the satellite and display TM.
-For example, ESA’s SCOS-2000 is an MCS that provides generic functionality that can be customized for a specific mission operation [47].
+For example, ESA’s SCOS-2000 is an MCS that provides generic functionality that can be customized for a specific mission operation [merri2007cutting].
 Some missions develop their own MCS.
-For example, the SUCHAI mission developed their own MCS application [45].
+For example, the SUCHAI mission developed their own MCS application [suchai-flight-software].
 Ground station control software (GSCS) 9 helps satellite operators track and visualize the satellite’s orbit and provide information about each satellite’s pass.
-For example, Gpredict is a popular open-source GSCS that performs real-time satellite tracking and orbit prediction [48].
+For example, Gpredict is a popular open-source GSCS that performs real-time satellite tracking and orbit prediction [csete2023gpredict].
 II-D Satellite Architecture
-Satellite architectures are varied and complex [49]; however, here we describe the most common terminology depicted in Fig. 1’s space segment 2.
+Satellite architectures are varied and complex [ivancic2003architecture]; however, here we describe the most common terminology depicted in Fig. 1’s space segment 2.
 When referring to satellite architecture there is a distinction between the platform that facilitates the successful operation of the satellite’s activities and the payload.
-The platform provides the possibility to run a payload that fulfills the purpose of these missions.
-Payloads are different depending on the satellite’s mission and can range from instruments to measure physical properties to communications systems providing wireless connectivity.
+The platform underpins the payload that fulfills the mission’s purpose.
+Payloads differ depending on the satellite’s mission and can range from measurement instruments to communications systems.
 Platform: The platform is composed of custom-designed or off-the-shelf subsystems necessary for critical satellite operations.
 These include the Attitude Determination and Control System (ADCS) to maintain the satellite’s orientation (i.e.
-attitude) and position to keep the satellite pointed towards antennas or solar panels illuminated; the Electrical Power Subsystem (EPS) for managing power generation and distribution; the Communication Subsystem (COMM) and the Command and Data Handling (C&DH) subsystems to facilitate communications for receiving TC and sending TM and controlling the satellite operations.
+attitude); the Electrical Power Subsystem (EPS) for managing power generation and distribution; the Communication Subsystem (COMM) and the Command and Data Handling (C&DH) subsystems to facilitate communications for receiving TC and sending TM.
 These subsystems are controlled via TCs sent to the satellite.
-TCs may be interpreted by a central C&DH System or merely forwarded to the recipient subsystem for further processing [50].
-This managerial duty is handled by the Flight Software (FS) running on an embedded system.
-Some examples include NASA’s Core Flight System (cFS) [51] and the University of Chile’s SUCHAI FS [52].
-Attackers aim to gain the ability to send TCs to the flight software as this may grant control over all spacecraft subsystems.
+TCs may be interpreted by a central C&DH System or merely forwarded to the recipient subsystem [yost2023nasa].
+This managerial duty is handled by the Flight Software (FS) running on an embedded system, for example, NASA’s Core Flight System (cFS) [mccomas2016core].
+Attackers aim to gain the ability to send TCs to the flight software as this may grant control over all subsystems.
 Payload.
 The payload is the equipment that a satellite employs to fulfill its mission.
-Due to satellites’ varied missions, payloads are heavily customized [53].
-For example, if a satellite’s mission is remote sensing, its payload may include an infrared camera [54].
+Due to satellites’ varied missions, payloads are heavily customized [esa2018payload].
+For example, if a satellite’s mission is remote sensing, its payload may include an infrared camera [nasa2019remote].
 II-E Small Satellite Protocol Ecosystems
@@
 Cubesat Space Protocol.
-The Cubesat Space Protocol (CSP) family of protocols is a one-stop solution for SmallSat missions [55].
-CSP is implemented as an open-source C library called libCSP [55] and follows the TCP/IP model, including transport and routing protocols and multiple layer 2 interfaces such as I2C (Inter-Integrated Circuit), and ZeroMQ (ZMQ) for transmission on TCP/IP networks [56].
+The Cubesat Space Protocol (CSP) family of protocols is a one-stop solution for SmallSat missions [LibcspLibcsp2024].
+CSP is implemented as an open-source C library called libCSP [LibcspLibcsp2024] and follows the TCP/IP model, including transport and routing protocols and multiple layer 2 interfaces such as I2C (Inter-Integrated Circuit), and ZeroMQ (ZMQ) for transmission on TCP/IP networks [GetStarted].
 In the context of a satellite mission, CSP connects the ground segment and satellite subsystems as part of a CSP network where each subsystem is identified as a node.
 Sending a TC is as simple as sending a CSP packet with the address of the corresponding subsystem node.
-In case the satellite is not currently performing a pass, the packet is dropped.
-There are also default services running on standardized ports that are enabled by default when building libCSP [55], for example for network diagnostic (ping) or basic operations (rebooting).
-The straightforwardness of CSP makes it particularly relevant for the honeypot use case as the internal structures of any CSP-based mission will look similar, so distinguishing two missions may be difficult.
 CCSDS Space Communication Protocols.
-The Consultative Committee for Space Data Systems (CCSDS) Space Communication protocols are a set of standardized protocols used for different purposes in space communications [57].
+The Consultative Committee for Space Data Systems (CCSDS) Space Communication is a set of standardized protocols used for different purposes in space communications [pusecosysytem].
 A relevant CCSDS protocol for SmallSat TM and TC is the Space Packet protocol.
-This protocol is used in combination with the ECSS Packet Utilization Standard (PUS) [58] to define how TCs and TM are encoded and transported.
-PUS defines services (and thus sets of TC/TM) for functionality that satellite missions require, including large data transfer or event reporting [57].
-Proprietary Protocols.
-In addition to CSP and CCSDS, vendor-specific protocols may be used on top of more standard protocols.
-Some basic small satellites may have entirely ad-hoc protocols used on a single mission, for example the ESTCube-1 uses a custom protocol [59].
-These proprietary protocols are not ideal candidates for constructing honeypots as the ecosystem of one mission will look out of place for other missions, preventing generalization.
-The choice of ecosystem also influences the choice of GSS as an MCS usually supports a single protocol stack [45].
-In addition to the space protocol ecosystems above, satellite operators use well-known network protocols to connect with the mission’s ground segment services.
-These protocols include Telnet, SSH, VNC, FTP, and HTTP.
-Telnet, SSH, and VNC are used to remotely connect to the MOC workstations [11].
-HTTP is used for web interfaces that operators use to visualize mission data.
-FTP is used to download TM data from the mission data repository.
-If an adversary were to compromise any of these services, they could use it as a stepping stone to target the satellite itself.
+This protocol is used in combination with the ECSS Packet Utilization Standard (PUS) [kaufeler1994esa] to define how TCs and TM are encoded and transported.
+PUS defines services (and thus sets of TC/TM) for functionality that satellite missions require, including large data transfer or event reporting [pusecosysytem].
 II-F Space Systems’ Tactics, Techniques and Procedures (TTPs)
-Tactics, Techniques, and Procedures (TTPs) describe a malicious actor’s behavior in a structured scheme to understand how they might execute an attack [60, 61].
-Several frameworks have been introduced to standardize space systems’ TTPs.
-There are two MITRE-style frameworks, the SPARTA matrix [62] and the SPACE-SHIELD matrix [63].
-These matrices list and describe space security-specific tactics and techniques such as initial access and ground segment compromise.
-For the remainder of this paper we use the SPACE-SHIELD matrix to analyze HoneySat.
+Tactics, Techniques, and Procedures (TTPs) describe a malicious actor’s behavior in a structured scheme to understand how they might execute an attack [raza2024what, nist2024tactics].
+The SPACE-SHIELD matrix [esa2024space-shield] is a framework used to standardize space systems’ TTPs, for example, ground segment compromise.
 III Threat Model
@@
 To gain initial access to the ground segment, an adversary needs to connect via one of the exposed network protocols discussed in Sec.
-II-E, namely, VNC, Telnet, TCP/IP, etc., which correspond to the operational protocols used in real satellite missions.
+II-E, which correspond to the operational protocols used in real satellite missions.
 From there, an adversary may try to launch different commands to take full control and/or compromise the services offered by the satellite’s mission as depicted in Fig.
-3.
+4.
 Finally, in this paper, the modeling of physical radio communication between the space and ground segments, which is commonly used in practice, is considered out of scope and left for future work.
@@
 As explained in Sec.
-I, the number one objective of any honeypot is to capture interaction data from which we can derive knowledge on adversaries’ TTPs.
-As such, our first objective for HoneySat is to have the capability to capture rich interaction data.
+I, the number one objective of any honeypot is to capture interaction data from which we derive knowledge on adversaries’ TTPs.
+As such, HoneySat’s first objective is to capture rich interaction data.
   2. DO-2
@@
 For example, a particular SmallSat may use CSP or CCSDS ecosystems.
-Additionally, each SmallSat mission has a particular orbit that must be customizable.
-As such, our third objective is for HoneySat to provide extensibility and customizability.
+As such, HoneySat’s third objective is extensibility and customizability.
 Figure 2: HoneySat framework high-level architecture design.
@@
     Realistic Simulation.
-As we discussed in Sec.
-II-B, satellites keep track of their orbit location, the position of the sun, among other details.
-These details must be simulated by our honeypot; otherwise, they may alert adversaries that they are interacting with a fake system.
-For this reason, we selected realistic simulation as our second design principle.
+As discussed in Sec.
+II-B, satellites track their orbit location, among others.
+These details must be simulated; otherwise, they may alert adversaries that they are interacting with a fake system.
 This design principle is based on design objective DO-2.
@@
 II, satellites are complex and diverse systems.
-There is no one-size-fits-all solution.
 To tackle this problem, we designed HoneySat to be modular.
-Modularity allows us to insert, remove, or change any part of our honeypot framework without implementing a new honeypot from scratch.
 This design principle is based on our objective DO-3.
@@
 2, consists of two sets of simulations, the space segment simulations and the ground segment simulations.
-Table II illustrates how HoneySat’s simulation components match the components of a real satellite mission.
-TABLE II: Correspondence between real satellite mission components and HoneySat’s design components.
-Segment Real Component HoneySat Component
-Ground Remote Operation Protocols VNC, Telnet
-Ground Mission Control Software Modified MCS
-Ground Ground Station Control Software Gpredict
-Ground Web Interface Web Interface
-Ground Ground Station Radio Simulator
-Space Flight Software F.S. Runtime + Services
-Space On Board Computer Docker Container
-Space Platform Satellite Simulator
-Space Payload Satellite Simulator
+Table VII illustrates how HoneySat’s simulation components match the components of a real satellite mission.
 IV-C Ground Segment Design
@@
 II-C, the ground software includes mission control software (MCS) and ground station control software (GSCS).
-We leverage existing ground software, allowing HoneySat to provide a high-interaction simulation; thereby following design principle DP-1.
+We leverage existing ground software, e.g., Gpredict, allowing HoneySat to provide a high-interaction simulation following design principle DP-1.
 3) Radio Simulator.
@@
 This produces a high-interaction honeypot simulation that follows design principle DP-1.
+Figure 3: High-level architecture of the Satellite Simulator. Due to space limitations, we do not show all the available sensors and simulations.
 2) Satellite Simulator.
 One of the biggest challenges when designing HoneySat was simulating all the physical processes, e.g., attitude, that satellites need to know.
-The Satellite Simulator is designed to solve this problem by simulating all the necessary satellite subsystems, e.g., the Electrical Power System, and the sensors, e.g., a GPS receiver, that a satellite requires to function.
-The purpose of the Satellite Simulator is to abstract away the various hardware peripherals that communicate with the flight software by simulating these peripherals’ data.
-For example, suppose that a particular flight software was implemented to utilize a specific lithium-ion battery pack.
-In that case, the Satellite Simulator’s power system simulation can simulate the basic behavior of the battery.
-We designed the Satellite Simulator to be modular so that new subsystems and sensors can be easily added, modified, or disabled according to our design principle DP-3.
+The Satellite Simulator solves this problem by simulating all the necessary satellite subsystems, e.g., the Electrical Power System, and the sensors, e.g., a GPS receiver.
+As shown in Fig.
+3, the Satellite Simulator includes five module types: simulation modules (orange), sensor modules (blue), subsystem state modules (green), satellite state modules (gray) and interface modules (yellow).
+Simulation Modules.
+These modules are abstractions of real-world processes required for a realistic satellite simulation (DP-2), e.g., the orbital simulation which uses orbital mechanics to calculate data like the satellite’s orbital position.
+These simulation modules can communicate with each other to facilitate proper functionality.
+For example, the power simulation queries the orbital simulation for data to determine if the satellite is positioned properly to draw power from its solar panels.
+The Satellite Simulator includes six simulations, the orbital, rotation, power system, thermal, magnetic, and payload simulations.
+  • Orbital Simulation.
+The orbital simulation calculates the satellite’s orbit [vallado2012two].
+It uses the TLE data configured in the satellite personality and it generates multiple values such as, latitude, and longitude of the satellite.
+The payload, EPS, and magnetic simulations rely on data from this simulation.
+  • Rotation Simulation.
+This simulation calculates the satellite’s attitude changes to provide its orientation by using a reference frame fixed to the satellite body and a non-rotational reference frame.
+The relation between the two reference frames is calculated using the conservation of angular momentum and the rigid-body Euler equation [euler].
+This simulation’s data is used by payloads like the Red, Green, Blue (RGB) camera to point towards Earth’s surface.
+  • Power System Simulation.
+This simulation manages the satellite’s power collection, consumption, and distribution.
+It tracks the battery capacity, the power draw and simulates battery charging whenever the orbital simulation tells it that the satellite is exposed to sunlight.
+  • Thermal Simulation.
+This simulation calculates the satellite’s temperature based on the total thermal energy and a user-defined specific heat capacity and uses a radiation loss formula [stefan-boltzmann] to calculate the thermal radiation emission.
+  • Magnetic Simulation.
+This simulation tracks and analyzes the interactions between the Earth’s magnetic field and the satellite’s own magnetic environment.
+It communicates with the orbital and rotation simulations to determine the satellite’s position and orientation to calculate the Earth’s magnetic field components.
+  • RGB Camera Simulation.
+This simulation replicates the functionality of an RGB camera pointed at Earth.
+It uses Earth observation satellite imagery from the U.S.
+Geological Survey (USGS) [geological2022earthexplorer].
+If a capture image command is issued, it will select and return an image.
+This simulation communicates with the orbital simulation to determine if satellite is in the presence of the sun.
+Sensor Modules.
+These modules are abstractions of the hardware sensors used by a satellite, i.e., temperature sensor.
+The sensor modules do not perform any computations, instead they collect data directly from the simulation modules following DP-2.
+For example, the voltage sensor will query the power simulation to collect data about the current state of the battery.
+Subsystem State Modules.
+These modules are abstractions of the satellite subsystems discussed in Sec.
+II-D.
+A given subsystem state is made up of one or more sensors.
+For example, in Fig.
+3 the Payload State (light green) includes the Camera Sensor (light blue).
+In this way, subsystem states serve two purposes.
+They can get data from their sensors or can set a specific configuration on their sensors.
+Satellite State Module.
+This module is an abstraction of an entire satellite.
+As depicted in Fig.
+3 the Satellite State (light gray) is made up of multiple subsystem states.
+Satellite State module routes messages between the Satellite Simulator and the modified flight software.
+For example, if the modified flight software sends a message to the Satellite Simulator requesting to provide the present voltage in the EPS, the satellite state will route that message to the EPS State.
+Interface Module.
+This module is an abstraction of the communication protocol between the modified flight software and the Satellite Simulator.
+As depicted in Fig.
+3, the Interface (light yellow) is the module that connects the flight software services with the Satellite Simulator simulations.
+This protocol must implement two basic message types, requests and replies.
+However, to meet DO-3, the underlying implementation of these messages is left open for the user to decide based on their requirements.
 3) Flight Software Services.
@@
 In this section, we provide a brief overview of how the designed framework components shown in Figure 2 interact by following the numbers in Fig. 4.
-An adversary gets initial access via one of the Exposed Network Protocols ①.
-On interacting with the Simulated Ground Segment, an attacker is presented with access to the Ground Software ②.
+An adversary gets initial access via one of the Exposed Network Protocols 1.
+On interacting with the Simulated Ground Segment, an attacker is presented with access to the Ground Software 2.
 The configuration of this ground software is defined by the Ground Configuration, which allows the ground software to look like one of many different missions from its protocol ecosystem to the attacker.
@@
 They might want to try to gain more privileges on the ground segment or try to send TC to the space segment.
-Instead of deploying a ground station with an RF transmitter and associated hardware, we employ the Radio Simulator ③ to handle all simulated radio frequency communications.
+Instead of deploying a ground station with an RF transmitter and associated hardware, we employ the Radio Simulator 3 to handle all simulated radio frequency communications.
 When an attacker uses the ground segment to send a valid TC or raw packets, the Radio Simulator checks whether the satellite configured in the Satellite Personality is currently passing over the designated ground station location.
-If so, the Radio Simulator forwards the TC to the Flight Software Runtime ④.
-When the Flight Software Runtime receives a command, it will run the corresponding Flight Software Service to compute a response ⑤.
-In case it requires either changes to an on-board system’s state or a sensor value to execute the TC, it will invoke the Satellite Simulator ⑥.
+If so, the Radio Simulator forwards the TC to the Flight Software Runtime 4.
+When the Flight Software Runtime receives a command, it will run the corresponding Flight Software Service to compute a response 5.
+In case it requires either changes to an on-board system’s state or a sensor value to execute the TC, it will invoke the Satellite Simulator 6.
 The Satellite Personality contains the configuration used by the Satellite Simulator.
-Once a TC is processed, the resulting TM is routed back through the Radio Simulator to either the attacker or the ground software used by the attacker ⑦-⑫, and is also recorded in our Logging Repository.
+Once a TC is processed, the resulting TM is routed back through the Radio Simulator to either the attacker or the ground software used by the attacker 7- 12, and is also recorded in our Logging Repository.
 V HoneySat’s Implementation
@@
 V-A Satellite Simulator Implementation
-We implemented the Satellite Simulator as a Python object-oriented programming application.
-We implemented 11 sensors, 4 subsystems, 1 satellite state, and 1 interface, based on our 6 simulations.
-These modules are generic enough so that we do not have to heavily modified them to support other satellites.
-Overall, satellite architectures, subsystems, and sensors are varied, so modularity is necessary when constructing a simulation framework.
-Many of the basic simulations and corresponding sensors can be shared, especially in a honeypot where attackers lack detailed information that would enable them to distinguish between small differences in sensor values.
-Further details
-can be found in Appendix A-B.
+We implemented the Satellite Simulator as a Python object-oriented programming application. We implemented 11 sensors, 4 subsystems, 1 satellite state, and 1 interface, based on our 6 simulations.
 V-B Generic CSP Mission Honeypot
-Following our design, we implemented HoneySat to support the CSP ecosystem.
-This implementation provides a generic CSP-based mission honeypot that can be customized to simulate a specific CPS-based satellite mission.
+We implemented HoneySat to support the CSP ecosystem providing a generic CSP-based honeypot that can simulate any specific CSP-based mission.
 V-B1 Ground Segment Implementation
@@
 1) Exposed Network Protocols.
-We implemented four exposed entry points.
-A VNC server, a Telnet server, a Web server, and ZeroMQ-based access to the CSP Network.
-We implemented the VNC server using TigerVNC [64] and used PyZMQ [65] for the Raw Packet Access.
-The Telnet server was implemented using Python to expose the command-line interface of the Ground Segment Software (GSS) on the network.
+We implemented a VNC server, a Telnet server, a Web server, and ZeroMQ-based access to the CSP Network.
+We implemented the VNC server using TigerVNC [ossman2024tigervnc] and used PyZMQ [granger2024pyzmq] for the raw packet access.
+The Telnet server was implemented using Python to expose the command-line interface of the Ground Segment Software.
 The web interface allows for the customization of text presented to the attacker through configuration.
-Using the above entry points, adversaries can interact with the GSS and the space segment simulations.
 Figure 5: The view an attacker would see upon connection to the VNC server. Ground Station Control Software (Red) Ecosystem/Mission Specific MCS (Yellow).
 2) Ground Software.
-To implement the MCS software for our CSP-based honeypot, we require an MCS that supports the CSP ecosystem.
-We leveraged the SUCHAI mission control software (SUCHAI MCS).
+To implement the MCS we leveraged the SUCHAI mission control software (SUCHAI MCS).
 This software is a node on the CSP network that has a command line-based interface that allows basic MCS functionality like scheduling telecommands or saving downlinked telemetry.
-This is a good choice for the generic CSP MCS as it can be passed off as the CLI of commercial ground segment appliances like the GomSpace NanoCom MS100 [46].
-Given HoneySat being a generic honeypot, we can set mission-specific parameters through the ground configuration.
-Additionally, we provide a desktop environment accessible via VNC, complete with supporting applications such as a configured Gpredict for the satellite mission.
-For certain deployments, this environment also includes a mission-specific GUI tool, both of which appear as open windows on the desktop.
+Additionally, we provide a desktop environment accessible via VNC that includes Gpredict and the SUCHAI MCS.
+Finally, we implemented a radio link and network delay simulation using a sleep function whenever the MCS receives data from the space segment simulations.
 3) Radio Simulator.
-The Radio Simulator was implemented as a ZMQ Hub [66].
+The Radio Simulator was implemented as a ZMQ Hub [ProtocolStackCubesatZMQ].
 The ZMQ Hub functions as a router for the CSP network simulating both the ground part of the CSP network including the ground station and the space segment.
 Depending on whether the satellite is currently passing, packets to the space segment are forwarded to the Flight Software Runtime.
-We decided to use ZeroMQ as it is the protocol to use for transmitting CSP packets over TCP/IP [66].
-4) Ground Configuration. We implemented the ground configuration as settings that can be modified through a Docker compose file explained in Appendix A-C.
+Additionally, the Radio Simulator dynamically simulates packet loss depending on the predicted elevation angle of the satellite (provided by the Satellite Simulator’s orbital simulation).
+For example, if an attacker sends a telecommand, the Radio Simulator will use a probability value based on the elevation angle and decide to drop or forward the packet.
+4) Ground Configuration. We implemented the ground configuration as settings that can be modified through a Docker compose file explained in Appendix A.
 5) Logging Repository.
-This implementation uses the MongoDB database [67] to aggregate and store logs.
+This implementation uses the MongoDB database [mongodb2024mongodb] to aggregate and store logs.
 The web interface records login attempts, while the Telnet server logs inputs on a per-connection basis.
 Additionally, network traffic to and from the host running HoneySat is captured.
+The logs are secured at points that are not observable by attackers, e.g., different Docker container, shown in Fig.
+6, and the isolated MongoDB database is accessed by an append-only user.
 V-B2 Space Segment Implementation
@@
 1) Flight Software Runtime.
-Although most of the flight software frameworks presented in Section II-D are open-source projects, it is difficult to find a real satellite flight software implementation that is available for study and modification [12].
-Thus, we use the University of Chile’s SUCHAI Flight Software (SUCHAI FS) [52], which has been deployed in four satellite missions [68].
-The SUCHAI FS provides flight software based on the CSP ecosystem that already features a mode where it can be run on a POSIX OS.
+We use the SUCHAI Flight Software (SUCHAI FS) [gonzalez2019architecture], which has been deployed in four satellite missions [Garrido2023] and provides flight software based on the CSP ecosystem.
 2) Flight Software Services.
-The SUCHAI FS uses libCSP [55] and thus features the libCSP default services that present a generic attack surface.
-Furthermore, there is a set of commands commonly used in satellites as a basic set of tasks to control common satellite operations such as communications.
-There are 96 of these commands and parameters that are parsed from CSP packets using a text-based protocol.
-We modified the command handlers to use the Satellite Simulator instead of commanding subsystems and querying real sensors.
+The SUCHAI FS uses libCSP [LibcspLibcsp2024] and thus features the libCSP default services that present a generic attack surface.
+SUCHAI FS includes a set of telecommands which we modified to use the Satellite Simulator instead of querying real sensors.
 3) Satellite Personality.
 The satellite personality was implemented as a Python class that holds multiple variables related to the space segment.
-The current implementation includes almost 50 configurable variables, such as the satellite’s name, the battery’s nominal cell capacity, etc.
+It includes almost 50 configurable variables, such as the satellite’s name.
 4) Logging Repository. Like the ground segment’s logging repository (Sec V-B1), the Satellite Simulator and the FS are connected to the MongoDB instance using a MongoDB client.
-Each logged data entry is created with a customized tag and a timestamp. The collected data is structured JSON [69] format and stored in the MongoDB database.
-Finally, we containerized and hardened this version of the honeypot implementation for secure and easy deployment, which we describe in Appendix A-C.
+Finally, we containerized and hardened this version of the honeypot implementation for secure and easy deployment, which we describe in Appendix A.
+V-C Summary of Existing and Newly Developed Software
+HoneySat’s implementation involved the integration of existing, modified, and newly developed software.
+As-is software includes TigerVNC, MongoDB, and Gpredict.
+Modified software includes the SUCHAI FS, SUCHAI MCS, the Telnet server, and the ZeroMQ Hub.
+Newly developed software includes the Satellite Simulator, including all 6 simulations, the web interface, the flight software services for FS integration, the satellite personality, the ground configuration, and all the necessary framework infrastructure, including Docker files, integration scripts, and configuration files.
 VI Evaluation
@@
   3. Q-3
-    Can HoneySat be extended to simulate different SmallSat missions and support hardware-in-the-loop simulations?
+    Can HoneySat be extended to support different SmallSat missions and hardware-in-the-loop simulations?
     HoneySat’s customization is important because it would allow users of our framework to implement their own honeypots and even support satellite hardware integration.
@@
 This experiment seeks to answer Q-1 by quantifying the interactions provided by HoneySat.
-To achieve this, we leveraged the SPACE-SHIELD matrix (Version 2.0) [63] discussed in Sec.
-II-F.
-SPACE-SHIELD provides a collection of adversary tactics and techniques for space systems.
-In this experiment, we determined the number of tactics and techniques that HoneySat supports.
+To achieve this, we leveraged the SPACE-SHIELD matrix (Version 2.0) [esa2024space-shield] which provides a collection of adversary tactics and techniques for space systems.
+We determined the number of tactics and techniques that HoneySat supports.
 SPACE-SHIELD consists of 14 tactics and 62 techniques.
@@
 Taking this into consideration, 14 tactics are applicable to HoneySat.
-From these tactics, HoneySat can feasibly offer up to 38 techniques as possible interactions for adversaries.
+From these tactics, HoneySat can feasibly offer up to 38 techniques as interactions for adversaries.
 Experiment Description.
@@
 The exploits were simple, using one TC, or complex with multiple TCs involved.
-For example, the technique System Service Discovery (ID T1007)’s description reads “Adversaries may try to gather information about registered local system services.
-Adversaries may obtain information about services using tools and OS utility commands.
-Adversaries may use the information from System Service Discovery during automated discovery to shape follow-on behaviors, including whether or not the adversary fully infects the target and/or attempts specific actions.”
+For example, the technique System Service Discovery (ID T1007) involves adversaries obtaining information about services using tools and OS utility commands.
 Based on this description, we designed the following exploit to capture information about the satellite’s running processes:
@@
 Experiment Results.
-We successfully crafted and ran an interaction or exploit on HoneySat for 33 techniques out of the feasible 38 as depicted in Table III.
+We successfully crafted and ran an interaction or exploit on HoneySat for 33 techniques out of the feasible 38 as depicted in Table II.
 The remaining 5 techniques were not implemented due to limitations in HoneySat’s implementation.
 For example, the Retrieve TT&C master/session keys (T2015.002) technique depends on a cryptographic protocol implementation which HoneySat’s flight software currently does not support.
-This is not surprising, as satellite flight software rarely if ever does [12].
-Due to space limitations, we do not describe all the exploits and interactions here. However, the complete exploit and interaction list is available in Table VIII in Appendix A.
+Due to space limitations, we do not describe all the exploits and interactions here.
+However, the complete exploit and interaction list is available in Table VIII in Appendix B.
 The key findings of our experiment are shown below, providing strong evidence for answering question Q-1 in the affirmative and design objective DO-1 as achieved.
 Key findings Q-1 HoneySat supports 87% of the SPACE-SHIELD matrix techniques possible in a virtual satellite honeypot. HoneySat supports 100% of the SPACE-SHIELD matrix tactics.
-TABLE III: Tactics and techniques supported by HoneySat compared to the ones included in the SPACE-SHIELD matrix that are applicable to virtual honeypots.
+TABLE II: Tactics and techniques supported by HoneySat.
 Tactics SPACE-SHIELD Techniques (Applicable to Virtual Honeypots) HoneySat Supported Techniques
@@
 I, satellites, including SmallSats, are very diverse.
-Second, there is no established metric or tool, such as Nmap’s OS detection [70], to quantify the level of realism of our honeypot.
+Second, there is no established metric or tool, such as Nmap’s OS detection [nmap2025nmap], to quantify the level of realism of our honeypot.
 To evaluate the realism and deception capabilities of HoneySat, we surveyed experienced SmallSat operators.
 Because operators interact with real-world SmallSat missions on a daily basis they are experts and thus are the best population to rigorously evaluate HoneySat.
-TABLE IV: Summary of questions in Section VI-C, satellite honeypot operation tasks of the survey.
+TABLE III: Summary of questions in Section VI-C, satellite honeypot operation tasks of the survey.
 Satellite Operation Evaluated Component No. Qts.
@@
 We divided our survey into five sections.
-The first section collected the participants’ background information including demographic data.
-The second section focused on their professional experience, including whether they worked in industry, government, or academia and to self-report their satellite-related and cybersecurity skills (e.g., satellite hardware engineering).
-The third section probed participants’ satellite operation experience, including how many missions they had operated and which tools they used.
+The first section collected the participants’ background information, e.g., demographic data, the second section focused on their professional experience, the third section probed participants’ satellite operation experience, e.g., how many missions they had operated.
 The fourth survey section involved giving participants hands-on access to a live instance of HoneySat.
 In this section, participants were asked to perform 4 hands-on tasks that correspond to specific satellite mission operations using a live instance of HoneySat as depicted in Fig.
-4.
+5.
 Each hands-on task was designed to feature different components of HoneySat by replicating a real-world satellite mission operation, as discussed in Section II.
 After participants completed each hands-on task, they answered questions designed to assess how realistic they found different aspects of the honeypot.
-Table IV lists the satellite operations included in the survey.
-Finally, in section five, once participants had interacted with multiple elements of our honeypot via these tasks, the survey concluded with participants answering questions about HoneySat’s overall realism and deception.
-Questions related to both the honeypot operation tasks and the overall evaluation used a 5-point Likert scale [71] to measure the participants’ reactions.
+Table III lists the satellite operations included in the survey.
+Finally, in section five, once participants had interacted with multiple elements of our honeypot, the survey concluded with participants answering questions about HoneySat’s overall realism and deception.
+Questions related to both the honeypot operation tasks and the overall evaluation used a 5-point Likert scale [likert2025qualtrics] to measure the participants’ reactions.
 VI-C2 Participants
-We conducted the survey via Qualtrics and Zoom videoconferencing and distributed the survey directly to experienced SmallSat mission operators from previously identified missions.
+We conducted the survey via Qualtrics and Zoom videoconferencing and distributed the survey directly to operators from previously identified missions.
 In total, we received responses from 10 satellite operators who have between 1 to 10 years of experience operating satellites and have operated between 1 to 5 unique missions.
@@
 Over a span of four months, we reached out to national and international institutions, as well as private corporations involved in operational satellite missions to identify suitable participants.
+Despite these challenges, our sample is diverse, including operators from Europe, North America, and South America, spanning industry, government, and academia, across 5 missions.
 VI-C3 Methodology and Key Results
@@
 First, whether the ground segment simulations are realistic; second, whether the space segment simulations are realistic; and third, whether HoneySat, as a whole, provides a convincing and realistic SmallSat mission simulation.
-Before describing the results, it is important to emphasize that the participants were informed that the system they interacted with was a simulation of a satellite mission and not a real one.
-This is because it was not feasible to synchronize participants’ availability to take the survey with the timing of the real satellite’s pass, which may happen only a few times per day and lasts only a few minutes.
+Before describing the results, it is important to emphasize that the participants were informed that the system they interacted with was a simulation.
+We informed participants for two reasons.
+First, it was not feasible to synchronize participants’ availability to take the survey with the timing of the real satellite’s pass, which may happen only a few times per day and lasts only a few minutes.
+Second, claiming to provide access to a real satellite would itself be seen as unrealistic.
 Ground Segment Realism.
 To understand if HoneySat’s ground segment is realistic, participants interacted with HoneySat by performing different satellite operations (discussed in Sec.
-II-C), to showcase the ground segment components listed in Table IV.
+II-C), to showcase the ground segment components listed in Table III.
 One of these tasks is the pass prediction operation which involves calculating when and where a satellite will be within the communication range of a specific ground station.
@@
 90% of the participants agreed or strongly agreed that the telemetry shown during the live telemetry download operation resembles that of a real mission, with the temperature telemetry achieving a 100% strongly agree response rate with some participants mentioning that the temperature values resemble a normal operation and align with the expected temperature values of a satellite.
-For the EPS telemetry, 80% strongly agreed, 10% agreed, and the last 10% neither agreed nor disagreed when asked if they perceived the EPS telemetry as realistic.
-In this case, one participant pointed out that in their specific mission they have more batteries, which is expected as satellite missions are highly diverse as we discussed in Sec.
-II-C.
 These results indicate that the telemetry generated by the Satellite Simulator is considered highly realistic by the vast majority of participants.
-We also asked the participants to assess the ping test operation.
-This operation is a diagnostic procedure used to verify the communication link between a ground station and the satellite.
-It involves sending CSP packets between two CSP nodes, namely the ground segment and the spacecraft.
-After the participants completed this operation’s hands-on task on HoneySat, we asked them to rate their perceived level of realism.
-70% of the participants strongly agreed and 30% agreed that the ping test operation resembles that of a real mission with some participants citing that communication was only possible when the satellite was performing a pass.
-The results of the ping test operation indicate that HoneySat’s radio simulator (which controls the communication between the ground and space segments) is considered to realistically simulate the communication of a real satellite mission by all participants.
 In summary, according to these results, the vast majority of participants, who are experienced satellite operators, perceive the space segment simulated by HoneySat as highly resembling a real satellite mission.
@@
 Additionally, when asked if they would be able to distinguish between HoneySat’s satellite mission simulation and a real mission, 70% strongly agreed, 20% agreed and 10% neither agreed nor disagreed that they would not be able to and again to emphasize that these are experienced satellite operators.
-Overall, our survey results provide convincing evidence for answering question Q-2 in the affirmative and design objective DO-2 as achieved.
-Key finding Q-2 90% of surveyed experienced satellite operators agreed or strongly agreed with the following statement: “I would not be able to distinguish between HoneySat’s satellite honeypot system from the real CubeSat satellite mission it is based on.”
+Overall, our survey results provide evidence for answering question Q-2 in the affirmative and design objective DO-2 as achieved.
+These findings serve as empirical evidence of HoneySat’s fidelity when simulating a real satellite mission.
+Our survey’s questions and results are available online [^3] https://github.com/HoneySat/honeysat-survey-data..
+Key finding Q-2 90% of surveyed satellite operators agreed with the statement: “I would not be able to distinguish between HoneySat’s satellite honeypot system from the real CubeSat satellite mission it is based on.”
 VI-D Internet Interaction Experiment
-This experiment explores HoneySat’s capabilities to entice external actors to interact with it by deploying our honeypot over the Internet.
+This experiment explores HoneySat’s capabilities to entice external actors by deploying it over the Internet.
 Experiment Methodology.
-We leveraged HoneySat’s customization and deployability features to deploy five instances of HoneySat over the Internet and exposed the necessary TCP ports used by HoneySat’s exposed network protocols discussed in Sec.
-IV-C, e.g., Telnet’s port 23.
+We leveraged HoneySat’s customization and deployability features to deploy five instances of HoneySat over the Internet and exposed TCP ports 23 for Telnet, 80 for the web interface, and 5901 for VNC.
+Deployment.
 Initially, we deployed our first HoneySat instance to simulate the SUCHAI-2 SmallSat on the premises of the University of Chile, where the mission’s real ground segment is.
@@
 Therefore, in January 2025, we deployed four additional HoneySat instances on cloud servers.
-We configured two instances to simulate NASA’s Advanced Composite Solar Sail System (ACS3) CubeSat [72], and two simulating DLR’s PIXL-1 CubeSat [73].
+We configured two instances to simulate NASA’s Advanced Composite Solar Sail System (ACS3) CubeSat [nasa2025acs3], and two simulating DLR’s PIXL-1 CubeSat [eoportal2025pixl1].
 We selected these two particular CubeSats because they use the CSP ecosystem (Sec.
@@
 We ensured these additional deployments were believable by configuring their satellite personalities and giving them public IP addresses in the regions belonging to the satellite mission origin, namely Germany and the United States.
-Table V summarizes all five HoneySat deployment details.
-TABLE V: HoneySat Internet deployments’ details.
+Table IV summarizes all five HoneySat deployment details.
+While IP assignment can affect our honeypots’ covert status [lopez2022honeyplc], the increasing use of cloud and web-based ground segment services, such as AWS Ground Station [amazon2024satellite] and YAMCS [yamcs], has made cloud-hosted satellite operations increasingly commonplace.
+As a result, a cloud deployment no longer inherently signals inauthenticity to adversaries.
+Furthermore, the part of the Ground Segment that we expose on cloud IPs is not dependent on being collocated with the GS and thus may as well be hosted with a cloud provider.
+TABLE IV: HoneySat Internet deployments’ details.
 Deployment
@@
 Cloud PIXL-1 Germany 6
-Cloud ACS3 United States 6
-Cloud ACS3 United States 6
+Cloud ACS3 USA 6
+Cloud ACS3 USA 6
 On-prem SUCHAI-2 University of Chile 12
-TABLE VI: Exposed Telnet interactions received.
+TABLE V: Exposed Telnet interactions received.
+Date
 Satellite
@@
 Received
-Engaged
 Time
-ACS3 United States Egypt 3 2 hr
-PIXL-1 Germany Sweden (Tor Node) 6 5 min
-ACS3 United States France (Tor Node) 5 4 min
-ACS3 United States USA 8 3 min
+Jan 18, 2025 ACS3 USA Egypt 4 2 hr
+Jan 24, 2025 PIXL-1 Germany Sweden (Tor) 6 4 min
+Jan 24, 2025 ACS3 USA France (Tor) 9 5 min
+Apr 3, 2025 ACS3 USA USA 8 3 min
 Experiment Results.
-Three of our honeypots successfully enticed external actors and captured four distinct interaction sessions (show in Table VI) via the exposed Telnet protocol discussed in Sec.
-IV-C.
-The four interaction sessions yielded 22 satellite flight software-specific commands.
-These flight software-specific commands show that the adversaries that interacted with HoneySat were purposefully trying to exploit our honeypot using commands that the flight software recognized, not just haphazardly sending random commands.
+We collected several gigabytes of network traffic data on ports 23, 80 and 5901 for each honeypot, however most of these data was generated by crawlers and bots and not satellite-specific.
+But, on January 18th, 2025, we received our first space-specific interaction via Telnet and in the next few months we received more.
+Overall, three of our honeypots successfully enticed external actors and captured four distinct interaction sessions (shown in Table V) all via Telnet.
+The four interaction sessions yielded 22 flight software-specific commands.
+These flight software-specific commands show that the adversaries that interacted with HoneySat were purposefully trying to exploit our honeypot using commands that the flight software recognized.
+Across the four Internet interaction sessions we observed, the first lasted two hours and originated from a non-Tor IP in Egypt.
+The next two sessions occurred within a week, shared their command set (e.g., help, ls, and fp_show), and transitioned to Tor exit nodes in Sweden and France.
+These three factors suggest that the same adversary returned using anonymized infrastructure after initial reconnaissance.
+The first session may have ended after the actor gathered sufficient information.
+In contrast, the fourth and final interaction in April, separated by over two months, from a non-Tor U.S.
+address, and using a very different command set appears unrelated and likely originates from a different actor.
+Overall, the first three interactions are best interpreted as a single adversary adapting its anonymity strategy, whereas the last session was likely initiated by a second actor.
 We now describe one of the most complex interaction sessions HoneySat captured and analyze it in terms of the SPACE-SHIELD matrix (discussed in Sec.
@@
 However, after analyzing all four interaction sessions, we identified the Spacecraft’s Components Discovery (ID T2034) technique, totaling six SPACE-SHIELD techniques based on real-world data captured by HoneySat after successfully enticing and deceiving human adversaries.
-It is worth noting that satellites and space systems in general are far less common compared to CPS such as ICS and UAVs making these interactions collected by HoneySat all the more valuable.
-Additionally, these interactions were produced by human adversaries and not crawlers.
-We know this because HoneySat’s exposed Telnet was implemented so that a human had to type “activate” before sending commands effectively filtering out Internet Telnet bots.
+While our dataset is relatively small, its significance becomes evident when contextualized.
+For example, according to a 2018 Open Platform Communications (OPC) group report [arc2018opcinstalledbase], there are an estimated 47 million OPC-enabled ICS devices deployed worldwide, compared to 11,000 operational satellites in 2025 [livescience2024satellites].
+This represents about 4,700 ICS devices for every one satellite.
+In other words, attackers have orders of magnitude more opportunities to target ICS devices than satellites.
+Thus, making our dataset a rare and valuable contribution to understanding space systems’ TTPs.
+In addition to its rarity, our dataset is the first empirical real-world dataset of TTPs against satellites.
+As such, it not only provides the first view into how adversaries interact with satellite systems, but also establishes a reproducible baseline reference for future studies.
+Lastly, our contribution complements prior work that provides limited insights, e.g., no specific telecommands [uscc2011annualreport].
 In summary, these results provide strong evidence for answering question Q-2 in the affirmative and design objective DO-2 as achieved.
 Key findings Q-2 HoneySat deceived human adversaries who sent 22 satellite flight software-specific commands.
-The interactions collected by our HoneySat Internet deployments account for six SPACE-SHIELD matrix techniques.
+The interactions collected by our HoneySat deployments comprise six SPACE-SHIELD techniques.
 VI-E Case Study: Extending HoneySat to CCSDS Ecosystem
-TABLE VII: SmallSat ecosystems (2) and example SmallSats (3) integrated to HoneySat.
+In Sec.
+V-B we described how we implemented HoneySat using one SmallSat protocol ecosystem, namely, CSP.
+In this case study, we are interested in testing the extensibility capabilities of HoneySat to support additional ecosystems (discussed in Sec.
+II-E) by adding a second ecosystem to HoneySat, namely the CCSDS ecosystem.
+Experiment Description.
+We selected CCSDS because it is a standard protocol suite used by other SmallSats [willbold2023space].
+To accomplish this, we leveraged an open-source CCSDS ecosystem-based flight software framework, RACCOON OS [raccoonos], and YAMCS, an open-source Mission Control software framework with built-in support for PUS [yamcs].
+Experiment Methodology. Building upon our HoneySat framework implementation, as detailed in Sec. V, we enhanced the system by integrating various components of the RACCOON OS and the YAMCS framework.
+Regarding the ground segment, we implemented the exposed network protocol using YAMCS’ built-in web interface.
+For the mission control software, we used YAMCS’s built-in Mission Control Software [yamcs].
+For the radio simulator, we used RACCOON’s communication application.
+For both the ground configuration and the logging repository, we again used YAMCS built-in features.
+For the space segment, we implemented the flight software runtime using the RACCOON framework.
+For the flight software services, we configured the RACCOON framework to connect it to the YAMCS’s MCS on the ground segment.
+The satellite personality and logging repository were based on the existing HoneySat implementations.
+Experiment Results.
+We successfully extended HoneySat to support the CCSDS ecosystem.
+The implementation was completed by a graduate student with no prior satellite-related experience in the span of two weeks.
+However, the majority of the effort involved in extending HoneySat to support the CCSDS ecosystem was dedicated to understanding the ecosystem itself, and analyzing the RACCOON flight software code and the YAMCS framework documentation.
+The only extra implementation that was required was the flight software services which we modified using Rust.
+Other than that, we reused several modules such as the Satellite Simulator.
+In summary, these results provide evidence for answering Q-3 in the affirmative and DO-3 as achieved.
+Key findings Q-3 Out of the box HoneySat supports CSP and CCSDS, the two most widely used space ecosystems, and was evaluated by simulating 3 real-world SmallSats.
+TABLE VI: SmallSat ecosystems (2) and example SmallSats (3) integrated to HoneySat.
 Honeypot
@@
 Institution
-HoneySat CSP
-ACS3
-NASA
-PIXL-1
-DLR
-SUCHAI-2
-U of Chile
-CCSDS
-PUS
-N/A
-In Sec.
-V-B we described how we implemented HoneySat using one SmallSat protocol ecosystem, namely, CSP.
-In this case study, we are interested in testing the extensibility capabilities of HoneySat to support additional ecosystems (discussed in Sec.
-II-E) by adding a second ecosystem to HoneySat, namely the CCSDS ecosystem.
-Experiment Description.
-We selected the CCSDS ecosystem because it is a standard protocol suite used by other SmallSats [12].
-To accomplish this, we leveraged an open-source CCSDS ecosystem-based flight software framework, RACCOON OS [74], and YAMCS, an open-source Mission Control software framework with built-in support for PUS [75].
-Experiment Methodology. Building upon our HoneySat framework implementation, as detailed in Sec. V, we enhanced the system by integrating various components of the RACCOON OS and the YAMCS framework.
-Regarding the ground segment, we implemented the exposed network protocol using YAMCS’ built-in web interface (Fig.
-8).
-For the mission control software, we used YAMCS’s built-in Mission Control Software [75].
-For the radio simulator, we used RACCOON’s communication application.
-For both the ground configuration and the logging repository, we again used YAMCS built-in features.
-For the space segment, we implemented the flight software runtime using the RACCOON framework.
-For the flight software services, we configured the RACCOON framework to connect it to the YAMCS’s MCS on the ground segment.
-The satellite personality and logging repository were based on the existing HoneySat implementations.
-Experiment Results.
-We successfully extended HoneySat to support the CCSDS ecosystem.
-The implementation was completed by a graduate student with no prior satellite-related in the span of two weeks.
-However, the majority of the effort involved in extending HoneySat to support the CCSDS ecosystem was dedicated to understanding the ecosystem itself, and analyzing the RACCOON flight software code and the YAMCS framework documentation.
-The only extra implementation that was required was the flight software services which we modified using Rust.
-Other than that, we reused several HoneySat modules such as the satellite personality and the Satellite Simulator.
-In summary, these results provide strong evidence for answering question Q-3 in the affirmative and design objective DO-3 as achieved.
-Key findings Q-3 Out of the box HoneySat supports CSP and CCSDS, the two most widely used standard space ecosystems, and was evaluated by simulating 3 real-world SmallSats.
+HoneySat CSP ACS3 NASA
+PIXL-1 DLR
+SUCHAI-2 U of Chile
+CCSDS PUS N/A
 VI-F Case Study: Hardware-in-the-loop Experiment
@@
 In this experiment we integrated HoneySat with an in-orbit, operational SmallSat mission.
-To achieve this, we collaborated with Quantum Galactics GmbH [76], an aerospace company that develops hardware subsystems for SmallSats.
-The experiment involved sending a telecommand from HoneySat’s ground segment simulation which was then sent to the in-orbit SmallSat, which would then process the telecommand and respond with sanitized telemetry.
-Due to confidentiality and security restrictions, we are unable to disclose the name of the in-orbit SmallSat which we refer to as SmallSat X.
+To achieve this, we collaborated with an aerospace company that develops hardware subsystems for SmallSats.
+The experiment involved sending a telecommand from HoneySat’s ground segment simulation which was then sent to the in-orbit SmallSat, which would then process the telecommand.
+Due to security restrictions, we are unable to disclose the name of the in-orbit SmallSat which we refer to as SmallSat X.
 Experiment Methodology.
@@
 Specifically, we deployed a proxy server that connected the Satellite Simulator to the SmallSat X’s mission production YAMCS instance.
-Conversely, the SmallSat X’s environment was customized by Quantum Galactics’ team by deploying a script in their YAMCS instance to complete the integration.
-Once the integration was completed, we coordinated with Quantum Galactics to send a telecommand from HoneySat during one of SmallSat X’s passes.
+Conversely, the SmallSat X’s environment was customized by aerospace company’s team by deploying a script in their YAMCS instance to complete the integration.
+Once the integration was completed, we coordinated with the aerospace company to send a telecommand from HoneySat during one of SmallSat X’s passes.
 Experiment Results.
 The telecommand successfully reached SmallSat X during one of the passes and HoneySat received the appropriate telemetry.
-Due to security restrictions, we are not able to disclose further details on the telecommand and the telemetry received.
+Due to security restrictions, we are not able to disclose details on the telecommand and the telemetry received.
 However, the entire communication was initiated by HoneySat which in turn received sanitized telemetry from SmallSat X, effectively closing the loop in the HIL experiment.
-Additionally, Quantum Galactics provided us with a snippet of decoded radio signal data shown in Fig.
-5, which confirms that the experiment was successful.
+Additionally, the aerospace company provided us with a snippet of radio signal data shown in Fig.
+7, which confirms that the experiment was successful.
 In summary, these results highlight HoneySat’s robust extensibility features and provide strong evidence for answering question Q-3 in the affirmative and design objective DO-3 as achieved.
-Figure 5: Waterfall plot that shows the radio signal spectrogram when HoneySat communicated with SmallSat X.
-The signal on top (blue box) belongs to the telecommand sent by HoneySat, and the signal on the bottom (red box) belongs to the telemetry received from the SmallSat X.
+These findings serve as empirical evidence of HoneySat’s fidelity when interacting with a real satellite mission.
 Key findings Q-3 HoneySat was successfully integrated into a satellite hardware-in-the-loop operation and communicated with an in-orbit, operational SmallSat.
 VII Discussion and Future Work
-HoneySat’s Generalizability.
-HoneySat was designed to be extensible and modular as we discussed in Sec.
-IV-A which makes our framework generalizable to additional SmallSat missions.
-We showed this by extending HoneySat to support the two most widely used SmallSat protocol ecosystems, namely CSP and CCSDS (Sec.
-VI-E), two flight software, SUCHAI and RACOON (Sec.
-VI-E), and we customized HoneySat to simulate three different SmallSats (Sec.
-VI-D).
-Additionally, we provide a plethora of customization tools including the Satellite Simulator (Sec IV-D), the satellite personality configuration, the ground configuration, and multiple exposed network protocols (Sec V-B2).
-Finally, in Table IX, we provide examples of different extensibility use cases, including high-level steps and the HoneySat modules involved so that researchers can use and extend HoneySat for their own experiments.
+Challenges of Creating the First Satellite Honeypot.
+During HoneySat’s design and implementation we encountered and overcame two main challenges which stemmed from fundamental differences between satellites and other Cyber-Physical Systems (CPS), e.g., ICS.
+First, time and link-constrained communication.
+Unlike ICS honeypots, which assume continuous network connectivity and stable control loops, e.g., scan cycles, satellite communication occurs only during orbital passes with variable duration and packet loss.
+To overcome this challenge, we developed the Radio Simulator discussed in Sec.
+IV-C.
+Second, dynamic physics-aware simulation.
+ICS honeypots simulate sensor readings through simple physics loops (e.g., tank level, valve position).
+In contrast, a satellite honeypot must dynamically maintain physically coherent telemetry across interdependent subsystems such as attitude, and power.
+We addressed this by creating the Satellite Simulator, which simulates subsystems and sensors that communicate between each other and generate data in real time.
+Satellite Honeynets.
+To increase attacker engagement and reflect emerging satellite networks [ma2023network], multiple HoneySat instances can be deployed as a honeynet [icsnet2024salazar].
+Each instance could simulate different satellites and communicate through virtual inter-satellite links (ISLs).
+In addition, HoneySat could also be part of a honeynet connected to other ground mission infrastructure such as database servers and radio equipment.
+Due to HoneySat’s extensibility, these honeynet scenarios could be implemented in the future.
+Satellite Honeypots’ Fingerprinting.
+Based on our experience designing, implementing and evaluating HoneySat we now discuss some qualitative detection vectors that can guide future satellite honeypot fingerprinting research.
+Space Protocol Uniformity.
+The use of default space protocol configurations may be used for fingerprinting.
+For example, CSP deployments built directly from libCSP expose identical port assignments and diagnostic services, e.g.
+ping.
+Perfectly periodic pass timing.
+Real satellites rarely follow identical pass schedules; orbital perturbations, TLE drift, and operational delays introduce small timing variations.
+A honeypot that enables communication at perfectly fixed intervals or with constant latency can thus be fingerprinted.
+To avoid this, HoneySat periodically downloads the latest TLE data to keep the orbital simulation from drifting.
+This, on the other hand, causes a jump in the orbit once the new TLE data is used to base the simulation on.
+Telecommand implementation.
+A given FS may have completely different telecommand names and functions depending on the satellite and thus a mismatch may reveal the honeypot.
+Additionally, telecommands may have revealing features that print information about the underlying honeypot host, e.g., Ubuntu, which may reveal the simulation.
+VII-A Initial Satellite Honeypot Anti-Detection Techniques
+Building upon lessons learned from HoneySat’s development and from other CPS honeypots [uitto2017survey, tay2023taxonomy], we identify concrete strategies to enhance HoneySat’s and future satellite honeypots’ resistance to fingerprinting and detection.
+Improve Environmental Diversity
+Previous research shows that static structural layouts and repeated host fingerprints enable easy detection [tay2023taxonomy].
+Future satellite honeypots and honeynets can incorporate diversity in its ground segment topology, and network services.
+Following the Purdue-style segmentation [298052] used in ICS honeypots, different HoneySat modules (mission control, station control, payload) can be isolated behind realistic network layers such as routers and firewalls.
+Implement Adaptive Reconfiguration
+Honeypot adaptive reconfiguration occurs when a honeypot recognizes that it is being probed or scanned, e.g., Nmap probes, and reconfigures itself to avoid detection [gabrys2024honeyganpotsdeeplearning, uitto2017survey].
+Adapting this technique to satellite honeypots would involve both the ground and space segments.
+For example, the honeypot could modify exposed services, e.g., CSP over ZeroMQ to appear as a different ground station node.
+Additionally, different telemetry datasets could be rotated, or switched between low interaction and high interaction versions of the honeypot [kyung2017honeyproxy].
+Use Real hardware and software identifiers
+Future satellite honeypots could use real hardware identity and fingerprints such as MAC Organizationally Unique Identifiers (OUIs), and telemetry headers such as APIDs (Application Process Identifiers) [SpacePacketProtocol2020].
+Obtaining these details is non-trivial as they are often proprietary or absent from public documentation, and may require reverse-engineering, or careful analysis of captured mission traffic.
 Limitations.
 Currently, the functionality of some subsystems within HoneySat’s Satellite Simulator are constrained by the quality of the data provided.
-For example, the resolution of Earth’s images generated by our camera payload depends on the resolution of its source data (USGS [77]).
+For example, the resolution of Earth’s images generated by our camera payload depends on the resolution of its source data (USGS [geological2022earthexplorer]).
 Consequently, creating a honeypot for a satellite with a high-resolution camera payload would not be feasible without addressing the underlying issues of data source quality.
@@
 Although we have yet to witness a Stuxnet-like cyberattack on space systems, security researchers need to develop effective countermeasures to secure satellites.
-In this paper, we introduced HoneySat, the first satellite honeypot that realistically simulates an entire satellite mission and provided strong evidence that shows that our honeypot can obtain real-world interaction data and can be extended and customized with real-world satellite software and hardware.
-Finally, we hope that security researchers use HoneySat’s open-source implementation as a foundation not only for satellite honeypot deployments but also for simulation, education, and training applications.
+In this paper, we introduced HoneySat, the first satellite honeypot that simulates a satellite mission and provided evidence that our honeypot can obtain real-world interaction data and can be extended with real-world satellite software and hardware.
+Finally, we hope that security researchers use HoneySat’s open-source implementation as a foundation not only for satellite honeypot deployments but also for simulation, and training applications.
+Acknowledgments
+We thank our shepherd and the anonymous reviewers for their helpful suggestions towards improving this paper.
+The project underlying this paper was funded by the Federal Ministry of Transport (BMV) under the code 45AVF5A011 and was partially supported by NSF awards No.
+2131263 and 2232911, and by the US Department of Transportation (USDOT) CYBER-CARE Grant No.
+69A3552348332.
+We further thank the Saarbrücken Graduate School of Computer Science for their funding and support.
+The authors are responsible for the content of this publication.
 IX Ethical Considerations
-In this paper, we consider the ethical consequences and possible negative outcomes of two main elements of our research.
-The satellite operator user study and the multiple deployments of our satellite honeypot exposed to the Internet.
-We now discuss the stakeholders, potential risks, and how we mitigated those risks for each of these elements.
+In this paper, we consider the ethical consequences and possible negative outcomes of the satellite operator user study and the HoneySat deployments.
+We now discuss the stakeholders, potential risks, and how we mitigated those risks.
 IX-A Satellite Operators User Study
-Prior to conducting any research, the study plan, survey protocol, and associated materials were submitted for review by our Institutional Review Board (IRB), which evaluated the study’s potential risks to participants.
-The protocol received ”exempt status” from the IRB, indicating that the research involves no more than minimal risk.
-This status reflects our careful evaluation of possible harms, which we discuss further below.
-Stakeholder Identification.
-The primary stakeholders in the survey are survey participants.
-Our study aims to gather voluntary responses regarding the fidelity and realism of our honeypot system.
-We carefully considered the potential risks to both the participants of the survey.
-Risk Mitigation.The main risk for the survey participants is the breach of privacy.
-To mitigate this risk, we ensured that no identifiable data (e.g., names, contact details) was collected from survey participants, and the survey responses remain fully anonymous.
+Prior to conducting any research, the survey protocol, and materials were submitted for review by our Institutional Review Board (IRB).
+The protocol received “exempt status”, indicating no more than minimal risk.
+Stakeholder Identification. The primary stakeholders in the survey are survey participants.
+Risk Mitigation.
+The main risk for the survey participants is the breach of privacy.
+To mitigate this risk, we ensured that no identifiable data (e.g., names) was collected, and the responses remain anonymous.
 Additionally, the survey itself includes an informed consent section that informs the respondents about the voluntary participation, the survey’s purpose, a description of the procedures, the risks involved, contact information, and the option to opt out of the survey.
-Finally, we provided participants the choice to skip questions using the response options “Prefer not to say” and “I do not know” to applicable questions.
-Due to the aforementioned risk mitigation strategies, the stakeholders’ breach of privacy risk is negligible.
+Finally, we provided participants the choice to skip questions using the response options “Prefer not to say” and “I do not know.”
 IX-B Honeypot Deployment Experiment
@@
 The primary stakeholders in the honeypot deployment are the external actors who interact with our honeypot and the infrastructure owners on whose infrastructure our honeypot is running.
-In the honeypot deployment experiment we deployed HoneySat on server infrastructure owned by real-world institutions.
 Risk Mitigation.
-The main risk for the infrastructure owners is that one of the external actors will use our honeypot as an stepping stone to breach their production infrastructure.
-For the deployment on the premises of the institutions, we worked with the local IT infrastructure administrators to ensure that we took all the necessary precautions to avoid any possibility that the external actors could gain access to the infrastructure via our honeypot.
-First, the infrastructure owners provided us with a server with a clean OS installation that did not have any production applications or sensitive data.
+The main risk for the infrastructure owners is that external actors will use our honeypot as a stepping stone to breach their infrastructure.
+For the on-prem deployments, we worked with the local IT administrators to ensure that we took all the necessary precautions to avoid this scenario.
+First, we were provided with a server with a clean OS installation that did not have any production applications or sensitive data.
 Second, the server was isolated on its own network segment.
@@
 Fourth, the VNC portion of the honeypot is configured so that external actors cannot use it to host malicious services by denying traffic forwarding.
-Fifth and final, we deployed our honeypot using two sandboxing layers (Docker containers and Virtual Machines (VMs) as we discussed in Appendix A-C.
-Due to the aforementioned risk mitigation strategies, the stakeholders’ breach of privacy risk is negligible.
+Fifth and final, we deployed our honeypot using two sandboxing layers (Docker containers and Virtual Machines (VMs) as we discussed in Appendix A.
 The main risk for the external actors is the breach of their privacy.
-However, it is generally accepted that trespassers of a computer system do not have reasonable expectation of privacy [78].
-In order to mitigate this risk, our honeypot system gives notice and warning to any user connecting to it indicating that ”This computer system is for authorized use only.” Additionally, the web services of our honeypot are protected by a username and password.
-This notice system gives external actors a chance to avoid interaction, which helps reduce the risk.
-Appendix A Appendix
-A-A Interaction Sequences and Exploits
+However, it is generally accepted that trespassers of a computer system do not have reasonable expectation of privacy [spitzner2002honeypots].
+In order to mitigate this risk, our honeypot system gives notice and warning to any user connecting to it indicating that “This computer system is for authorized use only.” Additionally, the web services of our honeypot are protected by a username and password.
+Appendix A Deployment and Security Hardening Implementation
+Figure 6: Architecture of dockerized Generic CSP Honeypot
+As we mentioned in Sec. II-A, high-interaction honeypots such as HoneySat present a high risk of adversary takeover. To mitigate this risk, we implemented HoneySat with two sandboxing layers.
+Virtual Machine. VMs provide the highest isolation level among sandboxing techniques. We implemented HoneySat in VM environments to leverage VMs’ robust security.
+Containerization.
+After we completed HoneySat’s development, we used Docker Compose [docker2024docker] to containerize each of our framework’s applications.
+Specifically, we created four different containers depicted in Fig.
+6.
+Containerizing HoneySat provides two benefits.
+First, it creates another sandboxing layer that prevents adversaries from using our honeypot to access the underlying system [marra2024feasibility].
+Second, it proves a convenient and flexible way to deploy HoneySat.
+TABLE VII: Correspondence between real satellite mission components and HoneySat’s design components.
+Segment Real Component HoneySat Component
+Ground Remote Operation Protocols VNC, Telnet
+Ground Mission Control Software Modified MCS
+Ground Ground Station Control Software Gpredict
+Ground Web Interface Web Interface
+Ground Ground Station Radio Simulator
+Space Flight Software F.S. Runtime + Services
+Space On Board Computer Docker Container
+Space Platform Satellite Simulator
+Space Payload Satellite Simulator
+Appendix B Interaction Sequences and Exploits
 Table VIII includes all the interaction sequences and exploits we performed during the experiments in Sec. VI-B.
@@
 Flight software.
-Exploit OS, libraries or software vulnarabilities.
+Exploit OS, libraries or software vulnerabilities.
 Deploy custom CSP application to forge TC/TM.
@@
 Use commands to upload artifacts modify system settings:
-tm_send_file articaft 1
+tm_send_file artifact 1
 1: obc_mv artifact /etc/config/artifact
@@
 Credential Access Communication Link Sniffing T2044 Ground software.
-Scape the ground software or docker and run tcpdump.
+Escape the ground software or docker and run tcpdump.
 Deploy a CSP application to capture/inject CSP packets.
@@
 Flight software.
-Scape the ground software or docker and run tcpdump.
+Escape the ground software or docker and run tcpdump.
 Deploy a CSP application to capture/inject CSP packets.
@@
 Flight software.
-Deploy an malicious application that sends data over CSP
-Command and Control Telecommand a Spacescraft T2019
+Deploy a malicious application that sends data over CSP
+Command and Control Telecommand a Spacecraft T2019
 Ground software.
@@
 Flight software.
-The atacker deploys a custom CSP node or a backdoor
+The attacker deploys a custom CSP node or a backdoor
 Exfiltration Optical link modification T2037 Threat Model Limitation N.A.
@@
 Flight software.
-Send TC to change communication paramters.
+Send TC to change communication parameters.
 Modify network configuration in the ground station.
@@
 Flight software.
-Send TC to lauch a fork bomb or a reset loop
+Send TC to launch a fork bomb or a reset loop
 1: obc_system :(){ :—:& };:
@@
 Impact Transmitted Data Manipulation T2024 Threat Model Limitation N.A.
-A-B Satellite Simulator Design Details
-Figure 6: High-level architecture of the Satellite Simulator.
-Due to space limitations, we do not show all the available sensors and simulations.
-Some sensors communicate with multiple simulations.
-Although the flight software services are not part of the Satellite Simulator we included it to depict the Interface module’s data flow better.
-As shown in Fig.
-6, the Satellite Simulator includes five module types: simulation modules (light orange), sensor modules (light blue), subsystem state modules (light green), satellite state modules (light gray) and interface modules (light yellow).
-Simulation Modules.
-These modules are abstractions of real-world processes required for a realistic satellite simulation (DP-2).
-An example of these simulations is the orbital simulation.
-This simulation uses orbital mechanics to calculate data such as the satellite’s orbital position at any given time.
-These simulation modules can communicate with each other to facilitate proper functionality.
-For example, the power simulation module will query the orbital simulation for orbital data to determine if the satellite is positioned properly in relation to the sun so that it can draw power from its solar panels.
-The subsystems and the sensors of the Satellite Simulator are designed around these simulations.
-The Satellite Simulator includes the orbital, rotation, power system, thermal, magnetic, and payload simulations.
-Due to space restrictions and their complex design, we explain each simulation in Appendix A-B.
-Overall, these simulations provide the foundation for any satellite operator to use our framework and implement their honeypot using their satellite flight software.
-The Satellite Simulator simulations are configured using the satellite personality, which we discuss below.
-Sensor Modules.
-These modules are abstractions of the hardware sensors used by a satellite, i.e., temperature sensor.
-The sensor modules do not perform any computations, instead they collect data directly from the simulation modules following DP-2.
-For example, the voltage sensor will query the power simulation to collect data about the current state of the battery.
-Subsystem State Modules.
-These modules are abstractions of the satellite subsystems discussed in Sec.
-II-D.
-A given subsystem state is made up of one or more sensors.
-For example, in Fig.
-6 the Payload State (light green) includes the Camera Sensor (light blue).
-In this way, subsystem states serve two purposes.
-They can get data from their sensors or can set a specific configuration on their sensors.
-Satellite State Module.
-This module is an abstraction of an entire satellite.
-As depicted in Fig.
-6 the Satellite State (light gray) is made up of multiple subsystem states.
-Satellite State module routes messages between the Satellite Simulator and the modified flight software.
-For example, if the modified flight software sends a message to the Satellite Simulator requesting to provide the present voltage in the EPS, the satellite state will route that message to the EPS State.
-Interface Module.
-This module is an abstraction of the communication protocol between the modified flight software and the Satellite Simulator.
-As depicted in Fig.
-6, the Interface (light yellow) is the module that connects the flight software services with the Satellite Simulator simulations.
-This protocol must implement two basic message types, requests and replies.
-However, to meet DO-3, the underlying implementation of these messages is left open for the user to decide based on their requirements.
-To finish, the Satellite Simulator’s built-in modular design allows subsystems, simulations, and sensor modules to be mixed ’n matched to meet the requirements of most satellites and their missions.
-Additionally, our flexible design allows users to easily add additional simulations, sensors, or subsystems as modules to the honeypot.
-  • Orbital Simulation.
-This simulation calculates and predicts the satellite’s orbit.
-To achieve this, it uses the two-line element set or TLE data configured in the satellite personality.
-The TLE data contains the necessary orbital information required to calculate the location of a satellite given a point in time [43].
-We rely on the Skyfield python library to perform advanced and precise calculations [79].
-This simulation calculates multiple orbital values which include the altitude angle, latitude, longitude of the satellite among other values.
-The payload, EPS, and magnetic simulations rely on data from this simulation.
-  • Rotation Simulation.
-This simulation simulates the satellite attitude changes to provide the satellite orientation.
-Determining and controlling the satellite orientation is relevant for payloads such as cameras or communication systems that require pointing to the earth surface or to other celestial bodies.
-This simulation calculates the rotation kinematics of the satellite using two reference frames: a reference frame fixed to the satellite body and a non-rotational reference frame.
-The relation between the two reference frames is calculated using the conservation of angular momentum and the rigid-body Euler equation [80].
-To simulate the satellite’s rotation, the initial instantaneous angular velocity of the satellite, the current orientation in quaternions, and the satellite’s inertia matrix must be defined.
-The simulation uses the fourth-order Runge Kutta integration method (RK4) [81] to perform these calculations to avoid divergences in the simulation with time steps greater than 1 second.
-  • Power System Simulation.
-This simulation manages the satellite’s power collection, consumption, and distribution.
-It tracks the battery capacity, the power draw and simulates battery charging whenever the orbital simulation tells it that the satellite is exposed to sunlight.
-A battery pack is modeled as a series or parallel configuration of smaller battery packs.
-The simplest battery pack consists of a single cell.
-A single cell is simulated to calculate the charge and discharge curves, assuming that the cell’s current will not change.
-This cell voltage is then used to calculate the pack’s voltage.
-The electrical power system simulation uses the pack’s voltage together with the input and output power to calculate the battery current.
-In case the input or output power of the power subsystem changes, a new cell current is calculated, and the battery cell simulation is rerun to reflect the change.
-  • Thermal Simulation.
-This simulation calculates the temperature of the satellite.
-The simulation is based on two values: the total thermal energy and a user-defined specific heat capacity.
-Together with emissivity, mass and surface area, the temperature of a satellite can be approximated based on the absorbed sunlight and the energy emitted through thermal radiation.
-This simulation uses a radiation loss formula [82] to calculate the thermal radiation emission.
-  • Magnetic Simulation.
-This simulation tracks and analyzes the interactions between the Earth’s magnetic field and the satellite’s own magnetic environment.
-It communicates with the orbital and rotation simulations to determine the satellite’s position and orientation in space so that it can calculate the Earth’s magnetic field components (north, east, vertical) and total intensity at the satellite’s current location.
-  • RGB Camera Simulation.
-This simulation replicates the functionality of an RGB camera pointed at Earth.
-It uses Earth observation satellite imagery from the EarthExplorer database from the U.S.
-Geological Survey (USGS) [77].
-If a capture image command is issued, it will select an image from one of two pools of images: daytime or nighttime.
-This simulation communicates with the orbital simulation to determine if satellite is in the presence of the sun.
-A-C Deployment and Security Hardening Implementation
-Figure 7: Architecture of dockerized Generic CSP Honeypot
-As we mentioned in Sec. II-A, high-interaction honeypots such as HoneySat present a high risk of adversary takeover. To mitigate this risk, we implemented HoneySat with two sandboxing layers.
-Virtual Machine. VMs provide the highest isolation level among sandboxing techniques. We implemented HoneySat in VM environments to leverage VMs’ robust security.
-Containerization.
-After we completed HoneySat’s development, we used Docker Compose [83] to containerize each of our framework’s applications.
-Specifically, we created four different containers depicted in Fig.
-7.
-Containerizing HoneySat provides two benefits.
-First, it creates another sandboxing layer that prevents adversaries from using our honeypot to access the underlying system [32].
-Second, it proves a convenient and flexible way to deploy HoneySat.
-Figure 8: Screenshot of the HoneySat YAMCS web interface, displaying a TM packet, returned from the Generic CCSDS Mission Honeypot.
-A-D Extensibility Example Use Cases
-TABLE IX: HoneySat example use cases for further extensibility.
-Extensibility Use Case Relevant HoneySat Module(s) High-level Steps
-Add new sensor Satellite Simulator To add a new sensor you can leverage the existing Sensor Abstract Base Class (ABC) included in the Satellite Simulator source code and write the logic for your new sensor.
-Add new payload Satellite Simulator To add a new payload you can leverage the existing PayloadSimulation Abstract Base Class (ABC) included in the Satellite Simulator source code and write the logic for your new payload including where to get the payload data.
-Integrate new flight software Satellite Simulator, ground configuration, satellite personality, radio simulator, flight software services To integrate a new flight software into HoneySat the first step is to analyze and understand the target flight software architecture and source code even before attempting integration.
-Then the target flight software must be integrated using HoneySat’s flight software services module.
-After integrating the flight software services, the Satellite Simulator can be connected using the Interface Python ABC locanted in the Satellite Simulator source code.
-Add a new exposed network protocol Ground Software, Docker compose A new exposed network protocol can be added by modifying the Docker compose file included with HoneySat.
-The compose file allows for new services and TCP ports to be easily added.
+Appendix C Hardware-in-the-loop Experiment Results
+Fig. 7 depicts a snippet of radio signal data which confirms that the hardware-in-the-loop experiment in Sec VI-F was successful. This was provided by the aerospace company.
+Figure 7: Waterfall plot that shows the radio signal spectrogram when HoneySat communicated with SmallSat X.
+The signal on top (blue box) belongs to the telecommand sent by HoneySat, and the signal on the bottom (red box) belongs to the telemetry received from the SmallSat X.
+Appendix D Artifact Appendix
+This appendix accompanies the paper “HoneySat: A Network-based Satellite Honeypot Framework” and provides detailed instructions for obtaining, installing, and evaluating the artifact submitted for NDSS 2026 Artifact Evaluation.
+D-A Description & Requirements
+Our artifact contains a Dockerized version of our honeypot framework, along with all the necessary components to bootstrap honeypots for satellite missions supporting two different space protocol stacks (CSP and CCSDS/YAMCS).
+To aid in evaluations, we also included utility scripts that facilitate the evaluation of data provided by the honeypot.
+D-A1 How to access
+The artifact can be found at https://doi.org/10.5281/zenodo.17871431.
+D-A2 Hardware dependencies
+  • 25GB of disk space
+  • at least 4 CPU cores
+  • 8GB RAM
+D-A3 Software dependencies
+Most Linux distributions will work. We verified the artifact on Ubuntu 24.04.2 LTS.
+  • Docker (with compose and related components, we recommend to follow https://docs.docker.com/engine/install/)
+  • Python 3.12 (more recent versions may also be compatible)
+  • Python virtual environment
+  • Bash
+  • Telnet client
+  • Any modern web browser
+D-A4 Benchmarks
+No External Benchmarks required.
+D-B Artifact Installation & Configuration
+Extract files
+Extract files from tar file downloaded from link:
+tar xzpvf ndss-artifact-eval.tar.gz
+Repository Layout
+The root directory contains two subdirectories:
+  • deployment/: Dockerized HoneySat services with detailed setup documentation.
+  • evaluation/: Python utilities and scripts to help evaluate HoneySat’s capabilities.
+The evaluation utilities require Python 3.12 and the packages listed in evaluation/requirements.txt.
+Create and Activate a Python Virtual Environment.
+From the repository root:
+python3 -m venv .
+source bin/activate
+Install Python Dependencies
+Install the required packages for the evaluation utilities:
+python3 -m pip install -r evaluation/requirements.txt
+Configure Docker user
+We need to add the current user to the Docker group to avoid using sudo with Docker all the time.
+sudo usermod -aG docker \$USER
+sudo reboot now
+Start the Honeypot Services.
+All remaining setup is handled by Docker containers. You can either:
+  • Run the provided convenience scripts (see “Quick Start for Pass Simulations” below), or
+  • Follow the guide in deployment/README.md to bring up the services via Docker Compose.
+Quick Start for Pass Simulations
+A key HoneySat feature is its realistic communication windows: the satellite is reachable only while passing over a ground station.
+Many experiments benefit from a setup in which the satellite becomes reachable immediately or within a short time window.
+We provide two helper scripts that compute a suitable ground-station location along the satellite’s predicted ground track so that a pass occurs shortly after startup.
+Each script builds and starts the necessary Docker Compose services and will shut them down cleanly on CTRL+C.
+  • CSP-based honeypot:
+    ./evaluation/experiment-1/run-experiment-csp.sh
+    This script builds and starts a CSP honeypot instance, and positions the ground station to enable communication almost immediately, allowing you to observe how an attacker would perceive a pass without waiting for a real one.
+  • CCSDS/YAMCS-based honeypot:
+    ./evaluation/experiment-1/run-experiment-ccsds.sh
+    This script builds and starts a CCSDS and YAMCS-based honeypot instance, using the same predicted ground-station placement to trigger an imminent pass.
+Please proceed with the experiments described in the following sections, or consult deployment/README.md for additional configuration details.
+D-C Experiment Workflow
+The following experiments all involve starting the honeypot system and interacting with it in some way (either programmatically or manually).
+The workflow typically involves issuing a startup command and then performing tasks outlined in the experiment description.
+For some tasks, timing matters; for example, the TCs have to be issued while the simulated satellite is reachable via the simulated ground station.
+D-D Major Claims
+We make the following claims in our paper
+  • (C1): HoneySat can be used to deceive adversaries and log activities.
+    • HoneySat’s simulator provides believable TM (Experiment 1.1)
+    • HoneySat has realistic communication windows (Experiment 1.2)
+    • HoneySat provides interaction capabilities (process TC, provide TM) (experiment 1.3)
+    • HoneySat Logs interaction details (Experiment 1.4)
+  • (C2): HoneySat Is extensible and supports two different protocol ecosystems
+    • HoneySat is configurable (Experiment 2)
+D-E Evaluation
+D-E1 Experiment 1 (E1.1)
+[Believable TM/TC]   [10 human-minutes]
+Goal. Demonstrate that HoneySat produces believable telemetry (TM) in response to telecommands (TC) for a CCSDS/YAMCS setup.
+Preparation. Ensure dependencies are installed as listed above.
+Execution.
+./evaluation/experiment-1/run-experiment-ccsds.sh
+After the script completes and prints telemetry values, you may also inspect the telemetry via the YAMCS web interface.
+Expected Results. Believable current, voltage, and temperature values for the selected battery configuration:
+  • Voltage:  8000 mV (normal test case)
+  • Temperature:  30 °C
+  • Current draw:  74 mA
+D-E2 Experiment 1.2 (E1.2)
+[Believable passes]   [20 human-minutes]
+Goal. Show that HoneySat enforces realistic communication windows: the satellite is reachable only during predicted passes over a ground station.
+Preparation.
+./evaluation/experiment-1/run-experiment-csp.sh
+This positions the ground station so that a pass begins approximately 2-3 minutes after startup (adjustable), and starts all required services.
+Execution.
+  1. 1.
+    Open the web interface at http://localhost:80.
+Log in with username admin and password admin.
+Click the ground-station icon and view the next predicted passes.
+The pass should soon show as “ongoing.”
+  2. 2.
+    In a separate terminal, connect to the CSP telnet interface and activate it:
+    telnet localhost 24
+    # In the telnet session:
+    activate
+  3. 3.
+    Probe satellite reachability every few seconds:
+    1: com_ping 10
+Expected Results. The satellite responds to pings only during the predicted pass and not before or after. Responses will indicate the expected addressing (source address 1, destination address 10).
+D-E3 Experiment 1.3 (E1.3)
+[Simulate Interaction]   [15 human-minutes]
+Goal. Demonstrate interactive capabilities once the satellite becomes reachable.
+Preparation. Repeat the setup from Experiment 1.2. Review the command reference in the experiment-1.3 directory.
+Execution. After the satellite becomes reachable (as verified in Experiment 1.2), issue commands via the telnet interface. For example, to execute arbitrary shell commands on the OBC:
+1: obc_system [shell command]
+Expected Results. The experimenter can successfully execute arbitrary shell commands via obc_system while the satellite is reachable.
+D-E4 Experiment 1.4 (E1.4)
+[View logging capabilities]   [5 human-minutes]
+Goal. Verify that HoneySat logs interactions and relevant parameters.
+Preparation. Keep a HoneySat CSP instance running (e.g., from Experiment 1.3), optionally after interacting with it.
+Execution.
+python3 ./evaluation/experiment-1/python_dump_mongodb/dump_mongodb.py
+Alternatively, connect to the MongoDB instance with a database client of your choice.
+Expected Results. The script prints MongoDB contents to stdout, showing logged parameters and interaction details.
+D-E5 Experiment 2 (E2)
+[Customization]
+Goal. Demonstrate HoneySat’s configurability across protocol stacks and scenarios.
+Preparation. Navigate to ./evaluation/experiment-2.
+Execution.
+  1. 1.
+    Create a baseline customization:
+    python3 ./honeysat.py [csp|ccsds] "{satellitename}" "{location}"
+    For example:
+    python3 ./honeysat.py csp "BEESAT" "Berlin"
+  2. 2.
+    Start services with the generated configuration (or configs in this directory):
+    python3 ./honeysat.py start [csp|ccsds] "{satellitename}" "{location}"
+    For example:
+    python3 ./honeysat.py start csp "BEESAT" "Berlin"
+  3. 3.
+    When finished, stop the services:
+    python3 ./honeysat.py stop [csp|ccsds]
+    For example:
+    python3 ./honeysat.py stop csp
+Expected Results. HoneySat can be configured with relatively low time effort to reflect different satellites, locations, and protocol ecosystems.